Warning: Can't find topic HCC.TabsJanP

Create new blog entry for Point


Created by Main.MirkoPojer on 2007-07-10 - 09:31

Beam Dump System - Installation and Commissioning Meeting _ 03/07/07

PDF RAT_20070703.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPointx2007x07x10x09x30/RAT_20070703.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-14 - 18:22

Powering Tests Sector 78 - Report 14.06.2007

PowerPoint RD2-Q4_14.06.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPointx2007x06x14x18x22/RD2-Q4_14.06.ppt

Test Report of Thursday 14.06.2007

  • ElQA
    • RQD.A78 and RQF.A78 circuits released by ElQA. Circuits OK for powering as soon as QPS gives green light.
    • MBA and MBB lines tested OK up to 1 kV with all instrumentation connected. Circuit OK for QPS IST

  • RD2.L8, RQ4.L8 and its correctors
    • Powered to nominal (6000 A) during the whole morning. Ramped down to beam operational current (5980 A) in the begining of the afternoon.
    • Keeping RD2.L8 at operational current, the five Q4 correctors available were ramped to nominal current (72 A) RQ4.L8 is powered acording to the procedure attached (see slide) up to 3.6 kA and 2.95 kA in apertures 1 and 2 respectively. No quench was observed.
    • Pushing for the quench: With D2 at 6kA, aperture 1 of Q4 at 3610 and four correctors at nominal
      • Ramp up from 2950A to 3610 at 1 A/s - OK
      • Ramp down from 3610A to 2950A at 5 A/s -OK
      • Ramp up from 2950A to 3610A at 10 A/s -OK
      • Ramp down from 3610A to 2950 at 10 A/s - OK
      • Ramp up from 2950A to 3610A at 10 A/s - OK
    • Magnet still doesnt quench, so we go on:
      • Ramp up Aperture B1 and B2 from 3610A to 2200A at 10 A/s - OK
      • Ramp up Aperture B2 from 2200A to 3610A at 10 A/s - OK
      • Ramp up Aperture B2 from 2200A to 3610A at 10 A/s - OK

  • RQ5.L8 and its correctors
    • Calibration tests in all the converters

  • 600 A Circuits
    • Very useful data was obtained during the morning on the RQ6.R7 quench detectors
    • Generation and analysis of the QPS data advances in parallel with PCC tests for the available 600A circuits in the DFBAO

  • 60 A type test performed but not finished. Half a day more needed.

Planning for Friday 15.06.2007

  • RQF.A78 and RQD.A78
    • Wait for green light from QPS to go ahead with PIC tests. All heaters and quench loop worked OK, 'Switch Open Failure' functionality to be corrected.
    • PIC1 for both circuits as soon as QPS gives green light.
    • AB/PO will use the characterisation values obtained during ElQA to avoid the connection of a 120A converter for PCC.
    • The DC cables will be connected to the leads once PIC1 is completed.

  • RB.A78
    • QPS IST to start today. They will take around 2 days.

  • 600A Circuits in the DFBAO
    • PCC will continue for the circuits available. QPS will take proffit of them to obtain the needed data for improving the filtering of the detector noise.
    • PIC2 for the circuits in which PCC is completed
    • AB/PO will do some measurements after PIC2 (below 20A)

  • Matching Section ML8
    • 8-hour heat run with all the circuits in the matching section at nominal current
    • Squeezing tests for RQ5.L8 and RQ4.L8 can be carried out in parallel

  • 60 A type test will start after QPS IST for RQF and RQD are completed.

  • Cryo-instrumentation
    • The time needed by QPS to implement the corrections in the 600A quench detector software will be used to test the electronica cards from cryo damaged during the ElQA tests.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-06-11 - 16:01

Powering Tests: Planning for today 11.06

PDF RAT_20070611.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPointx2007x06x11x15x54/RAT_20070611.pdf

PowerPoint _Direct_Quench_Detection_with_Shunt.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPointx2007x06x11x15x54/_Direct_Quench_Detection_with_Shunt.ppt

Test Report of Monday 11/06/2007

  • AB-PO performed "type test" on RPLA.12L8.RCBH12.L8B2
  • Special test on RD2.L8: Quench Heater discharge at 200A (Reiner) - Signal cable to be fixed (Vincent)
  • 600A PC de-consigne: RQT12.L8B1 RQTD.A78.B1 RQTD.A78.B2 RQTF.A78.B1 RQTF.A78.B2 RSD1.A78.B1 RSD1.A78.B2 RQS.L8.B1 RSD2.A78.B1 RSD2.A78.B2 RSF1.A78.B1 RSF1.A78.B2 RSF2.A78.B1 RCD.A78.B1 RCS.A78.B1 RCS.A78.B2
  • AB-CO performed sequencer procedure of 600A circuits - not terminated, half day needed.

Planning for Tuesday 12/06/2007

  • ElQA on 13kA circuits
  • ElQA on 600A
  • Implementation on RQ6.R7 and on L8 600A circutis of the second order monitoring software by AT-MEL - See attached presentation given by Bob Flora during the afternoon RAT in CCC
  • QPS Software test on RQ6.R7
  • Sequencer verification for 600A circuits - second part
  • PCC on 600A when released by AT-MEL
  • PCS on the first bunch of 600A circuits

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-06 - 18:58

Powering Tests Sector 78 - Morning Meeting 06/06/2007

PDF RAT_20070606.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPointx2007x06x06x18x56/RAT_20070606.pdf

Comments:

Edit, attach file or add comment


Created by Main.IvanMoyaMartinez on 2007-11-01 - 15:14

Sector 8-1

Going-on Activities

* It has been decided to repair the flexibles in the DFBAs of Sector 8-1 before the global leak tests, allocated dates are: o DFBAP repair: Wk 44 o DFBAA repair: End of wk 44 - begining of 45

* SSS Interconnection

* Inner Triplets: o Left 1: closed, being leak tested o Right 8: closed, being leak tested

* Electrical non-conformities on instrumentation cables of the power converters will be closed by the end of the week.

* The procedure for cleaning the DFBXH is currently in engineering check. Its expected that the document will be ready for the half of the next week.

Planned activities

* PIC-PC interface test - Wk 45 * Room temperature vacuum and collimator works (wks 43-46), bake-out (wks 47 -48) * ATLAS-RP cable installation - Wks 48-49 * TS/EL: o New water-cooled cable DFBXH: Wk 45 - Saturday & Wk 46 o Replacement of water-cooled cable of RQ6 in the DFBAP - Wk 44 o Week 46: Presentation of the cables and installation of the cable supports on the DFBMs in RA83. * ElQA For the ELQA it was talked today to start the test in the L1 and in the arc at the end of the week 45 and will continue the week 46 also. There will not be interference with other groups, and the conditions for the ELQA will be granted for the date. o TP4-A: Inner triplet L1 - Week 46 o TP4-A: Inner triplet R8 - Week 46

General

* Cooling water shut-down - Wks 42-44 - All the cryogeninc activities will be stopped.

Comments:

Sector 7-8

* It will be possible that the time for the consolidation time will be longer than expected. Because of this it will be possible that the leak test will not be performed

-- Main.IvanMoyaMartinez - 05 Nov 2007

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-10-25 - 09:30

From Pressure Tests tp Cooldown - Sector 8-1 - Thursday October 25, 2007

Pptx Sector81_status_20071025.pptx:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x10x25x08x43/Sector81_status_20071025.pptx

Sector 8-1

Going-on Activities

  • It has been decided to repair the flexibles in the DFBAs of Sector 8-1 before the global leak tests, allocated dates are:
    • DFBAP repair: Wk 44
    • DFBAA repair: Wk 45

  • SSS Interconnection

  • Inner Triplets:
    • Left 1: closed, being leak tested
    • Right 8: closed, being leak tested

  • Electrical non-conformities on instrumentation cables of the power converters will be closed by the end of the week.

  • The procedure for cleaning the DFBXH is currently being completed. Due to unavailabily of the person responsible to do the works, cleaning cannot start on week 45 as expected but will have to be postponed till 46.

Planned activities

  • PIC-PC interface test - Wk 45
  • Room temperature vacuum and collimator works (wks 43-46), bake-out (wks 47 -48)
  • ATLAS-RP cable installation - Wks 48-49
  • TS/EL:
    • New water-cooled cable DFBXH: Wk 45 - Saturday & Wk 46
    • Replacement of water-cooled cable of RQ6 in the DFBAP - Wk 44
    • Week 46: Presentation of the cables and installation of the cable supports on the DFBMs in RA83.
  • ElQA
    • TP4-A: Inner triplet L1 - Week 46
    • TP4-A: Inner triplet R8 - Week 46

General

  • Cooling water shut-down - Wks 42-44 - All the cryogeninc activities will be stopped.

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-10-22 - 16:27

From the Pressure Tests to the Cooldown - Sector 8-1 - October 19, 2007

Sector 8-1

Going-on Activities

  • SSS Replacement

  • Inner Triplets:
    • Left 1: closed, being leak tested
    • Right 8: closed, being leak tested

  • Electrical non-conformities on instrumentation cables of the power converters will be closed by the end of the week.

  • The procedure for cleaning the DFBXH is currently being completed. In principle, the instrumentation wont have to be removed.

Planned activities

  • PIC-PC interface test - Wk 43
  • Room temperature vacuum and collimator works (wks 43-46), bake-out (wks 47 -48)
  • ATLAS-RP cable installation - Wks 48-49
  • TS/EL:
    • DFBXH Ineo cabling - To be done on Monday 22.
    • Installation of the removed water-cooled cable and RQ6 - Flohe Wk 44 - 45 Once the cleaning is completed.
    • NC DFBXH: The water cooled cable associated to the current lead #3 cannot be fixed to the lead on the left side. There is a mechanical interference between the water tube on the lug and a connector necessary for the routing of the protection signal from the intrumentation port. The connector cannot be moved to another position, it is therefore proposed to fix the water cooled cable on the right side of the current lead lug.
  • ElQA
    • TP4-A: Inner triplet L1 - as soon as possible since it will be the first time they are performed
    • TP4-A: Inner triplet R8 - not before the DFBX is repaired
  • Alarm test in the whole sector will be carried out on Wednesday 24.

General

  • Cooling water shut-down - Wks 42-44 - All the cryogeninc activities will be stopped.
  • A latest date for taking a decission about repairing or not the DFBAs has to be set for not jeopardizing the starting date of the flushing

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-10-17 - 17:02

From Pressure Tests to Cooldown - Sector 81 - October 16, 2007

Sector 8-1

Going-on Activities

  • SSS Replacement

  • Inner Triplets:
    • Left 1: closed, being leak tested
    • Right 8: closed, being leak tested

  • Electrical non-conformities on instrumentation cables of the power converters will be closed by the end of the week.

Planned activities

  • PIC-PC interface test - Wk 43
  • Room temperature vacuum and collimator works (wks 43-46), bake-out (wks 47 -48)
  • ATLAS-RP cable installation - Wks 48-49
  • DFBXH:
    • Ineo cabling - Wk 43
    • Installation of the removed water-cooled cable and RQ6 - Flohe Wk 45
    • New cleaning of the DFBXH, the instrumentation will have to be removed and the DFBX re-validated by AT-MEL afterwards - Wk ?
  • ElQA
    • TP4-A: Inner triplet L1 - as soon as possible since it will be the first time they are performed
    • TP4-A: Inner triplet R8 - not before the DFBX is repaired

General

  • AUG test on October 19
  • Cooling water shut-down - Wks 42-44 - All the cryogeninc activities will be stopped.
  • A latest date for taking a decission about repairing or not the DFBAs has to be set for not jeopardizing the starting date of the flushing

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-10-11 - 09:34

From Pressure Tests to Cooldown - Point 8 - Thursday Oct 11, 2007

Sector 8-1

Going-on Activities

  • SSS Replacement, including:
    • Replacement and interconnection works
    • Leak tests
    • Local ElQA tests

  • Inner Triplets:
    • Left 1: closed, being leak tested
    • Right 8: leak localised in C'. Repair going on. Final test wil be done once the SSS repairs are finished.

  • Electrical non-conformities on instrumentation cables of the power converters will be closed by the end of the week.

  • LHCb compensator tests will last till tomorrow Friday

  • Installation of the instrumentation rods in DFBs Right 8 - COMPLETED

  • Water pump check Point 8 planned for this week

Planned activities

  • PIC-PC interface test - Wk 43
  • Room temperature vacuum and collimator works (wks 43-46), bake-out (wks 47 -48)
  • ATLAS-RP cable installation - Wks 48-49
  • DFBXH:
    • Ineo cabling - Wk 43
    • Installation of the removed water-cooled cable - Flohe Wk ?
    • New cleaning of the DFBXH, the instrumentation will have to be removed and the DFBX re-validated by AT-MEL afterwards - Wk ?
  • MB1158 diode: electrical continuity tests - Monday 15 Oct
  • Flexible water cable reconnection DFBs Left 1 - Week 42

General

  • AUG test on October 19 - Real impact will be described today
  • Cooling water shut-down - Wks 42-44 - All the cryogeninc activities will be stopped. The purging acitivities required to continue with he vacuum tests could be carried out using a temporary solution. Paul will ask Francois to go ahead with it.
  • A latest date for taking a decission about repairing or not the DFBAs has to be set for not jeopardizing the starting date of the flushing

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-10-09 - 11:51

From Pressure Tests to Cooldown - Point 8 - Monday 09.10.2007

Sector 8-1

  • Inner Triplets
    • Right 8: the leak has been located in line C', precisely on a weld in the middle of the flexible. It is currently being replaced by the US collaborators. Once it is corrected and the works in the SSS completed, the inner triplet will be pressurised again for testing. It should be sufficient for this test to have the C' line in the SSS welded, however it is not sure that only one line can be pressurised without affecting the other lines, hence most likely the test will be carried out at the end of the SSS works (wk 47).
    • Left 1: inner triplet closed and being leak-tested.
  • SSS Q17R8
    • The replacement of the SSS started this week and will last till wk 46.
    • Local electrical tests will progress in parallel with the assembly of the SSS.
  • MB1158
    • The diode has been cleaned up. A continuity test will be done before closure beginning of next week. The final validation cannot be done before the TP4-A before the flushing.
  • DFBAs repairs
    • A latest date for taking a decission about repairing or not the DFBAs has to be set for not jeopardizing the starting date of the flushing.
    • Paolo reminds that the time used for reparing the DFBAs will be different for each sector due to the different configurations.
  • ElQA
    • TP4-A is planned for week 47, just before the flushing.
    • Low voltage continuity tests of the dipole lines will be carried ut during the SSS replacement between the DFBAA and C18R8
    • The non-conformities being corrected by INEO this week will be re-checked once they are closed.
  • AUG Test
    • The yearly AUG test for the whole Point 8 is planned for October 19. HCC will study how can we get the maximum profit from it.
  • A4L1
    • Vacuum mechanical works will start on week 43 and will last till week 47.
    • Collimator to be installed week 44.
    • Neg-activation weeks 48 and 49 (overlaps with one week of flushing)
  • LHCb Compensator Magnet Tests
    • The three compensator magnets will be on today and tomorrow. No interference with other activities.
  • Cryogenics
    • Nothing pressurized in the cold circuits.
    • Warm helium line a few mbars above atm pressure
    • Installation of the instrumentation rods in the DFBs of Right 8 will be done tomorrow Wednesday

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-07-05 - 16:38

Powering Tests Sector 7-8 - Thursday 05.07.2007

Microsoft Excel Spreadsheet PlanningW27_050707.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x07x05x16x36/PlanningW27_050707.xls

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-07-05 - 00:13

Powering tests Sector 7-8 _ 04/07/2007

Morning shift activity.

  • RSD2.A78B2 done up to PCS4
  • RQS.A78B2 wrong polarity
  • RQS.R7B1 wrong polarity
  • RQTL10.R7B1 wrong polarity
  • RSD1.A78B1 high noise
  • RSD1.A78B2 card problems
  • RSF2.A78B1 done up to PCS2
  • RSF2.A78B2 tripped during PCS1
  • RSF1.A78B2 done up to PCS3

Performance test done on: RQS.L8B1 RCS.A78B2 RCS.A78B1

24h run stopped at 11:30.

Afternoon shift.

The following circuits have been testes up to PCS4:
  • RQS.A78B2
  • RQS.R7B1
  • RQTL10.R7B1
  • RSF1.A78B2
  • RSF2.A78B1
  • RSF2.A78B2

The following circuits are affected by a problem in a QPS card: RSF1.A78B1 RSD1.A78B1 RSD1.A78B2

RQ6.R7B1/B2 were tested but tripped due to QPS threshold. Reiner needs to increase the threshold there. In agreement with MPP, it has been decided to raise their threshold to 50 mV.

Activity for tomorrow, Thursday 5th.

  • condamnation of all coverters of point 7, to allow ElQA to start their tests there (RQ6, in DFBMH, is on the same electrical depart as other circuits of DFBAN);
  • Reiner will work on the QPS system for RSD1.A78B1/B2 and RSF1.A78B1 tomorrow morning around 9.00 (change of detection card required); as soon as he finishes, those circuits will be powered up to PCS4;
  • as soon as the three circuits above are over, condamnation of all converters of DFBAO;
  • as soon as Reiner finishes the intervention on RQ6 (probably early afternoon), the corresponding converters will be de-condamned and RQ6 powered for PCS4 (THIS HAS TO BE DONE IN ACCORDANCE WITH ElQA FOR WHAT MENTIONED ABOVE ON THE SAME ELECTRICAL DEPART!!!); the converters will be condamned again right after.

  • ElQA activity:
    • matching section to be completed;
    • RB to qualify (after disconnection by Philip Denis of leads in point 7)
    • RQF/D qualification
    • 60 A COC

    • DFBAN ready from the morning
    • DFBAO ready from the afteernoon

  • Installation of passive heating system on one 600A of DFBAN (Vincent Fontanive)

  • DFBXG.3L8 will be cold-tested tomorrow afternoon.

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-07-04 - 16:47

Powering Tests Sector 7-8 - Test Progress Wednesday 04.07.2007 Afternoon

Microsoft Excel Spreadsheet PlanningW27_040707_afternoon.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x07x04x16x26/PlanningW27_040707_afternoon.xls

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2007-07-03 - 21:13

Report from Tuesday 3rd July 2007 - Powering Tests Sector 78

Microsoft Excel Spreadsheet PlanningW27_040707.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x07x03x19x54/PlanningW27_040707.xls

Tests carried out on Tuesday 03.07.2007

  • 600A Correctors
    • PCS completed OK for: RCS.A78B2, RCS.A78B1, RQTD.A78B1, RQTD.A78B2, RSD2.A78B1
    • PCS.3 (100A) completed OK for RSF1.A78B1
    • PLI1 (50A), PLI2 (100A) and PNO (550A) for RQS.L8B1 - Test OK
    • ElQA, PCC and PIC2 completed for RQS.A78B2, RQS.R7B1 and RQTL10.R7B1 - Test OK
    • Stability tests on RQT12.L8B1 completed - Test OK
    • Stability tests on RQTF.A78B1 ongoing

  • Manual Tracking tests for RQF.178, RQD.A78 and RB.A78 completed

  • LSA tracking tests for RQF.178, RQD.A78, RB.A78, RQ4, RQ5 and RQS.L8B1 completed

  • Long Run including the following circuits launched around 19:30 and ongoing.

RCBYHS4.L8B1 RCBYVS4.L8B2 RCBYVS4.L8B1 RCBYV4.L8B1 RCBYH4.L8B2 RCBCH5.L8B1 RCBCV5.L8B2

RCBV23.L8B2 RCBH23.L8B1 RCBH25.L8B1 RCBV27.L8B2 RCBH27.L8B1 RCBV29.L8B2 RCBH31.L8B1 RCBV32.L8B1 RCBH17.L8B1 RCBV17.L8B2 RCBH16.L8B2 RCBV14.L8B1 RCBH13.L8B1 RCBV13.L8B2 RCBH12.L8B2 RCBV11.L8B2 RCBV24.L8B1 RCBH24.L8B2 RCBV25.L8B2 RCBV28.L8B1 RCBH28.L8B2 RCBH29.L8B1 RCBV31.L8B2 RCBH32.L8B2 RCBH18.L8B2 RCBV18.L8B1 RCBV16.L8B1 RCBH15.L8B1 RCBV15.L8B2 RCBH14.L8B2 RCBV12.L8B1 RCBH11.L8B1

RQD.A78 RQF.A78 RB.A78 RD2.L8 RQ4.L8B2 RQ4.L8B1 RQ5.L8B1 RQ5.L8B2

RQS.L8B1

Test Plan for Wednesday 04.07.2007

See attached file

Maria Paz Casas, Antonio Vergara and Blanca Perea for Hardware Commissioning- 03 Jul 2007

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-07-03 - 16:39

ElQA Sector 7-8 - Last Week of Powering Tests

Microsoft Excel Spreadsheet ElQAS78.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x07x03x16x21/ElQAS78.xls

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-07-03 - 08:52

Powering Tests - Sector 7-8 Morning Plan

Microsoft Excel Spreadsheet PlanningW27_030707.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x07x03x08x51/PlanningW27_030707.xls

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-07-02 - 22:12

Powering Tests Sector 7-8 - Report 02.07.2007

Microsoft Excel Spreadsheet PlanningW27_020707.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x07x02x21x18/PlanningW27_020707.xls

Test Report from Monday 02.07.2007

* RQF.A78, RQF.A78 and RB.A78: Stability tests completed.

* RCBCH5.L8B1 stability test completed.

* 60A Closed Orbit Correctors: PNO completed for all the circuits released by ElQA. Tests OK except for two natural quenches in the circuits: RCBH28.L8B2 and RCBH25.L8B1

* 600A Correctors:

  • PCC and PIC2 completed for all the circuits available in DFBAO except RQTD.A78B1 and RQTD.A78B2
  • PCS completed successfully for RQTF.A78B1 and RQTF.A78B2
  • RCS.A78B2 - PCS.2 (50A) PCS.3 (100A) - Test OK
  • RCS.A78B1 - PCS.2 failed due to inversed polarity

Test Plan for Tuesday 03.07.2007

See file attached

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-07-02 - 12:11

Powering Test Sector 78 - 2nd July 2007

Microsoft Excel Spreadsheet PlanningW27_020707_.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x07x02x12x09/PlanningW27_020707_.xls

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-07-02 - 11:04

Cooldown Sector 78

We are now in the last week before the warm-up of sector 78 for the repair and connection of the inner triplet magnets and some consolidation work which is expected to last six weeks.

During the past week, the commissioning of sector 78 has continued on several fronts: the main magnet circuits, the 600 A corrector circuits, the 60A dipole correctors as well as the matching section magnets. This activity was made possible thanks to the very stable operation of the cryogenic system and of the utilities almost throughout the week. Unfortunately starting from Thursday morning two independent faults (return heater, 24V power supply) in the cryogenic system compromised the powering tests at 1.9K for the rest of the week.

After the two quadrupole circuits, which were both powered to 6.5 kA, the dipole circuit was successfully taken to 2 kA on Wednesday (June 27th) morning. The first cycle is shown below. This cycle ends with a discharge of the circuit in the dump resistor.

The in depth analysis of the data collected during these runs is ongoing to validate the correct operation of all the equipment.

The commissioning of the 600A circuit could finally start after the careful tuning of the quench detection system and the successful interlock test with the power converters connected to the magnets. The quench detection system is set to run with very low detection thresholds. The tests on these circuits are aimed at validating the intermagnet splices some of which could have been executed not to the required standards. These tests will continue this week.

Although not at full thrust, the commissioning of the 60 A orbit corrector circuits is progressing. In addition to all this, detailed studies on many circuits were performed to better identify the powering parameters of these circuits.

Studies to better understand the coupling between the Q4 and the D2 magnets did not succeed to reproduce the phenomenon which was initially observed despite all the efforts.

The preparation for the cool down of sector 45 is well advanced: the electrical quality assurance identified and localized a short in the dipole circuit which was very quickly repaired. A mechanical non conformity was also detected during the flushing: it concerned the swap of two lines in the interconnect between the low and high current module of the DFBAH. This was also quickly repaired and the conditioning of the lines continued over the weekend.

The puppets of the Joule-Thomson CV910 were also replaced and the cooldown is expected to start on Wednesday July 4th.

PNG Cooldown_78_Jul02_1.PNG:
Cooldown_78_Jul02_1.PNG

PNG Cooldown_78_Jul02_2.PNG:
Cooldown_78_Jul02_2.PNG

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-06-29 - 20:33

Powering Tests Sector 7-8 _ 29/06/2007

D2-Q4 study in the afternoon

The following cycle has been executed on D2-Q4, to study stability at minimum liquid helium level in DFBMA:

  • with nominal liquil helium level inside the DFBMA, RD2.L8 and RQ4.L8B1/B2 to I_min_op
  • lower liquid helium level to 31%
  • ramp of RD2.L8 to 6kA with 18 A/s ramp rate; once at nominal, ramp of RQ4.L8B1/B2 to 3610 A with 10A/s ramp rate
  • plateau of 30 min, then ramp RQ4.L8B1/B2 to I_min_op (RD2.L8 always at nominal)
  • wait 2 min and ramp RQ4.L8B2 to 3610A at 10 A/s and RQ4.L8B1 to 1900A at 10A/s
  • plateau of 20 min
  • power permit removed on all PCs and SPA initiated withou quench.

Test result: successful!

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-29 - 17:16

Powering Tests Sector 7-8 - 29.06.2007

Microsoft Excel Spreadsheet PlanningW27.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x29x16x22/PlanningW27.xls

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-06-28 - 19:42

Powering test

Priorities and Tests for next week chategorized by subject

  600A Circuit Program  
1 PCS tests for RQTF.A78B1, RQTF.A78B2 Reiner, Bob, Sandor
1 PCC & PIC2 for RSD1.A78B1, RSD2.A78B2, RSF2.A78B1, RSF2.A78B2, RCS.A78B1, RCS.A78B2 Markus, Valerie
2 PCS tests for RSD1.A78B1, RSD1.A78B2, Reiner, Bob, Sandor
2 PCC & PIC2 for RQTD.A78B1, RQTD.A78B1 Markus, Valerie
3 PCS tests for the rest (RSD2.A78B1, RSD2.A78B2, RSF2.A78B1, RSF2.A78B2, RCS.A78B1, RCS.A78B2, RQTD.A78B1, RQTD.A78B2, RSF1.A78B1, RSF1.A78B2) Reiner, Bob, Sandor
4 PCC, PIC2 and PCS tests of RQS.A78.B2, RQS.R7B1, RQTL10.R7B1 Markus, Valerie, Reiner, Bob, Sandor
5 PCS tests of RQ6.R7B1, RQ6.R7B2 Reiner, Bob, Sandor
6 Performance and type tests of RQS.L8B1, RQT12.L8B1, RQTF.A78B1, RCS.A78B1 Valerie, David

  Main magnets  
1 Quadruple quadrupole quench AB/PO, AT/MEL, HCC

  60A orbit correctors  
1 P2N AB/PO, HCC

  Power Converter Studies  
1 RQF.A78 and RB.A78 stability and reproducibility Valerie, David
2 RCBCH5.L8B1 and RCBYH4.L8B2 stability and reproducibility Valerie, David

  Sector studies  
1 Power converter tracking tests (RB, RQF, RQD) Valerie, David
2 LSA tracking tests (RB, RQF, RQD, Q4, Q5, D2 and some 600 A corrector circuits) Mike, David
3 Matching Section: Powering of RD2.L8 and RQ4.L8 with DFBMA at minimum liquid operational level (31%) - RD2.L8 quench detector threshold at 20mV Amalia, Glyn, Andzrej
4 Long duration tests of all available circuits – Minimum 8-hour run (overnight tests?) HCC

  ElQA  
1 RQS.A78.B2, RQS.R7B1, RQTL10.R7B1 Davide
2 All the circuits in the sector via DFBs 2 days Davide
3 3006 additional diagnostic 1.5 days Davide
4 Q5 correctors (4) in DFBMC 1 day Davide
5 Installation of monitoring system for warm-up Davide
6 Boiloff test on DFBMA on the two chimneys 120A circuits Davide
7 Remaining two third of 60A orbit correctors at cold Davide

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-27 - 19:09

Powering Tests Sector 7-8 - Wednesday 27.06.2007

PDF Progress_PoweringS78_070628.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x27x18x48/Progress_PoweringS78_070628.pdf

Test Report from Wednesday 27.06.2007

  • RB.A78
    • PLI1.4 - FPA from PIC - Test OK
    • PLI1.5 - Quench heater discharge on A28.R7 - Test OK
    • The bus-bar detection threshold is increased from 1V to 2V in order to ease the ramp up
    • Circuit successfully ramped to 2000A!!
      • PLI2.1 - Current cycle - Test OK
      • PLI2.2 - Energy extraction discharge (quench to a current lead) - Test OK
      • PLI2.2b - SPA from PIC - Test OK
      • PLI2.3 - Quench heater discharge on LBBRA.13L8 (magnet 3002) and LBBRA.13R7 (actually we wanted to quench LBARA.13R7) - Test OK

  • RQF.A78 and RQD.A78 ramped up together to 6500A. Quench on 15L8 - Test OK

  • Simmultaneous powering of RQF.A78, RQD.A78 and RB.A78 - Strated @ 21:00 - Stopped @ 22:00 - Test OK

  • 600A Correctors - 9 new circuits released by QPS (see pdf file attached)
    • PIC2 done on 9 circuits
    • PCS.2/.3/.4 on RQS.L8B1 and RCO.A78B1

  • 60A Closed Orbit Correctors
    • P2N continued for 2 circuits

  • Matching Section
    • Reproduction of the quench event of May 31st - Magnet did not quench.
    • Last commissioning step of RQ5.L8 via the sequencer - Test OK

  • Stability test on RQ4.L8B1/B2

Test Plan for Thursday 28.06.2007 - Noon Update

Tests will resume once ElQA and connection of the current leads on 8 circuits of the DFBAO (see file attached) are completed.

  • PCS tests for all the circuits released by PIC2

  • In parallel: PIC2 for the eight circuits that have not gone through it yet (whenever possible)

  • RQF.A78, RB.A78 - Performance tests

  • Local tests of AB/PO in one corrector circuit of DFBMC (the test implies disconnection-connection of active parts in the converter side)

  • Matching Section: Powering of RD2.L8 and RQ4.L8 with DFBMA at minimum liquid operational level (31%) - RD2.L8 quench detector level at 20mV

  • P2N continues for 60 closed orbit correctors

  • Quadruple quadrupole quench post-poned for tomorrow before lunch.

Outstanding activities

  • Performance tests of two 600A circuits once released, 4 hours per circuit are required.
  • Tracking tests of RB/QF/QD using PO function generation. Provisionally scheduled for Friday 29/07, from 14:00
  • Tracking tests of all available circuits (RB/QF/QD/Q4/Q5/D2/600A) using LSA application. Provisionally scheduled for Monday 2/7/07, from 08:00
  • Long duration tests of all available electrical circuits, minimum of 8H test
  • ElQA for circuits not powered in this run
  • AUG - Scheduled for Monday 9/7/07

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-26 - 20:06

Powering Tests Sector 7-8 - Tuesday 26.06.2007

PDF Progress_PoweringS78_070627.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x26x19x45/Progress_PoweringS78_070627.pdf

Test Report from Tuesday 26.06.2007

  • RB.A78
    • PIC2 - Test completed OK
    • During the whole day, the noise read by the bus-bar quench detector of the circuit tripped several times. The source of the noise (period of about 20 minutes) has not been identified yet.
    • PLI1 (760A)
      • PLI1.1 - Current cycle - Test OK
      • PLI1.2 - Energy extraction discharge (quench to a current lead) - Test OK
      • PLI1.3 - Powering Failure from PC - Test OK

  • RQ5.L8: Commissioned up to PLI3 using the sequencer - Tests OK

  • 60A closed orbit correctors - P2N completed for one corrector using new automated procedure

  • RQF.A78 - Type test completed - Test OK

  • 600A correctors
    • RCO.A78 and RQT12.L8B1 - PCS completed - Test OK

Test Plan for Wednesday 27.06.2007

  • Show goes from 7:00 to 22:00

  • RB.A78
    • PLI1.4 - FPA from PIC
    • PLI1.5 - Quench heater discharge on A28.R7
    • PLI2 (2000A)
      • PLI2.1 - Current cycle
      • PLI2.2 - Energy extraction discharge (quench to a current lead)
      • PLI2.2b - SPA from PIC
      • PLI2.3 - Quench heater discharge on LBBRA.13L8 (magnet 3002) and LBARA.13R7 (magnet 1054)

  • RQF.A78 and RQD.A78 (once RB.A78 tests are completed)
      • RQF and RQD ramped up together to 6500A. Quench on 15L8
      • RQF and RQD ramped up together to 6500A. Quadruple quench 20L8, 19L8, 19R7 and 22R7 .

  • Simmultaneous powering of RQF.A78, RQD.A78 and RB.A78.

  • 600A Correctors - More circuits will be released in the afternoon by QPS for PCS tests

  • 60A Closed Orbit Correctors
    • P2N will continue
    • Type tests for some converters in the evening after quenches in the main circuits

  • Matching Section
    • Reproduction of the quench event of May 31st
    • Last commissioning step of RQ5.L8 via the sequencer

  • AB/PO (David) will prepare a list of the activities to be carried out in the conveters already commissioned before the warm-up

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-25 - 19:34

Powering Tests Sector 7-8 - Report Monday 25.06.2007

Test Report of Monday 25.06.2007

  • RB.A78
    • Provisional correction of the detection threshold of the bus-bar quench detector.
    • TF measurement - Test OK after some corrections done in one QPS crate
    • RB converter connected to the circuit
    • Circuit ramped up to 350A with the bus-bar quench detection threshold increased. Loop properly closed!
    • Some modifications needed in the QPS bus-bar quench detector - Circuit ramped up again to 350A with QPS in operational conditions!!
    • PCC Circuit Config - TEST OK!

  • RQF.A78 (6500A)
    • PLI3.2 - Energy extraction discharge - Test OK
    • PLI3.3 - Quench heater discharge on 12L8 SSS - Test OK

  • 600A Correctors
    • PCS tests on the circuits: RQT12.L8B1, RQS.L8B1, RCO.A78B1 - Going on

  • DFBMA refilled and cooled down in a couple of hours

  • 60A Closed Orbit Correctors: PCC in 16 converters

Test Plan for Tuesday 26.06.2007

  • RB.A78
    • PIC2
    • PLI1 (760A)
      • PLI1.1 - Current cycle
      • PLI1.2 - Energy extraction discharge (quench to a current lead)
      • PLI1.3 - Powering Failure from PC
      • PLI1.4 - FPA from PIC
      • PLI1.5 - Quench heater discharge on LBBRG.10L8 (magnet 1030)
    • PLI2 (2000A)
      • PLI2.1 - Current cycle
      • PLI2.2 - Energy extraction discharge (quench to a current lead)
      • PLI2.2b - SPA from PIC
      • PLI2.3 - Quench heater discharge on LBBRA.13L8 (magnet 3002) and LBARA.13R7 (magnet 1054)

  • Matching Section: Powered to the same conditions of the triple quench on 31/06 once the quench detection threshold has been increased to operational values.

  • 60A closed orbit correctors: PCC and P2N

  • RQF.A78 or RQD.A78
    • Type tests
    • Quench campaign:
      • RQF and RQD ramped up together to 6500A. Quench TWO quadrupoles in different cryogenic cells simmultaneously.
      • RQF and RQD ramped up together to 6500A. Quench FOUR quadrupoles in different cryogenic cells simmultaneously.

  • 600A correctors
    • Study and debugging of the quench detectors going on.

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-24 - 14:47

Powering Tests Sector 7-8 - Report 22.06.2007

Test Report of Friday 22.06.2007

  • RB.A78
    • PIC1 - Completed OK after a the solution of an adjustment problem in the current loop controller of the QPS (few mA too low).
    • Connection of the 60A conveter to the RB circuit in the UA83 side. The TF measurements could not be completed because the converter noise made the bus-bar quench detector trip everytime it was switched on. The detection threshold will be increased on Monday.
    • The time out preventing from closing the switches too early after a switch opening has been lowered in order to speed the tests up.

  • RQF.A78
    • PLI1 (760A) completed - Test OK
    • PLI2 (2000A) completed - Test OK
    • PLI3.1 (current cycle up to 6500A) completed - Test OK

  • 60 Closed Orbit Correctors
    • PCC carried out in one circuit.
    • All the circuits released by ElQA have been connected to the power converter.

  • AB/PO current stability measurements in RD2.L8 - Completed

  • Empty of DFBMA and Q4/D2 started in the evening

Test Plan for Monday 25.06.2007

  • RB.A78
    • Provisional correction of the detection threshold of the bus-bar quench detector in order to allow TF measurement with the 60A converter.
    • TF measurement
    • Desconnection of the 60A converter and final connection of the RB converter to the circuit; PCC tests
    • Set detection threshold back to operation value
    • PIC2
    • PLI1 (760A)
      • PLI1.1 - Current cycle
      • PLI1.2 - Energy extraction discharge (quench to a current lead)
      • PLI1.3 - Powering Failure from PC
      • PLI1.4 - FPA from PIC
      • PLI1.5 - Quench heater discharge on a selected magnet
    • PLI2 (2000A)
      • PLI2.1 - Current cycle
      • PLI2.2 - Energy extraction discharge (quench to a current lead)
      • PLI2.2b - SPA from PIC
      • PLI2.3 - Quench heater discharge on a selected magnet

  • RQF.A78 (6500A)
    • PLI3.2 - Energy extraction discharge
    • PLI3.3 - Quench heater discharge on 12L8 SSS

  • 600A Correctors
    • Software upgrade by QPS
    • PCS tests on the circuits: RQT12.L8B1, RQS.L8B1, RCO.A78B1

  • DFBMA will me refilled in the afternoon

  • 60A Closed Orbit Correctors: PCC and P2N whenever possible

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-21 - 20:12

Powering Tests Sector 7-8 - Report 21.06.2007

Test Report of Thursday 21.06.2007

  • RQD.A78
    • PLI3.2 Discharge Request - Test OK
    • PLI3.3 Quench heater discharge in 12R7 - Test OK - Cryo_Maintain was recovered after 10 minutes and Cryo_Start in about 30 minutes.

  • RQF.A78
    • PIC2 completed - Test OK
    • PLI1.1 - Current cycle at 760A completed - Test OK
    • PLI1.3 - PC Failure at 760A completed - Test OK
    • Communication problems with the energy extraction controller prevented from closing the switches remotely. Tests stopped.

  • RB.A78
    • QPS IST completed - 2 DQHDS had to be replaced
    • A communication problem came up due to a corrupted code line maybe caused during the downloading of the last software version.
    • Current leads connected to DC cables in DFBAO and DFBAN
    • Last corrections being done by QPS experts

  • Matching Section
    • Heat runs continue during the whole week with RD2 and RQ5
    • Type tests of the converters by AB/PO completed.

Planning for Friday 22.06.2007

  • RB.A78
    • PIC1
    • Connection of the 60A conveter and TF measurement
    • Connection of the RB converter to the circuit+ PCC
    • PIC2

  • RQF.A78 (in parallel with RB whenever possible)
    • Complete PLI1 (760A)
      • Energy extraction discharge
      • FPA
      • Discharge request from PC
      • Heater firing
    • PLI2 (2000A)
    • PLI3 (6500A)

  • 60 Closed Orbit Correctors
    • PCC and P2N on selected circuits

  • AB/PO current stability measurements in RD2.L8

  • Empty of DFBMA and Q4/D2 cryostats over the weekend

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-20 - 20:20

Powering Tests Sector 7-8 - 21.06.2007

Test Report of Wednesday 20.6.2007

  • RQF.A78
    • PIC2 completed up to Fast Power Abort from PIC - Test OK

  • RQD.A78
    • PIC2 tests completed - OK
    • PLI1 tests at 760A completed - OK
    • PLI2 tests at 2000A completed - OK
    • PLI3 started. RQD.A78 circuit ramped up successfully to 6400A, kept there for more than one hour. - OK
    • Champagne and pictures
    • PLI3.1 test completed - Magnet ramped up to 6500A and down to 350A successfully!!

  • Matching section
    • Type tests on different circuits performed by AB/PO

Planning for Thursday 21.06.2007

  • RQD.A78
    • PLI3.2 Discharge request
    • PLI3.3 Quench heater discharge

  • RB.A78
    • After tests completed in RQD.A78 (in parallel):
      • QPS IST tests
      • Connection of the current leads in DFBAO and DFBAN
    • PIC1
    • Connection of the 60A converter - Transfer function measurement
    • Cabling repair 600A - RCS, RCD circuits.
    • Wait for DFBAN release from ACR

  • 600 Circuits - Further measurements by QPS - Depending on Reiner's progress in RB, which has full pripority.

  • Matching section
    • Type tests AB/PO continue
    • ACR checks on the DFBMA current leads with D2 at nominal current

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-19 - 19:55

Powering Tests Sector 7-8 - 19.07.2007

Test Report from Tuesday 19.07.2007

  • RQD.A78:
    • Transfer function measurements with 60A converter completed OK
    • PCC Circuit Config tests completed
    • PIC2 - All steps completed OK except:
      • Circuit discharge from power converter: switches were not rearmed after PM data generation - Test to be repeated
      • Discharge request from PIC - to be done

  • RQF.A78:
    • Transfer function measurements with 60A converter completed OK
    • PCC Circuit Config tests completed
    • PIC2 - All steps completed OK up to Powering Failure test (included)

  • RQ4.L8 and RD2.L8:
    • Tests 1 and 2 proposed by MPP completed but without provoking an artificial quench at the end of the second. Detection threshold still at 20mV. Still no quench. Further research to be done.

  • RCO.A78B1 powered below 20A for QPS analysis. The circuit is still blocked till the new sofware is uploaded.

Test Plan for Wednesday 20.07.2007

  • Resolution of cable signal/interlock swap F/D

  • RQD.A78 (from the CCC):
    • Last two steps of PIC2
    • PLI1 (750A)
    • PLI2 (2000A)
    • PLI3 (6500A)

  • RQF.A78 (from the CCC):
    • Complete PIC2
    • First steps of PLI1,2,3 if possible

  • RB.A78
    • IST Energy extraction tests, PIC1, connection of the leads

  • Matching section:
    • Type analysis of the power converters
    • Analyse boil-off data and eventually repeat a more complete boil-off
    • Waiting for new proposals from MPP
    • 120A converter on Q5 type tests

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-19 - 15:21

Powering RQD.A78 and RQF.A78

  • PIC2 - I = I_MIN_OP = 350A

  • PLI1 - I = I_INJECTION = 760A
    • PLI1.1 - Current cycle: 350A to 760A at 10 A/s
    • PLI1.2 - Energy extraction discharge at 760A (bus-bar or current lead quench)
    • PLI1.3 - Powering failure from power converter
    • PLI1.4 - FPA from PIC
    • PLI1.5 - Quench heater firing
      • SSS049 15L8

  • PLI2 - I = I_INTERM_1 = 2000A
    • PLI2.1 - Current cycle: 350A to 2000A at 10 A/s
    • PLI2.2 - Energy extraction discharge at 2000A (bus-bar or current lead quench)
    • PLI2.3 - Slow Power Abort from PIC
    • PLI2.4 - Quench heater firing
      • SSS016 24R7

  • PLI3 - I = I_INTERM_2 = 6500A
    • PLI3.1a - Currrent cycle: 350A to 6500A at 10 A/s
    • PLI3.2 - Energy extraction discharge at 6500A (bus-bar or current lead quench)
    • PLI3.3 - Quench heater firing

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-06-18 - 18:15

Update on Powering tests sector 78 - Afternoon 18 June 2007

Progress during Monday 18.06.2007 from the 8:30 meeting till the 16:30 meeting

  • RQF.A78 and RQD.A78
    • Wait for green light from QPS - OK
    • PIC1 for both circuits: OK but several points to note:
      • PM problem noted: the data, although complete, does not display correctly in the PM viewer. Sandrine will look at data manually and Adriaan will correct the problem
      • after a discharge request from the PC, the PC fault and discharge request signals arrive out of sequence. According to Valerie this is understood (due to cable delays)
      • QPS_OK does not take into account the switches status allowing powering of the magnets with the switches open
    • Connection of the DC cables to the current leads has not happened yet. Will be done tomorrow at 7:30

  • RB.A78
    • QPS IST. All heaters charged (2 out of 700 needed intervention) Discharging started at 16:45 (will need 3-4hours to complete)

  • Matching Section
    • MPP presented the next tests to investigate the triple-quench event. It includes various ramping schemes, forced quench of Q4 and a boiloff test of DFBMA, to be completed during the next few days. Tonight: DFBMA boiloff and re-fill.

  • 600A correctors in DFBAO
    • ACR/MEL tests on the damaged cryo-instrumentation cards (RSF2.A78B2, RCS.A78B1) - ONGOING
    • If possible, solution of problem in RCO QPS communication -DONE
    • QPS Software upgrade - Will not done before tomorrow

  • Tests of the 60A CRYO_START/CRYO_MANTAIN signals - OK although for the 60A circuits all tests will have to be done in conjuction with the cryo operators with a manual CRYO_START/MAINTAIN

A.O.B.: Francois reminded the meeting that it is preferable that the cryo EIC is informed (as well as the cryo operators) of any pending tests.

Programme for tomorrow Tuesday 19 June 2007

  • RQF.A78 and RQD.A78
    • Connection of DC cables to the current leads
    • connection of a 60A converter to the RQD.A78 circuit from the UA83
    • PCC tests for RQD.A78 (MAXIMUM CURRENT: 350A) - Experts from AB/PO, PIC, QPS and Energy Extraction required
    • PIC2 tests for RQD.A78 in parallel with PCC for RQF.A78 (MAXIMUM CURRENT FOR BOTH TESTS: 350A)

  • RB.A78 QPS IST tests will continue during the whole day

  • Matching Section
    • Test 1 proposed by MPP (see slides attached)
    • Test 2 proposed by MPP, but:
      • If detection threshold can be increased up to 50 mV: full test
      • If detection threshold cannot be modified yet: full cycle without provoked quench at the end

  • 600A correctors in DFBAO:
    • ACR/MEL tests on the electronic cards will go on till RQD.A78 is released by PIC2
    • Software upgrade of the quench detectors

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-15 - 17:25

Powering Tests Sector 78 - Evening 15.06.2007

Planning for Monday 18.06.2007

  • RQF.A78 and RQD.A78
    • Wait for green light from QPS
    • PIC1 for both circuits
    • Connection of the DC cables to the current leads
    • Connection of a 60A converter to the DC cables in the UA side
    • PCC tests for both circuits

  • RB.78
    • QPS IST. Discharge of the quench heaters, quench loop tests, energy extraction type tests

  • Matching Section
    • Further tests for investigating the triple-quench event to be discussed in MPP meeting on Monday

  • 600A correctors in DFBAO
    • ACR/MEL tests on the damaged cryo-instrumentation cards (RSF2.A78B2, RCS.A78B1)
    • If possible, solution of problem in RCO QPS communication
    • QPS Software upgrade

  • Tests of the 60A CRYO_START/CRYO_MANTAIN signals

Test Report of Friday 15.06.2007

  • Matching Section
    • In the morning RD2.L8 and RQ4.L8 and the 7 correctors ramped to nominal.
    • While starting ramp up, RQ5.L8 quench detector tripped due to the very high acceleration in the converter.
    • After some corrections, RQ5.L8 was also ramped to nominal current. The softaware of the quench detector of this circuit still needs to be upgraded
    • During the afternoon Squeezing tests from LSA were performed

  • 600A correctors in DFBAO
    • Eight out of the nine circuits released by ElQA and ACR went through the PCC tests OK. QPS gathered the data needed to update the detector and released the circuits. PIC2 performed in the afternoon in all of them. Test OK.
    • RCO has a problem with the QPS communication

  • 60A type tests completed OK.

  • RQF.A78 and RQD.A78
    • Switch opening failure issue already solved. Small communication problem to be fixed (world-FIP connection). The two circuits should be released for PCC tests on Monday morning.

  • RB.A78
    • Quench heaters are now charged.

  • Boil-off tests on the DFBAN successfully carried out on Saturday.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-06-13 - 20:48

D2 Quench

Bitmap quench_d2_q4.bmp:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x13x20x32/quench_d2_q4.bmp

Bitmap D2-Q4.bmp:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x13x20x32/D2-Q4.bmp

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-06-13 - 15:55

Powering of sector 78 - Report of 13.06.2007

Report of Wednesday 13.06.2007

  • ElQA
    • Tests in MBA during all the day.
  • RQ5.L8:
    • Calibration test performed by AB/PO. Test ongoing.
  • RD2.L8:
    • Ramped to 200A. Quench heater fired. Post-mortem data acquired properly. Test OK
    • Ramped to 1kA. Quench heater fired. Post-mortem data acquired properly. Test OK
  • RQF and RQD
    • Quench heater charge.
  • RD2.L8
    • Test 1 on the ML8 performed. D2 quenched when R4.L8B1 was at 2.2kA and R4.L8B2 was around 3kA ramping to 3.61kA. Test NOT OK.

Planning for Thursday 14.06.2007

AM:

  • ElQA on MB with instrumentation + QF & QD
  • RQ6.R7B1 and RQ6.R7B2: Repeat test
  • QPS IST RQF.A78 and RQD.A78: discharge of the heaters
  • RD2.L8 to nominal
  • RQ5.L8 and their correctors calibration by PO (UA)
PM:
  • 600 A circuits of L8: PCC (<20A) for all
  • 600 A circuits of L8: PIC2 for all
  • 600 A circuits of L8:PCS.1 for all
  • List of the 600A circuits released by AT/MEL and AT/ACR
    • RQTL12.L8B1
    • RQTF.A78B1
    • RQTF.A78B2
    • RSD1.A78B2
    • RQS.L8B1
    • RSD2.A78B1
    • RSF1.A78B1
    • RSF1.A78B2
    • RCO.A78B1
  • AB/PO Type tests on 60A converter

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-06-12 - 18:25

Powering of sector 78 - Report of 12.06.2007

Test Report of Tuesday 12.06.2007

  • ElQA
    • Tests in MBA line will continue tomorrow till noon.

  • RD2.L8:
    • Ramped to 200A. Quench heater fired. Post-mortem data acquired properly. Test OK
    • Ramped to 1kA. Quench heater fired. Post-mortem data does not arrive properly to the viewer. Test NOT OK
    • Second ramp to 1kA. Quench heater fired. Post-mortem data created is corrupted: the data is an exact copy of the one created in the previous with no new data. The QPS controller needs to be restarted. Test NOT OK

  • 600A QPS Controllers:
    • RQ6.R7B1 and RQ6.R7B2 are now ready for powering. Next step: PCS.4 (plateau at 200A)
    • The available 600A correctors in the DFBAO are released by AT/MEL for powering.

  • Sequencer Test 600A sequence - Completed using RQTL12.L8B2

  • RQ6.R7:
    • Different ramp rate applied (up to nominal ramp rate of 1.3 A/s). Current cycles between +10A -10A.
    • Noise level: 20 mV
    • Circuit ramped at 5 A/s: trip at 20A. Test NOT OK.

Planning for Wednesday 13.06.2007

  • AM
    • ELQA Test of MBA line will continue till 14h00.
    • RQ6.R7B1 and RQ6.R7B2: Repair of the cable

  • PM (14:00 in CCC)
    • RQ6.R7B1 and RQ6.R7B2: Repeat test
    • 600 A circuits of L8: PCC (<20A) for all
    • 600 A circuits of L8: PIC2 for all
    • 600 A circuits of L8:PCS.1 for all
    • QPS IST RQF.A78 and RQD.A78
    • Quench heater charge/discharge RQF and RQD circuits
    • RD2.L8, RQ4.L8 and RQ5.L8 and their correctors: Powering possible once Reiner has solved the problem with the PM data generation
    • AB/PO Type tests on 60A converter
    • In the late afternoon: ElQA on the 600A circuits of DFBAN.

  • List of the 600A circuits released by AT/MEL and AT/ACR

    • RQTL12.L8B1
    • RQTF.A78B1
    • RQTF.A78B2
    • RSD1.A78B1
    • RQS.L8B1
    • RSD2.A78B1
    • RSD2.A78B2
    • RSF1.A78B1
    • RSF1.A78B2
    • RCO.A78B1

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-08 - 18:27

Powering Tests Sector 78 - Report 08/06/2007

Test Report of Friday 08/06/2007

  • RQ6.R7B1 and RQ6.R7B2
    • PCS.1: Sinus excitation at 1 and 10 mHz OK, the detector tripped for 200 mHz and could not be tested up to 1 Hz as expected. Hoewever, the test is considered as OK for the QPS experts since they gather enough information about the circuit and the dectector performance - TEST OK
    • PCS.2 (Extraction at 50 A): Detector tripped for ramp rates of 0.5, 0.4 and 0.3 A/s, it worked OK at 0.25 A/s. There was a problem in the transmission of the Post Mortem data, which did not reach the server the first time. The problem was fixed and no data was lost. The issue has not been understood yet - TEST OK
    • PCS.3 (Extraction at 100 A): Test OK
    • PCS.4 (Plateu at 200A): The detector tripped while the current was reaching the plateau due to the very well known issue with the d2I/dt2. Energy extraction worked ok - TEST NOT OK.

  • 80-120A Correctors of Q4-Q5:
    • PNO.4 (Slow Power Abort from PIC): Battery of seven circuits test together - Test OK
    • PNO.2 (Full current bi-polar cycle) repeated after the meeting - Test OK

  • RQ5.L8:
    • Squeezing Tests: Done after the meeting - Test OK

  • ElQA:
    • Scope installed and system callibrated using artificial 100 mV spikes.
    • The measurements on the lead heaters insulation gave correct results.
    • Reiner will provide a way to measure the insulation between the quench heaters and the coils.

Planning for Monday 11/06/2007

  • RD2.L8, RQ4.L8. RQ5.L8:
    • Special test explained on Wednesday will be applied as soon as AT/MEL has completed new instrumentation in the RD2.L8 circuit.

  • 600A Correctors in the DFBAO:
    • Reiner will try to release some circuits for PCS tests on Tuesday. Circuits released by ElQA and connected to the leads are:

RQT12.L8B1 RQTD.A78.B1 RQTD.A78.B2 RQTF.A78.B1 RQTF.A78.B2 RSD1.A78.B1 RSD1.A78.B2 RQS.L8.B1 RSD2.A78.B1 RSD2.A78.B2 RSF1.A78.B1 RSF1.A78.B2 RSF2.A78.B1 RCD.A78.B1 RCS.A78.B1 RCS.A78.B2

  • ElQA:
    • Tests will continue in MBA line.

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-08 - 15:45

Powering Tests Sector 78 - Morning Meeting 07/06/07

PDF RAT_20070607.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x08x15x44/RAT_20070607.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-08 - 10:03

Powering Tests Sector 78 - Morning Meeting 08/06/2007

PDF RAT_20070608.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x08x10x02/RAT_20070608.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-07 - 19:18

Powering Tests Sector 78 - Report 07/07/2007

Test Report of Thursday 07/06/2007

  • RQ6.R7:
    • PIC2 tests were blocked due to a damaged electronic card in the QPS controller which affected the creation of Post-Mortem data and the proper generation of interlock signals. The problem is understood now and the card changed. PIC2 will be completed today after the evening meeting.

  • RD2.L8:
    • The circuit was powered and the heater discharged at 200 A in order to check the new instrumentation implemented by AT/MEL. The system still needs some hardware work. The protection system is working properly. RD2.L8 stays blocked till Monday since priority is given to RQ6.
    • Antonio Perin presented the non-conformity in the internal cabling of the RD2 and RQ4 circuits in the DFBMA.L2 and DFBMA.L8 which may be the cause of the triple quench event of last week in the matching section.

  • RQ5.L8 and RQ4.L8
    • Liberated now for further powering after the Powering Subsector OFF functionality has been disabled in the ML8 powering subsector.
    • SPA from PIC in RQ4.L8 at nominal current: Test OK
    • SPA from PIC in RQ5.L8 at nominal current: Test OK

  • 80-120A correctors in Q4 and Q5:
    • Ramped all together after the meeting at nominal current and kept there for a couple of hours.

  • ElQA
    • MBB line to be retested with all the instrumentation connected.
    • MBA line: the test with the cryogening system and one QPS rack disconnected could not start today since the boiling test in the DFBAN is taking longer than expected. If the test fails, it will be repeated with 50% of the QPS racks in the MBB line (38 units) disconnected. If the test is still failing, all the QPS racks will be disconnected.

Planning for Friday 08/06/2007

  • RQ6.R7B1 and RQ6.R7B2
    • PCS.1-PCS.4 tests (Sinus at +/-10A, energy discharge at 50, 100 and 200A) from the CCC
    • Upgrade of the controller software after the test.

  • RQ4.L8 and RQ5.L8
    • Squeezing tests if possible.

  • 80-120A correctors Q4 and Q5:
    • PNO.4 test in battery for the seven circuits available.

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-06-06 - 17:50

Powreing Test Sector 7-8 - Report 06/06/07

Test Report Wednesday 06/06/2007

  • RQ4.L8
    • Manual Squeezing test (PSQ). Test OK. (The test will be considered as fully passed only once automatic function is used)
    • When Squeezing was finished, the magnet was ramped to nominal current (3610 A) and kept there for "endurance test"; after one hour, the B2 converter tripped and a FPA was initiated: analysis of the PM signals was indicating an external interlock alarm - red emergency button pressed. Nobody was in the area at that moment.
    • The two converters were ramped back to nominal and kept there; after three hours, the same trip as before was experienced on B2 and the FPA initiated. The circuit was blocked for powering.
    • Further investigations showed that the two trips were provoked by a mispositioned pin in the connector on the emergency red button. Already fixed by PO.

  • RQ6.R7B1 and RQ6.R7B2
    • PCC completed by PO and QPS in the tunnel of point 7.
    • PIC2 CRYO-OK passed on RQ6.R7B1, RQ6.R7B2, RCBH6.R7B2 and RCBV6.R7B1.
    • Remaining steps of PIC2 passed on RCBH6.R7B2. Not possible to continue with the others due to a communication problem with QPS detectors; a manual reset in UA83 is needed.

Planning for Thursday 07/06/2007

  • RQ6.R7B1 and RQ6.R7B2
    • Complete PIC2 on the two 600 A circuits and the corrector RCBV6.R7B1
    • PCS.1, PCS.2, PCS.3, PCS.4 (Ramp rate 0.2-0.5 A/s)

  • 600 A correctors
    • Cable connection on the circuits released by ElQA (after verification by PO of the condemnation of those converters)

  • 80 A correctors of Q4 and Q5: "Endurance test" after Reiner's intervention

  • ElQA tests:
    • Further investigations on RB circuit/instrumentation

  • ML8: once D2 is available, special cycle with D2, Q4 and Q5.

  • Powering of RQ4.L8 and RQ5.L8 to validate the FGC software update

  • Boil-off tests on DFBAN in the morning.

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-06 - 14:59

Powering RD2.L8, RQ4.L8, RQ5.L8 Proposal after Triple Quench

PowerPoint D2Q4Q5tests_070606.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x06x14x58/D2Q4Q5tests_070606.ppt

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-05 - 18:36

Powering Tests Sector 78 - Morning Meeting 05/06/2007

PDF RAT_20070605.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x05x18x34/RAT_20070605.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-06-05 - 18:33

Powering Test Sector 78 - Evening Report 05/06/2007

Test Report Tuesday 05/06/2007

  • RQ4.L8
    • Special tests carried out by AB/PO from the tunnel in the converter in order to further study the fault in the FWD seen yesterday during the Fast Power Abort test (PNO.9)
    • While trying to ramp the converter up during the decay after a provoked fast power abort, the quench detector tripped and the quench heaters were fired. RQ5.L8 was going through the squeezing tests at the moment of the Q4 sporious quench, hence it performed a Fast Power Abort and its quench detector did not trip although it was at 3 kA.
    • Fast Power Abort from PIC Test repeated after diode was repaired (PNO.9). Test OK.

  • RQ5.L8
    • Manual Squeezing tests (PSQ). Test OK. It will not be considered as fully passed till it is done with the automatic function.

  • RQ6.R7B1 and RQ6.R7B2
    • PCS test up to +10 A OK. The quench detector performed OK at this current level. There is a problem in the supervision that is being fixed. The transfer functions measured by AB/PO match very well the ones measured during ElQA tests.
    • Due to a missunderstanding with MPP and HCC, the circuit was powered up to 60A in order to check the crowbar. At 53 A, as expected, the QPS provoked a FPA that opened the switches. The procedure has to be reviewed since, so far, powering above 10 A is not allowed before carrying out PIC2. HCC will find an agreement by tomorrow between Mr Circuit, AB/PO and MPP.

  • ElQA RB.A78
    • At 600 V, with all the DFB instrumentation except the lead heaters disconnected, the resistance of the leads was 346 and 260 Ohms for 5 minutes which is a very promising result. Tests will continue tomorrow when 1.9 K are expected.

Planning for Tomorrow 06/06/2007

  • RQ4.L8
    • Powering at nominal plateau for some hours
    • If functions are available: Squeezing test

  • RQ5.L8
    • After modification in the quench detector: Fast Power Abort at nominal current (PNO.9)
    • If functions are available: Squeezing test

  • RD2.L8
    • Once new instrumentation is operational: Power to nominal current step by step.

  • ElQA
    • Tests up to 1.9 V in RB.A78

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-06-04 - 18:44

Powering Tests Sector 78 - Afternoon Meeting 04/06/2007

Test Report Monday 04/06/2007

  • RQ5.L8: Powered up to nominal. During the first trial, the quench detector triggered at 27A. The magnet was powered at its nominal current for 4 hours. Tests successful. MPP is asked to give permission to increase the detection threshold to 100 mV (nominal) to carry out the Fast Power Abort test (PNO.9)

  • RQ4.L8:
    • Powered up to 250 A while RQ5.L8 was at nominal in order to check the integrity of the circuit. Test OK.
    • Once RQ5.L8 test was completed and with the agreement of MPP, RQ4.L8 was powered at different current levels up to its nominal current (250, 600, 2000 and 3610 A) and kept there for 20 minutes. Test OK.
    • Squeezing tests (PSQ). Test OK.
    • Fast Power Abort from PIC at nominal current (PNO.9). The circuit did not quench, however, AB/PO considers the test as failed due to the FWD circuit creating a fault. Under investigation. Test NOT OK.

  • RQ6.R7B1 and RQ6.R7B2
    • PCC Circuit Config tests going on. Some correction in the quench detectors have to be carried out. The noise level in the detector looks promising. Tests will continue tomorrow Tuesday.

  • RD2.R8
    • Amalia has analysed previous quenches in this magnet. None of them showed any effect on the level and temperatures in the DFBMA, hence they can conclude that the quench ocurred in the DFBMA but they cannot confirm its cause yet.
    • AB/CO has completed the interlock and circuit synoptics modifications needed for the new instrumentation in D2. Reiner will get the hardwire ready as soon as he finishes the correction in Q6.

  • Boil-off Tests
    • DFBAO Low Current Module: test completed, further studies need to be done in the He level gauges
    • DFBAO High Current Module: Test going out. They should be completed this evening. Promising results so far.
    • DFBAN being refilled, the calibration data is being checked.

Planning for Tuesday (05/06/2007)

  • RQ5.L8:
    • Squeezing Tests (PSQ)

  • RQ4.L8:
    • Run at nominal current for some hours
    • Repeat PNO.9 (FPA) if required by AB/PO

  • RQ4.L8 & RQ5.L8 & Correctors
    • Powering at nominal current together (?)

  • ElQA RB.A78
    • Tests will be restarted tomorrow Tuesday. First priority activity.

  • 60A Converter Type test
    • If possible in coordination with the ElQA.

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-06-04 - 11:07

Powering Tests Sector 78 - Minutes 04/06/2007

PDF RAT_20070604_2.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x04x11x06/RAT_20070604_2.pdf

Pptm D2_6000A.pptm:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x04x11x06/D2_6000A.pptm

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-06-04 - 08:54

Cooldown Sector 78

The commissioning programme on Q4, Q5, D2 and the associated correctors continued during this week. While the Q4, Q5 and D2 commissioning was completed, the commissioning of the 120 A correctors associated to these magnets could not be completed: some additional tuning of the software in the FGC remains to be done.

Yesterday (Thursday May30th) Q4, Q5, D2 were taken together to nominal current. At about 80% of the ramp, D2 quenched (5268 A). As expected, this caused the shutdown of the converters in the matching section and during the ramp down Q4 and Q5 also quenched. The data of this event is still being analysed.

The Electrical Quality Assurance team qualified the focusing and the defocusing quadrupole circuits and the matching section quadrupole Q6 at Point 7.

All the DFBs of the sector have been equipped with dry-air bags to ensure no condensation forms on the current leads and their chimneys.

Unfortunately, on Friday (June 1st) morning, the safety chain of cabinet was accidentally broken and the sector was disconnected from the pumping group. The cryo team is now busy recovering the situation to re-establish nominal temperature on the magnets for Monday morning.

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-06-03 - 20:28

Powering Tests Sector 7-8

PDF RAT_20070601.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x03x20x22/RAT_20070601.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-06-01 - 18:19

Powering Tests 7-8 - Evening Meeting 01/06/2007

ACTIVITY OF FRIDAY

* Triple quench in the matching section (Q5, Q4 and D2) is still under investigation. The current leads have been excluded as possible quench origin. New hardware will be implemented on Monday in order to better analyse the future events. In principle the quench in Q4 was not induced by the previous quench in D2, the delay is too short for that.

* The circuits in the DFBMH have been released by ElQA and QPS.

* 1.9K in the arc are not granted for Monday

PLAN FOR TODAY MONDAY

TEAM 1 - RR77 (& CCC)

* Connection of the four 600A current leads of the DFBMH (RQ6.R7B1 and RQ6.R7B2).

* PCC-Circuit Configuration tests in the circuits of DFBMH. Remark: the PC Permit signal from PIC will have to forced since the DFBMH belongs to the A78 powering subsector

* PIC2 in the DFBMH circuits (CCC)

TEAM 2 - CCC

* Powering of RQ5.L8 and its correctors to nominal current in intermidiate steps. Run for some hours at nominal current for validating the splices. If the test is successful the detection threshold of its quench detector will be risen to its normal operational value (100 mV)

ElQA

* If cryogenic conditions are recovered (2.1 K) ElQA of the RB.A78 circuit will continue. First, all the electronics of the DFBAs will be disconnected. If the test still fails the QPS electronics in the tunnel will be disconnected. Reiner ask for one hour advice for carrying out the disconnection.

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-31 - 18:34

Powering test in sector 7-8. 31 May - 01 June 2007

COMPLETED ON 31/05/07

Q5 Complete cycle of Q5 up to nominal current (4310 A) PNO.8d. Test OK.
Correctors of Q4 and Q5
Battery test of six correctors: PNO.3. Test OK.
Battery test of six correctors: PNO.4 (SPA). Not OK for FGC software problem.
Q4 corrector released by MPP after quench analysis. PNO.3. Test OK.
D2, Q4, Q5 Powering to nominal cycle of all the matching section ML8 together. Test not OK for quench in D2 and Q4. The quench occurred when all the circuits were at about 80% of the nominal current. Maximum was 5268 A for D2.
ElQA 13kA Completed for RQF and RQD.
In RB discharged at 1.7 kV (1.9 kV nominal). Cryo instrumentation was disconnected and next discharged was at 300 V. In the evening it was foreseen to repeat it with all the electronic disconnected.

PLAN FOR 01/06/07

ElQA HV tests dipole circuit
RQ6.R7 High voltage test to be repeated
RQ6.R7 in DFBMH Once ElQA finished and insulation plug restored
* QPS controller to be reprogrammed
* Current leads to be connected
* PCC Circuit Config
Matching Section ML8 Blocked by MPP till deeper understanding of the quench event. No powering tests till MPP releases it.

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-30 - 18:52

Powering test in sector 78 - 30-31 May 2007

COMPLETED ON 30/05/07

CORRECTOR CIRCUITS (Q4,Q5) RCBYVS4.L8B1; RCBYVS4.L8B2; RCBYV4.L8B1 went through PNO.1, PNO.2 and partially PNO.3
RCBYVS4.L8B2 had a training quench during PNO.2
RCBCH5.L8B1; RCBCV5.L8B2; RCBYH4.L8B2; RCBYHS4.L8B1 went partially through PNO.3
RQ4.L8
Unbalanced ramp up to nominal and unbalanced quench (PNO.8c)

Third trial for the unbalanced cycle up to nominal: B1 and B2 got to nominal current. In the ramp down of B1, due to a too the fast negative slope the converter tripped. B2 slow power abort commanded by PIC.
(Since B2 reached nominal current, there is not need to repeat the quench at nominal current in both apertures, PNO.7a is considered OK)
Provoked unbalanced quench with (B1 at nominal). PNO.8c OK.
Current cycle up to nominal PNO.9
RQ5.L8 Provoked unbalanced quench with (B1 at nominal). Ramp down from nominal this time worked OK. PNO.8c OK.

PROPOSAL OF THE ACTIVITY OF THURSDAY 31/05/07

CORRECTOR CIRCUITS (Q4,Q5) RCBYVS4.L8B1; RCBYVS4.L8B2; RCBYV4.L8B1; RCBCH5.L8B1; RCBCV5.L8B2; RCBYH4.L8B2; RCBYHS4.L8B1; should go through PNO.3 and PNO.4
RQ4.L8
RQ5.L8
Current cycle up to nominal PNO.9 on Q5.
As soon as they are released, the squeeze function test will be performed on these 2 circuits.

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-29 - 18:52

Powering test of sector 78 - minutes of the meeting of 29th Mat 2007 afternoon

PDF RAT_20070529b.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x29x18x50/RAT_20070529b.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-29 - 17:27

Powering Tests - Sector 7-8 - 30/05/2007

Yesterday was devoted to perform tests on RD2.L8 and RQ4.L8

    • RD2.L8 has been successfully commissioned up to PNO.8d (Current cycle up to nominal current)
    • Missing tests for RD2.L8: Fast Power Abort from PIC (PNO.9). The detection threshold needs to be modified. Possibly done while powering at unison (like PNO.9 for RQ5)
    • RQ4.L8:
        • Unbalanced cycle up to nominal (3610 A) for B1 and 2000 A for B2. While ramping B1 converter tripped. The attempt of PO expert to recover control made the quench detector trip (same problem as last week)
        • Second trial of unbalanced cycle up to nominal: with B1 at nominal current and B2 ramping up to nominal, the latter quenched at 3590 A. Test failed.

Proposal for the activity of today (30/05/2007)

  • Complete P2N of all the 120-80A correctors of Q4 and Q5.
  • Continue P2N of RQ4.L8. Two options: * Unbalanced ramp up and quench with both apertures at nominal current (PNO.7a) or/and * Unbalanced cycle and quench with one aperture at nominal and the other at 50%.

  • ElQA activities: Morning: * Full qualification of the 3 circuits selected on DFBAN * TFM on circuits powered via DFBMH * preparation for the HV qualification of the 600A circuits Afternoon * HV qualification on 600 A circuits powered via DFBAO

  • QPS: Interlock tests .

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2007-05-29 - 10:57

Powering Tests of Sector 7-8: minutes of the meetings of 29/05

PDF RAT_20070529.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x29x10x56/RAT_20070529.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-25 - 21:28

Powering Tests in Sector 7-8: activity of Friday 25/05.

The day was fully devoted to perform tests on D2.

The following tests have been executed on D2 (MANUAL execution, due to change in procedure):
  • PIC2 (SEQUENCER execution); test successful;
  • PLI1: ramp to 350 A with 10 A/s ramp rate, we stayed on the plateau 4 min and then provoked a FPA; test successful;
  • PLI2: ramp to 350 A with 10 A/s ramp rate, then ramp to 1000 A with 10 A/s, and quench driven by QPS, generating an artificial quench signal from one current lead detector; the test was successful;
  • PNO: during the first attempt to ramp to 3000 A, the power converter went to ‘faults’ and produced a FPA when reaching 1008 A; investigations of PO showed that a hardware limitation was present. The second attempt was successful:
    • Ramp to 350 A with 10 A/s ramp rate;
    • Continue ramping to 3000 A; plateau of 4 min;
    • Continue ramping to 6000 A (nominal current); plateau of 30 min to check voltage drift;
    • Provoked quench at nominal current.

Proposal for the activity of next week.

  • Continue powering Q4 and D2
  • Continue powering the correctors (4 circuits have already passed the first 2 steps, three more are available).
  • Possible powering of the whole matching section.
  • Powering of the arc 60 A orbit correctors, according to humidity conditions.

  • Continue installation of plastic bags on DFBAN and DFBMH.
  • Check the status of condensation of the 13kA leads of DFBAO which have copper laminations and no plastic bags; check the status of the 6kA lead of DFBMA which has copper laminations and no plastic bags.
  • Installation of ventilators on the 60 A orbit correctors leads.

  • ElQA on the selected circuits: see presentation of Davide in the last post.

  • Quench heaters qualification, without interference with ElQA.

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-25 - 21:15

Powering Tests of Sector 7-8: minutes of the meetings of 25/05.

PDF Q4.L8.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x25x21x07/Q4.L8.pdf

PDF RAT_20070525.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x25x21x07/RAT_20070525.pdf

PDF ELQA_CCC_25052007.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x25x21x07/ELQA_CCC_25052007.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-24 - 19:26

Powering tests in Sector 7-8: activity of Thursday 24/05.

The day was fully devoted to perform tests on Q4, which had experienced an unexpected quench at 300 A on Monday (after ramping at 600 A).

The following tests have been executed on Q4 (MANUAL execution, due to change in procedure):
  • ramp to 600 A with 5 A/s and quench driven by QPS, generating an artificial quench signal from one current lead detector; the test was successful;
  • ramp of B1 power converter to nominal current (3610 A) with 10 A/s ramp rate and simultaneous ramp of B2 power converter to 2000 A with nearly 5 A/s (unbalanced configuration):
    • when B1 current reached 1025 A, its reference current went anormally down to 0;
    • later analysis showed that a fault was sent from the converter and a consequent FPA was started;
    • PO attempt to get control of B2 power converter, to ramp it down contemporarily to 0, provoked a sudden voltage peak in B1, which initiated a quench;
  • at 11.48 (according to PIC supervision) a power cut occured on the general services of UA83, which eventually provoked the loss of controls on cryogenics and lead heating system;
  • after waiting for the good cryo conditions to be recovered, and in agreement with PO, QPS and MPP experts, we decided to ramp again Q4 with the same unbalanced cycle:
    • B1 at 3610 A with 10 A/s ramp rate and B2 to 2000 A with a nearly 5 A/s ramp rate;the test was successful and NO quench was activated;
    • while ramping B2 power converter to the same current as B1 (the idea was to stay on plateau for half an hour and then provoke a quench), and with a current in B2 of 2900 A, a quench generated; preliminary analysis indicated that it was started in B1, possibly due to electromagnetic coupling (this configuration has never been tested before).
  • end of the tests due to long recovery time.

Proposal for the activity of tomorrow.

  • Continue powering Q4 and D2.

  • Continue installation of plastic bags on 600 A current leads.
  • Lower the temperature on the 13 kA leads to check the absence of condensation with AT/MEL copper laminations.

  • ElQA on the selected circuits: list will be provided during the morning.

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-05-24 - 15:02

Powering Tests Sector 7-8 - Minutes 24/05/2007

PDF RAT_20070524b.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x24x14x54/RAT_20070524b.pdf

PowerPoint Q5_test_240507.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x24x14x54/Q5_test_240507.ppt

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-24 - 10:06

Sector 7-8 cooldown

After becoming the first magnet at 4.5K and in the kA range to be powered last week, the Q5 matching section quadrupole (MQM) left of Point 8 became the first one to reach nominal current on May 23 @ 15:04:28. The two apertures of the magnet were then powered at different current levels and a quench was provoked. Following that, the magnet was taken to nominal current. This magnet, which was suspected to require training quenches, reached nominal current after two quenches (3910A and 4224 A) before reaching nominal current (4310A) [see the HC Twiki pages for details].

Commissioning now proceeds with Q4 and then D2 as well as the associated orbit correctors the same region.

In the meantime, the cryo team has continued with the tuning of the pumping group and the 4.5K refrigerator and has started cooling down the arc down to 1.9K. This morning (Thursday, May 24) half of the arc was below lambda point. The Power Converter, Interlock, Current Lead, Electrical Quality Assurance and Quench Protection System teams have started the preparation for the powering of the arc early next week.

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-23 - 19:02

Powering Tests of Sector 7-8: activity of Wednesday 23/05.

Today the activity was concentrated on Q5, which was released by MPP after analysis of the quench of Monday morning.

The following tests have been executed on Q5 (MANUAL execution):
  • ramp to nominal (4310 A) with 5 A/s ramp rate: quenched occurred at 4210 A on B2 aperture; preliminary analysis confirmed it was a training quench;
  • ramp to nominal (4310 A) with 5 A/s ramp rate: NO quench; a plateau of 10 min was mantained for possible voltage drift recording;
  • keeping B1 power converter at nominal current, B2 power converter was lowered to 55% of the nominal current (2370 A); with this unbalanced configuration, a quench was driven by QPS generating an artificial quench signal from one current lead detector; test was successful;
  • ramp to nominal (4310 A) with 5 A/s ramp rate: NO quench; a FPA was provoked from power converter, with an apparently too high ramp rate which triggered the quench detector.

Proposal for the activity of tomorrow.

Q4 and D2 will be tested from CCC.

Works towards powering the main circuits in the arc and the 600 A correctors (before reaching 1.9 K):

  • plastic bags on 600 A leads (action S. Kockorowski);
  • IST of QPS for which 1.9 K is not required (action R. Denz);
Once 1.9 K are reached:
  • ElQA for the three 13 kA circuits and some selected 600 A correctors (list to be provided by HCC; action D. Bozzini);
  • IST of QPS at 1.9 K (action R. Denz);
  • OK from PIC, PC, EE and Current Leads.

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-05-23 - 12:01

Powering Tests Sector 78 - Minutes 23/05/2007

PDF RAT_20070524.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x23x11x59/RAT_20070524.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-22 - 22:46

Powering Tests in Sector 7-8: activity of Tuesday 22/05.

The Powering to Nominal procedure was started fro the following circuits:

  PNO.1 PNO.2
RCBCH5.L8B1 O O
RCBCv5.L8B1 O O
RCBYH4.L8B2 O O
RCBYHS4.L8B1 O O

Several problems were faced:

  • cryogenics instabilities which delayed the CRYO START signal:
    • oscillating level on DFBMC
    • high temperature on the first 6 kA in DFBMA (54 K)
  • delay for PNO.2 step for RCBYH4.L8B2, whose origin was:
    • hardware limitation of the PC, which blocked the cycle to 52 A; the PC went on fault and the current drop to zero (intervention of PO to remove the limitation)
    • probably due to the problem just mentioned, a very long PM file was produced which was preventing the cycle from being restarted (manual reset of the PC was needed)
  • RCBYHS4.L8B2 showed a possible NON-CONFORMITY: while preparing this circuit for PNO.1, a high voltage value was detected on a lead (250 mV); a verification of PO showed that this is probably due to an open voltage tap. The circuit has been removed from the SOC.

PROPOSAL FOR WEDNESDAY 23/05:

  • continuation of tests on RCBCH5.L8B1, RCBCv5.L8B1, RCBYH4.L8B2 and RCBYHS4.L8B1;
  • Powering to Nominal procedure for circuits RCBYV4.L8B1, RCBYVS4.L8B1 and RCBYVS4.L8B2;
  • continuation of tests on Q4 and Q5 if released by MPP;
  • continuation of tests on D2.

POSSIBLE SCENARIOS FOR NEXT WEEK (N.B.: PC, PIC, QPS, EE systems and experts have to be ready for next Tuesday!!!)

RB.A78, RQF.A78 (?), RQD.A78 (?) Powered by their own converter DFBA, QPS, EE, PC, PIC Preparation (this week)
ARC @ 1.9K ElQA @ 1.9K RB & RQD
QPS Heater Discharge
PIC1 RB & RQF
Connection DC cable - lead
PC_CONFIG RQ - 350A
PIC 2 RQ - 350A
PC_CONFIG RB - 350A
PIC 2 RB - 350A
P2N RQ - Max 760A
P2N RB - Max 760A
 
Powered by a 120A converter DFBA, QPS, EE, PC, PIC Preparation (this week)
ARC @ 4.5K PIC1 RB & RQD - Without Quench Heaters
Connection DC cable - lead
PC_CONFIG RQ - 100A
PC_CONFIG RB - 100A
 
600A Circuits - Priority 1, 25 Circuits ARC @ 4.5K or 1.9K DFBA, QPS, EE, PC, PIC Preparation (this week)
ElQA @ 4.5K-1.9K first Circuits
QPS Last Conditioning
PC_CONFIG first circuits
ARC @ Superfluid Helium DFBA, QPS, EE, PC, PIC Preparation (this week)
ElQA @ 4.5K-1.9K All Circuits
QPS Last Conditioning
PIC2 First Circuits
PCS First Circuits
PC_Config, PIC2, PCS rest of circuits - Parallel

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-22 - 17:47

Warm Magnets Commissioning Point 8 - List of activities for this week (21)

Comments:

List of activities agreed at the Warm Magnets Commissioning Point 8 meeting on 23rd May (8:30 in 936/R-035)

  WM8 LHCb
23.05 MORNING Verification signalization in the tunnel.

Consignation all TI8 LHC PC.

Disconnection of the DC cables to the TI8 LHC PC.

WELQA on the MSI magnets
 
   
   
   
   
   
   
23.05 AFTERNOON WELQA on the rest of the TI8 LHC magnets.

Connection of the DC cables to the TI8 LHC PC.

Déconsignation all TI8 PC.
 
24.05 MORNING Verification signalization in the tunnel.

Repeat the WELQA in the circuits where it could not be completed.

Connection of the DC cables to the TI8 LHC PC.

Déconsignation all TI8 PC.

Polarity check TI8 LHC magnets.

From 13:00 to 14 :30

Déconsignation PC compensators L8:

1. RPTN.SR8.RBXWSH.L8
1. RPTN.SR8.RBXWH.L8

Switch on RPTN.SR8.RBXWSH.L8 and verification at the magnet level.

Switch off RPTN.SR8.RBXWSH.L8

Switch on RPTN.SR8.RBXWH.L8 and verification at the magnet level.

Switch off RPTN.SR8.RBXWH.L8







From 13:00 to 14:30
No work around the ANY compensator magnets
24.05 AFTERNOON Start 24h Heat Run (TI8 LHC + COMPENSATORS L8) (I ultimate) No work around the compensator magnets L8
25.05
MORNING
24h Heat Run ongoing

Temperatures measurements at the end 24h Heat Run
No work around the compensator magnets L8
25.05
AFTERNOON
Turn on all TI8 magnets PC, RPTN.SR8.RBXWSH.L8 and RPTN.SR8.RBXWH.L8.

Start 1st Interferences Evaluation

Start solution on non conformity on power converter of MQIF87800 (3e-4 ripple)
No work around the compensator magnets L8

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-22 - 11:42

Powering Tests Sector 7-8 - Minutes 22/05/2007

PDF RAT_20070523.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x22x11x37/RAT_20070523.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-22 - 00:23

Powering tests of Sector 7-8: activity of Monday, 21

Morning activity. The following tests have been performed on Q5 (MANUAL execution):

  • verification of p-p voltage vs ramp rate (problems with high ramp rates had been discovered last week): ramping up to 200A with 5 A/s, to 400A 7 A/s and 600A with 9 A/s;
  • HCA PLI3.1a: ramp to 2500 A with 5 A/s ramp rate;
  • HCA PLI3.1b, unbalanced cycle: starting from 120 A, one converter ramped to 2500 A at 5 A/s, the other ramped to 1375 at 2.5 A/s;
  • HCA PLI3.1c: skipped after agreement between experts from the different systems;
  • HCA PLI3.4: no icing confirmed by Amalia;
  • HCA PNO.5: ramp to nominal (4310 A); during the ramp, and at the current value of 3941 A, a quench occurred (analysis on going).
Afternoon activity. The following tests have been performed on Q4 (SEQUENCER execution):
  • HCA PLI1.1:
    • first test aborted because the FGC server didn't autorized lhcop to access the power supplies (intervention of S.Page);
    • test stopped, this time because of a cmw time out exception;
    • test failed again because the software current limit have not been set correctly in the FGC;
    • test OK;
  • HCA PLI1.4: FPA at 250 A;
  • HCA PLI1.7: ice absence verified;
  • HCA PLI2.1: current to interm1 (600 A); current reached without problems, but during ramping down the sequencer blocked because of an attemp to reduce the needed time to stand-by
  • HCA PLI2.1: repeated, but during the ramp, around 300 A, a quench occurred (investigations on going).

Proposal for tomorrow:

  • MPP will analyse the data collected on Q4 and Q5, with the feedback from QPS experts;
  • the test activity will focus on 120 A correctors.

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-21 - 10:38

Powering Tests Sector 7-8 - Minutes 21/05/2007

PDF RAT_20070522.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x21x10x36/RAT_20070522.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-20 - 22:20

Powering Tests - 1000A in Q5 -

Done on Friday 18.05.2007

9:30 Powering Procedure for RQ5B1(B2).L8 started
11:30 PLI1.1-4-7 I_inj: 350A
PLI2.1-3-4 I_int_1: 1000A
procedures performed manually and NOT using the sequencer
13:00 Boil-off test on DFBMA-C

Remarks on the work of the last week (14-18 May)

The current was ramped to 1000 A, and the magnet was quenched by triggering the firing of the quench heaters.
The different teams performed a preliminary analysis of the data that had been taken, and some observations were reported:
The DECOUPLING BETWEEN THE TWO CONVERTERS WAS EXCELLENT. By applying a modulation on one power converter of 10A at a DC current of 1000A it was measured that the coupling is negligible (ppm level). The modulation was seen by the quench detectors, but was below the threshold. For future tests, it was suggested to include the current modulation into the procedures, if required (Action: D.Nisbet).
The balancing of the QUENCH DETECTORS was very good. The THRESHOLD was set to 20mV, and the detectors did not trigger. However, it was noted that the PEAK AMPLITUDE OF THE BALANCED SIGNAL WAS SLIGHTLY ABOVE 20mV, BUT FOR A TIME LESS THAN THE VALIDATION TIME OF THE QUENCH DETECTORS OF 10ms. This effect depends on the speed of the ramp, and decreased between the first ramp at 13A/s AND THE SECOND RAMP AT 10A/s.
FOR THE FUTURE COMMISSIONING OF Q5 A RAMP OF 5A/S WILL BE APPLIED, in order to reduce the noise level for the quench detectors. The modified value should be integrated into the sequence. In the second phase the sensitivity of the quench detectors as a function of ramp speed will be further investigated. It might be required to increase the quench detector threshold from 20mV to a higher value. The objective is to ramp the magnet with the nominal rate of about 10A/s.
At the first glance it appears that all required data after the quench were recorded and written into the PM system.
The heater discharge worked as expected. It was correctly recorded, but due to one point with the wrong time the start of graph was wrong. This will be corrected (Action: R.Denz and A.Rijllart).
THE DATA FROM THE CURRENT LEADS (VOLTAGE ETC.) LOOKED OK. It was noticed that the voltage was high during the trigger of the heaters. This is due to the fact that the quench heaters are triggered using a test signal, and that the effect does not indicate an increased voltage across the leads.
AFTER TRIGGERING THE QUENCH, THE HELIUM LEVEL IN THE Q5 MAGNET CRYOSTAT DECREASED IN A SHORT TIME (~MIN) TO ZERO, and increased in about the same time to the nominal value. No change of temperature and pressure were observed. The valve opened after the quench. Since similar changes of the helium level were recorded at other occasions without quenching magnets, it is questions if this effect is real. This needs follow-up, since such drop of helium level would trigger a power abort (Action: ACR). OTHERWISE NO CONSEQUENCES OF THE QUENCH WERE OBSEREVED IN THE CRYOGENIC SYSTEM.
Since the procedures were executed without the sequencer, that data that was recorded needs to be put into MTF manually (Action: HCC)

Program for Monday 21.05.2007

Continue commissioning of Q5, manually without the sequencer.
Start commissioning of the 120A circuits using the sequencer. This can be done in parallel to Q5 commissioning, although there might be some interference.
When Q5 is commissioned, Q4 and D2 will follow. Before starting to commission D2,y the QPS team needs to intervene in order to correct a problem that was reported after PIC2.

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-05-18 - 17:39

RAT meeting 18/5/2007

PDF RAT_20070518.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x18x17x29/RAT_20070518.pdf

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-18 - 15:01

Warm Magnet Commissining Point 8 - Minutes 10.05.07 and planning

PDF Planning.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x18x14x31/Planning.pdf

PDF WM8minutes-presentation.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x18x14x31/WM8minutes-presentation.pdf

PowerPoint Warm_magnets_commissioning_point_8.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x18x14x31/Warm_magnets_commissioning_point_8.ppt

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-17 - 23:49

Powering Tests - Sector 78 -

Done on Wednesday 16.05.2007

9:30 PIC2 on corrector circuits on DFBMA/C
10:30 PIC2 on corrector circuits on DFBMA/C completed
11:00 Problems on RQ4.L8 solved.
12:00 PCC on RQ4.L8 completed
14:30 Problems on RQ5.L8 solved.
15:20 PCC on RQ5.L8 completed
15:30 PIC2 on D2-Q4-Q5.L8
18:00 PIC2 on RD2-RQ4-RQ5.L8 completed.
18:00 Last test of PIC2 not passed “PIC2.10 Test hardware links”
18:15 Tests end

Schedule for Friday 18.05.2007 in sector 7-8

WHEN CONDITIONS FOR THE ElQA IN THE DFBAs ARE MET:

DFBAO ElQA continuity test
DFBAN ElQA continuity test

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL) IN THE POWERING SUBSECTOR ML8:

Sub sector ML8 ready for Powering to Nominal
D2-Q4-Q5.L8 Powering to nominal up to PLI2 (<1000A) – Manually
Corrector circuit 120A DFBMC-A Powering to nominal –Via sequencer

WHEN CONDITIONS IN ARC ARE MET GO ON WITH THE PROGRAMME FOR 60A CONVERTERS (ORBIT CORRECTORS) AND 120A (Q7 TO Q10 CORRECTORS) AS WELL AS THE ELQAs IN THE DFBAs X: previous day - X : last day - O : Proposal

  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-16 - 15:50

Powering Tests Sector 7-8 - Minutes 16/05/2007

PDF RAT_20070516.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x16x15x44/RAT_20070516.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-16 - 12:25

Powering Tests: 16 May 2007

Program for 16.05.2007 - afternoon

14:00 PCC for Q5
15:00 PIC2 for D2,Q4,Q5: Ivan and Boris from the CCC, the rest in the FCC If things go smoothly
16:00 Start P2N, preferably Q5, final decision around 15:00. In the sequence a quench at 1000A will be provoked (if we reach this far)

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-15 - 18:59

Powering test - Sector 78 - Tuesday 15th

Done on Tuesday 15.05.2007

9:30 Test start on D2 – QH on.
9:40 Up to ~90A apparently without QPS protection – Actually It was not the case.
15.24 QH enabled for RD2.L8, QPS interlock problem on D2 understood. Only the displayed ref. voltage was a sum of the voltage picked up from the two apertures instead of a difference. At the level of protection the signal was right.
15:34 PCC start on D2 - up to 200A at 16:00
17:30 Loss of Cryo-Start and Cryo-Maintain due to a hardware intervention. Physical cryo conditions were stable. Signals recovered after few minutes.
16:23 PCC for D2 finished successfully
17:40 HQ of RQ4.L8 on – Reiner changed
17:50 Still problems on Q4 – Picked voltage exceed 20mV when PC tries to start up – QPS trips the PC.
   
19:00 Tests end

Schedule for Wednesday 16.05.2007 in sector 7-8

WHEN CONDITIONS FOR THE ElQA IN THE DFBAs ARE MET:

DFBAO ElQA continuity test
DFBAN ElQA continuity test

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL) IN THE POWERING SUBSECTOR ML8:

RQ4.L8 AT-MEL General re-check
AB-PO - Special test with a 60A power converter – charatizarion
DFBMC-A PIC2 on the 120A corrector circuits.

PC CONFIGURATION IN RQ4

DFBMA RQ4.L8
RD2.L8

Balisage with flashing lights HCC
RQ4.L8
RD2.L8
Warn the QPS team that the main circuit will be powered HCC
RQ4.L8
RD2.L8
Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ4.L8
RD2.L8
Unlock and switch on the corresponding power converters AB/PO
RQ4.L8
RD2.L8
PC configuration
UP TO 100 A
AB/PO
PC CONFIGURATION IN RQ5
DFBMC RQ5.L8

Balisage with flashing lights HCC
RQ5.L8 Warn the QPS team that the main circuit will be powered HCC
RQ5.L8 Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ5.L8 Unlock and switch on the corresponding power converters AB/PO
RQ5.L8 PC configuration
UP TO 100 A
AB/PO
WHEN CONDITIONS IN ARC ARE MET GO ON WITH THE PROGRAMME FOR 60A CONVERTERS (ORBIT CORRECTORS) AND 120A (Q7 TO Q10 CORRECTORS) AS WELL AS THE ELQAs IN THE DFBAs X: previous day - X : last day - O : Proposal
  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-15 - 16:13

Powering Tests Sector 7-8 - Minutes 15/05/2007

PDF RAT_20070515.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x15x16x12/RAT_20070515.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-14 - 20:42

Powering test - Monday 14.05 -

Done on Monday 14.05.2007

9:30 PCC on 120A correctors of DFBMC/A
10:30 Deconsignation of Q5-Q4 PC
11:00 Problems with Q5 – Water faults – TS/CV came at 13:30 – they cleaned the filters – Problem recovered.
14.30 AB-PO started to tests D2 with a 60A PC connected to the magnet (circuit definition)– DC cable disconnected from PC side.
14:42 Logging not working. Need to disactivate and re-activate the SOC from SOC editor to restore it. Monitoring problem RCBCH5 and RCBCV5 stopped transmitting data to the monitoring system. Expects contacted. System restarted. Problem disappeared. Systected 'middleware' problem in the network - problem is known and happens once a week.
14.50 Create and activate new SOC named PCC_IPDQ140507 with circuits RQ4.L8 RQ5.L8 and RD2.L8.
15.30 Deconsignation of QH for Q5-Q4-D2
16.30 PCC tests for SOC PCC_070510 finished (corrector circuits DFBMA/C)
16:40 Testing communication between cryo-maintain signal and PIC, DFBMA,DFBMC
16.45 End of the RD2.L8 charaterization
17.30 RQ4.L8B1 and RQ4.L8B2 ready for PCC
17.30 QPS detecting voltages of around ~40-50mV from the magnet protection and trip the PC (not current leads). These trips occur immediately after attempting to ramp the power converters. Note that the QPS trip threshold was reduced from 100mV to 20mV.
18.30 Same problem on Q5.
19:00 Tests end

Schedule for Tuesday 15.05.2007 in sector 7-8

WHEN CONDITIONS FOR THE ElQA IN THE DFBAs ARE MET:

DFBAO ElQA continuity test
DFBAN ElQA continuity test

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL) IN THE POWERING SUBSECTOR ML8:

PC CONFIGURATION IN RQ4, RD2 AND RQ5

DFBMA RQ4.L8
RD2.L8

Balisage with flashing lights HCC
RQ4.L8
RD2.L8
Warn the QPS team that the main circuit will be powered HCC
RQ4.L8
RD2.L8
Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ4.L8
RD2.L8
Unlock and switch on the corresponding power converters AB/PO
RQ4.L8
RD2.L8
PC configuration
UP TO 100 A
AB/PO

DFBMC RQ5.L8

Balisage with flashing lights HCC
RQ5.L8 Warn the QPS team that the main circuit will be powered HCC
RQ5.L8 Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ5.L8 Unlock and switch on the corresponding power converters AB/PO
RQ5.L8 PC configuration
UP TO 100 A
AB/PO
WHEN CONDITIONS IN ARC ARE MET GO ON WITH THE PROGRAMME FOR 60A CONVERTERS (ORBIT CORRECTORS) AND 120A (Q7 TO Q10 CORRECTORS) AS WELL AS THE ELQAs IN THE DFBAs X: previous day - X : last day - O : Proposal
  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

**

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-14 - 10:54

Powering Tests Sector 7-8. Minutes 14/05/2007

PDF RAT_20070514.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x14x10x53/RAT_20070514.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-11 - 16:26

Schedule for Monday 14.05.07 in sector 78

Schedule for Monday 14.05.2007 in sector 7-8

WHEN CONDITIONS FOR THE ElQA IN THE DFBAs ARE MET:

DFBAO ElQA continuity test
DFBAN ElQA continuity test

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL) IN THE POWERING SUBSECTOR ML8:

DFBMA RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
Balisage with flashing lights HCC
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
PC Configuration (test of PC and magnet at minimum current)
UP TO 10 A
AB/PO

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Balisage with flashing lights HCC
RCBCH5.L8B1
RCBCV5.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
PC Configuration (test of PC and magnet at minimum current)
UP TO 10 A
AB/PO

PC CONFIGURATION IN RQ4, RD2 AND RQ5

DFBMA RQ4.L8
RD2.L8

Balisage with flashing lights HCC
RQ4.L8
RD2.L8
Warn the QPS team that the main circuit will be powered HCC
RQ4.L8
RD2.L8
Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ4.L8
RD2.L8
Unlock and switch on the corresponding power converters AB/PO
RQ4.L8
RD2.L8
PC configuration
UP TO 100 A
AB/PO

DFBMC RQ5.L8

Balisage with flashing lights HCC
RQ5.L8 Warn the QPS team that the main circuit will be powered HCC
RQ5.L8 Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ5.L8 Unlock and switch on the corresponding power converters AB/PO
RQ5.L8 PC configuration
UP TO 100 A
AB/PO
WHEN CONDITIONS IN ARC ARE MET GO ON WITH THE PROGRAMME FOR 60A CONVERTERS (ORBIT CORRECTORS) AND 120A (Q7 TO Q10 CORRECTORS) AS WELL AS THE ELQAs IN THE DFBAs X: previous day - X : last day - O : Proposal
  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

Done on Thursday 10.05.2007

9:30 Swith the Power Converters on: RPMBA.UA83.RQT12.L8B1 and RPMBA.UA83.RQT12.L8B2
Beginning of the validation of the sequencer for the procedure of Q2.
Test on the communication between the PIC and the CRYO signals (CRYO start and CRYO maintain).
10:00 Deconsignation of the PCs of the corrector circuits of Q4 and Q5.
CRYO control in the FCR is not available, except a viewer.
A voltage tap in the circuit RCBYHS4.L8B2 has been detected, the problem will be investigated.
11:30 Waiting for solving the voltage problem.
Consignation of the PCs of the circuits: RPHGA.UA83.RQ9.L8B2 RPHGA.UA83.RQ9.L8B1 RPHGA.UA83.RQ10.L8B2 RPHGA.UA83.RQ10.L8B1 in order to allow AT/MEL to install new heaters.
14:00 Consignation of all the circuits of the converters of the Q4 in order to allow Vincent Chareyre to check the problem of the voltage drop in the circuit RCBYHS4.L8B2.
15:00 Vincent Chareyre has fixed the problem.
15:20 start with the PC configuration of the circuit RCBCH5.L8B1.
16:00 16:10 CRYO-OK lost.
16:20 CRYO-OK found again.
PC configuration of the circuit RCBCH5.L8B1 is ongoing.
17:00 Start with the PC configuration of the RCBCV5.L8B2.
17:10 CRYO-maintain forced LOST and PIC removed all the Power permits. Consequently all the PCs tripped off.
17:30 Cryo Operator reported on the situation: the CRYO OK was lost and the CRYO maintain was going to be lost, consequently they forced the CRYO-MAINTAIN as lost.
18 :00 Switch the Power Converters off: RPMBA.UA83.RQT12.L8B1 and RPMBA.UA83.RQT12.L8B2.

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-11 - 11:27

Powering Tests Sector 78 - RAT Minutes - 11/05/2007

PDF RAT_20070511_2.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x11x11x24/RAT_20070511_2.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-10 - 18:50

Schedule for Friday 11.05.2007 in sector 7-8

Schedule for Friday 11.05.2007 in sector 7-8

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL) IN THE POWERING SUBSECTOR ML8:

PC CONFIGURATION IN THE Q4 AND Q5 CORRECTOR CIRCUITS

DFBMA RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
Balisage with flashing lights HCC
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
PC Configuration (test of PC and magnet at minimum current)
UP TO 10 A
AB/PO

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Balisage with flashing lights HCC
RCBCH5.L8B1
RCBCV5.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
PC Configuration (test of PC and magnet at minimum current)
UP TO 10 A
AB/PO

PC CONFIGURATION IN RQ4, RD2 AND RQ5

DFBMA RQ4.L8
RD2.L8

Balisage with flashing lights HCC
RQ4.L8
RD2.L8
Warn the QPS team that the main circuit will be powered HCC
RQ4.L8
RD2.L8
Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ4.L8
RD2.L8
Unlock and switch on the corresponding power converters AB/PO
RQ4.L8
RD2.L8
PC configuration
UP TO 100 A
AB/PO

DFBMC RQ4.L8

Balisage with flashing lights HCC
RQ4.L8 Warn the QPS team that the main circuit will be powered HCC
RQ4.L8 Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ4.L8 Unlock and switch on the corresponding power converters AB/PO
RQ4.L8 PC configuration
UP TO 100 A
AB/PO
WHEN CONDITIONS IN ARC ARE MET GO ON WITH THE PROGRAMME FOR 60A CONVERTERS (ORBIT CORRECTORS) AND 120A (Q7 TO Q10 CORRECTORS) AS WELL AS THE ELQAs IN THE DFBAs

X: previous day - X : last day - O : Proposal

  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

REPORT ON THE ACTIVITIES OF THURSDAY 10TH MAY

09:30 Swith on of the Power Converters: RPMBA.UA83.RQT12.L8B1 and RPMBA.UA83.RQT12.L8B2
Beginning of the validation of the sequencer for the procedure of D2.
Test on the communication between the PIC and the CRYO signals (CRYO start and CRYO maintain).
10:00 Deconsignation of the PCs of the corrector circuits of Q4 and Q5.
CRYO control in the FCR is not available, except a viewer.
A voltage tap problem in the circuit RCBYHS4.L8B2 has been detected, the problem will be investigated.
11:30 Waiting for solving the voltage problem.
Consignation of the PCs of the circuits: RPHGA.UA83.RQ9.L8B2 RPHGA.UA83.RQ9.L8B1 RPHGA.UA83.RQ10.L8B2 RPHGA.UA83.RQ10.L8B1 in order to allow AT/MEL to install new heaters.
14:00 Consignation of all the circuits of the converters of the Q4 in order to allow Vincent Chareyre to check the problem of the voltage drop in the circuit RCBYHS4.L8B2.
15:00 Vincent Chareyre has fixed the problem.
15:20 start with the PC configuration of the circuit RCBCH5.L8B1.
16:00 16:10 CRYO-OK operation lost.
16:20 CRYO-OK operation found again.
PC configuration of the circuit RCBCH5.L8B1 is ongoing.
17:00 Start with the PC configuration of the RCBCV5.L8B2.
17:10 CRYO-maintain forced LOST and PIC removed all the Power permits. Consequently all the PCs tripped off.
17:30 Cryo Operator reported on the situation: the CRYO OK was lost and the CRYO maintain was going to be lost, consequently they forced the CRYO-MAINTAIN as lost.
18 :00 Switch off of the Power Converters: RPMBA.UA83.RQT12.L8B1 and RPMBA.UA83.RQT12.L8B2.

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-10 - 18:26

Powering Test Sector 78 - RAT Meeting - 10/05/2007

PDF RAT_20070510.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x10x18x17/RAT_20070510.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-10 - 08:37

Schedule for Thursday 10.05.2007 in sector 7-8

Schedule for Thursday 10.05.2007 in sector 7-8

FINAL CHECK ON CIRCUITS RPMBB.UA83.RCD.A78B1 AND RPMBB.UA83.RCD.A78B2 TO MAKE THEM AVAILABLE FOR SEQUENCER TESTS ON D2 TYPE PROCEDURE.

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL) IN THE POWERING SUBSECTOR ML8:

PC CONFIGURATION IN THE Q4 AND Q5 CORRECTOR CIRCUITS

DFBMA RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
Balisage with flashing lights HCC
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
PC Configuration (test of PC and magnet at minimum current)
UP TO 10 A
AB/PO

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Balisage with flashing lights HCC
RCBCH5.L8B1
RCBCV5.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
PC Configuration (test of PC and magnet at minimum current)
UP TO 10 A
AB/PO

PC CONFIGURATION IN RQ4, RD2 AND RQ5

DFBMA RQ4.L8
RD2.L8

Balisage with flashing lights HCC
RQ4.L8
RD2.L8
Warn the QPS team that the main circuit will be powered HCC
RQ4.L8
RD2.L8
Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ4.L8
RD2.L8
Unlock and switch on the corresponding power converters AB/PO
RQ4.L8
RD2.L8
PC configuration
UP TO 100 A
AB/PO

DFBMC RQ4.L8

Balisage with flashing lights HCC
RQ4.L8 Warn the QPS team that the main circuit will be powered HCC
RQ4.L8 Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ4.L8 Unlock and switch on the corresponding power converters AB/PO
RQ4.L8 PC configuration
UP TO 100 A
AB/PO
WHEN CONDITIONS IN ARC ARE MET GO ON WITH THE PROGRAMME FOR 60A CONVERTERS (ORBIT CORRECTORS) AND 120A (Q7 TO Q10 CORRECTORS) AS WELL AS THE ELQAs IN THE DFBAs

X: previous day - X : last day - O : Proposal

  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

Done on Monday 07.05.2007

SEQUENCER TESTS FOR THE PROCEDURE Q4Q5

DFBAO RQ6.L8B1
RQ6.L8B2
Balisage around the chimneys HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. Check of the short-circuit and the insulation. AB/PO
Sequencer tests on powering procedures for 6kA magnets with two power converters ("Q4Q5" procedure) AB/CO
Lock and switch off the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters AB/PO

INSULATION TEST AT THE DFBMA : RESULT OK

CIRCUIT 60A RCBV12.L8B1 INTERVENTION BY AT/MEL: CUPPER COIL PLACED AROUND A CURRENT LEAD

DRY BAGS AND VENTILATORS INTALLATION AT THE DFBMC AND DFBMA BY AT/ACR

“BAD QUALITY OF THE HELIUM LIQUID IN THE LINE CPREVENTED THE POWERING IN ML8 IN THE AFTERNOON”

Done on Tuesday AND Wednesday 08/09.05.2007

GO ON WITH THE VALIDATION OF THE SEQUENCER FOR THE PROCEDURE Q4Q5

DFBAO RQ6.L8B1
RQ6.L8B2
Balisage around the chimneys HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. Check of the short-circuit and the insulation. AB/PO
Sequencer tests on powering procedures for 6kA magnets with two power converters ("Q4Q5" procedure) AB/CO
Lock and switch off the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters AB/PO

INTERVENTION OF THE GEOMETERS AT D2 Q4 (Beginning AT 9:00)

MEASUREMENTS AT THE DFBMs (ML8)

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-05-08 - 17:36

RAT 7/5/2007

PowerPoint RAT2.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x08x17x17/RAT2.ppt
Presentation by A. Ballarino
PDF RAT_20070508.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x08x17x17/RAT_20070508.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-07 - 22:54

Anti-condensation systems tested on DFBMA and DFBMC.

JPEG DSC03896b.jpg:
DSC03896b.jpg
Status of the DFBMA _ Plastic envelopes on all chimneys plus additional heater on top of the 120A-8leads-assembly and air blower in front of it (maximum power).
JPEG DSC03893b.jpg:
DSC03893b.jpg
Status of the DFBMC _ Plastic envelopes on all chimneys plus additional heater on top of both 120A-assemblies and air blower in front of them (minimum power).

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-07 - 22:44

Temperatures, humidity and condensation status in Sector 7-8, Mon 7.

Position T [°C] T_meteo [°C] RH_meteo [%] T_portable [°C] RH_portable [%] Condensation Dew Point [°C] Time
DFBMC   20.6 48.0 20.4 50.8     20.45
Q6_L8 (120A) 9     19.4 52.2 Yes 9.3  
Q8_L8 (120A) 9     19.4 53.1 Yes 9.7 19.45
Q9_L8 (120A) 9.5     19.5 52.7 Yes 9.5  
Q10_L8 (120A) 9     19.0 53.7 Yes 9.4  
Q11_L8 (60A) 9.5     19.0 53.9 Yes 9.4  
Q12_L8 (60A) 9.5 18.6 51.4 18.3 55.5 Low 9.2 20.00
Q13_L8 (60A) 10     18.5 55.0 Yes 9.3  
Q14_L8 (60A) 9.5     18.4 55.4 Yes 9.3  
Q15_L8 (60A) 8.5     18.2 56.0 Yes 9.2 20.15

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-05-07 - 19:34

RAT 7/5/2007

PDF RAT_20070507_last.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x07x19x23/RAT_20070507_last.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-07 - 19:12

Programme for Tuesday 08.05.2007 in Sector 78

Schedule for Thursday 08.05.2007 in sector 7-8

GO ON WITH THE VALIDATION OF THE SEQUENCER FOR THE PROCEDURE Q4Q5

DFBAO RQ6.L8B1
RQ6.L8B2
Switch on the flashing lights around the chimneys of RQ6.L8 circuit in the DFBAO HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. AB/PO
Sequencer tests on powering procedures for 6kA magnets with 2 PCs ("Q4Q5" procedure) AB/CO

INTERVENTION OF THE GEOMETERS AT D2 Q4 (Beginning AT 9:00)

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL), EITHER IN THE DFBMA OR DFBMC, BEGIN WITH THE POWERING TESTS IN ML8

DFBMA RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
RQ4.L8B1

Balisage with flashing lights HCC
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
PC Configuration (test of PC and magnet at minimum current) AB/PO
RQ4.L8 Warn the QPS team that RQ4 will be powered HCC
RQ4.L8 Unlock the quench heaters power supplies of Q5 and charge AT/MEL
RQ4.L8 Unlock and switch on the corresponding power converters AB/PO
RQ4.L8 PC configuration AB/PO
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
RQ4.L8B1
PIC2 HCC
AB/CO
AB/PO
AT/MEL
AT/ACR
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
RQ4.L8B1
P2N
(Up to PL1.31B)
HCC
AB/CO
AB/PO
AT/MEL
AT/ACR

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
RQ5.L8
Balisage with flashing lights HCC
RCBCH5.L8B1
RCBCV5.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
PC Configuration (test of PC and magnet at minimum current) AB/PO
RQ5.L8 Warn the QPS team that RQ5 will be powered HCC
RQ5.L8 Unlock the quench heaters power supplies of Q5 and charge AT/MEL
RQ5.L8 Unlock and switch on the corresponding power converters AB/PO
RQ5.L8 PC configuration AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
RQ5.L8
PIC2 HCC
AB/CO
AB/PO
AT/MEL
AT/ACR
RCBCH5.L8B1
RCBCV5.L8B2
RQ5.L8
P2N
(Up to PL1.31B)
HCC
AB/CO
AB/PO
AT/MEL
AT/ACR

WHEN CONDITIONS IN ARC ARE MET GO ON WITH THE PROGRAMME FOR 60A CONVERTERS (ORBIT CORRECTORS) AND 120A (Q7 TO Q10 CORRECTORS) AS WELL AS THE ELQAs IN THE DFBAs

X: previous day - X : last day - O : Proposal

  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

Done on Monday 07.05.2007

SEQUENCER TESTS FOR THE PROCEDURE Q4Q5

DFBAO RQ6.L8B1
RQ6.L8B2
Balisage around the chimneys HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. Check of the short-circuit and the insulation. AB/PO
Sequencer tests on powering procedures for 6kA magnets with two power converters ("Q4Q5" procedure) AB/CO
Lock and switch off the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters AB/PO

INSULATION TEST AT THE DFBMA : RESULT OK

CIRCUIT 60A RCBV12.L8B1 INTERVENTION BY AT/MEL: COPPER COIL PLACED AROUND A CURRENT LEAD

DRY BAGS AND VENTILATORS INTALLATION AT THE DFBMC AND DFBMA BY AT/ACR

“BAD QUALITY OF THE HELIUM LIQUID IN THE LINE C PREVENTED THE POWERING IN ML8 IN THE AFTERNOON”

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-04 - 23:08

Temperatures, humidity and condensation status in Sector 7-8, Fri 4.

Position T [°C] T_meteo [°C] RH_meteo [%] T_portable [°C] RH_portable [%] Condensation Dew Point [°C] Air Speed [m/s] Time
DFBMC   20.4 48.1         0.3-0.8 17.30
    20.4 48.3 20.1 51.1       19.30
Q6_L8 (120A) 8.5     19.7 53.1 Low 9.9    
Q8_L8 (120A) 9     19.8 53.2 Low 10.0    
Q9_L8 (120A) 10     19.3 54.4 No 10.0 1.1-1.4  
Q10_L8 (120A) 9     19.0 54.2 Low 9.5   18.50
Q11_L8 (60A) 9     19.0 54.6 No 9.5    
Q12_L8 (60A) 9 18.5 52 18.6 54.5 Yes 9.2 1.0-1.3  
Q13_L8 (60A) 10     18.4 55.2 No 9.2    
Q14_L8 (60A) 9     18.5 55.1 Yes 9.3    
Q15_L8 (60A) 8.5     18.2 55.9 Yes 9.2   19.15

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-04 - 18:35

Programme for Monday 07.05.2007

Programme for Monday 07.05.2007

DFBAO RQ6.L8B1
RQ6.L8B2
Switch on the flashing lights around the chimneys of RQ6.L8 circuit in the DFBAO HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. AB/PO
Sequencer tests on powering procedures for 6kA magnets with 2 PCs ("Q4Q5" procedure) AB/CO

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL)

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Balisage with flashing lights HCC
PC Configuration (test of PC and magnet at minimum current) AB/PO
RQ5.L8 Unlock of the quench heater power supplies (DQHDS) of Q5 and charge. AT/MEL
PC Configuration AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
RQ5.L8
PIC2 HCC
AB/CO
AB/PO
AT/MEL
AT/ACR
P2N HCC
AB/CO
AB/PO
AT/MEL
AT/ACR

WHEN CONDITIONS IN ARC ARE MET
go on with the programme for the 60A converters (orbit correctors) and 120A (Q7 to Q10 correctors) as well as ElQAs in DFBAs

X: previous day - X : last day - O : Proposal

  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

Done on Friday 04.05.2007

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Connection of the DC cables to the power converters for these circuits (before 9:00) AB/PO
Verification of the connections at the DFBMC side AT/MEL
DFBAO RQ6.L8B1
RQ6.L8B2
Switch on the flashing lights around the chimneys of RQ6.L8 circuit in the DFBAO HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. AB/PO
Sequencer tests on powering procedures for 6kA magnets with 2 PCs ("Q4Q5" procedure) AB/CO

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-05-04 - 16:35

RAT meeting 4/5/2007

PowerPoint Current_lead_operating_conditions_20070504.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x04x16x33/Current_lead_operating_conditions_20070504.ppt
Presentation by A. Ballarino - conditions for powering
PDF RAT_20070504.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x04x16x33/RAT_20070504.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-04 - 16:03

Cooldown Sector 78

Despite the long series of faults (network, cooling water, pumping group) experienced at the end of last week, the cryo team managed to bring down the average temperature of the magnets to 3 K during the week-end. The switch over to the new cryoplant took place on Monday but the connection of the new pumping group was delayed until Wednesday morning. Stabilization of temperatures and pressures as well as fine adjustments of the plants took place most of Wednesday until Thursday at 11.30 when the pumping group was set at 15 mbar. Unfortunately, at 11.30 an attempt to restart the UPS system on the surface caused a power cut which stopped all the PLCs. The cryo team is still recovering from that failure and we expect to be at nominal cryogenic condition in the arc tonight Friday May 4th.

Following a top management meeting which took place on Monday, a series of measures towards the clearing of the difficulties with the commissioning of the DFBs were taken: two different ways to avoid condensation on the chimneys are now in place: blowers on 6kA leads + heaters on the 120A leads on the DFBMC and dry-air bags around the chimneys of the DFBMA.

Although the powering front was very quiet during the week, significant progress was achieved with the preparation of the sequencer for the independently powered quadrupoles and the 120A correctors: we now have the sequencer tuned for the 60 A and 120 A correctors as well as the independently powered quadrupoles and D2.

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-03 - 23:46

Anti-Condensation Systems tested on DFBMA and DFBMC

JPEG DSC03835b.jpg:
DSC03835b.jpg
DFBMA: Additional heater on the 120A-8leads-assembly and plastic envelope with forced air on the 120A-4leads-assembly
JPEG DSC03830b.jpg:
DSC03830b.jpg
DFBMC: Detail of the air blowers in front of the 6 kA chimneys/leads
JPEG DSC03828.JPG:
DSC03828.JPG
DFBMC: Additional heaters on the 120A chimneys and air blowers in front of the 6kA chimneys/leads
JPEG DSC03840b.jpg:
DSC03840b.jpg
DFBMA: Plastic envelopes with forced air on the 6 kA

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-03 - 22:55

Temperatures, relative humidity and condensation status in Sector 7-8, Thu 3.

Position T [°C] T_meteo [°C] RH_meteo [%] Condensation Comments
DFBMC   20.7 38.8    
Q6_L8 (120A) 9     Yes  
Q8_L8 (60A) 9.5     No  
Q9_L8 (60A) 13     No  
Q10_L8 (60A) 10     Yes  
Q11_L8 (60A) 9.5     No  
Q12_L8 (60A) 9 18.5 41.4 Yes  
Q13_L8 (60A) 10     No  
Q14_L8 (60A) 9     No  
Q15_L8 (60A) 9     Yes  
Q20_L8 (60A) 9     No  
Q25_L8 (60A) 10     No  
Q30_L8 (60A) 10     No  
Q30_R7 (60A) 9.5     Yes  
Q25_R7 (60A) 10.5     No  
Q20_R7 (60A) 10     No  
Q15_R7 (60A) 9     Yes  
Q14_R7 (60A) 9     Yes  
Q13_R7 (60A) 9.5     Yes  
Q12_R7 (60A) 9.5 17.3 49 Yes  
Q11_R7 (60A) 9     Yes  
Q10_R7 (60A) 9     Yes  
Q9_R7 (60A) 9.5     Yes  
Q8_R7 (60A) 9     Yes  
Q7_R7 (60A) 7     Yes  
DFBAN   18 45.9    

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-03 - 18:03

RAT meeting 3/5/2007

Programme for Friday 04.05.2007

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Connection of the DC cables to the power converters for these circuits (before 9:00) AB/PO
Verification of the connections at the DFBMC side AT/MEL
DFBAO RQ6.L8B1
RQ6.L8B2
Switch on the flashing lights around the chimneys of RQ6.L8 circuit in the DFBAO HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. AB/PO
Sequencer tests on powering procedures for 6kA magnets with 2 PCs ("Q4Q5" procedure) AB/CO

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL)

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Balisage with flashing lights HCC
PC Configuration (test of PC and magnet at minimum current) AB/PO
RQ5.L8 Unlock of the quench heater power supplies (DQHDS) of Q5 and charge. AT/MEL
PC Configuration AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
RQ5.L8
PIC2 HCC
AB/CO
AB/PO
AT/MEL
AT/ACR
P2N HCC
AB/CO
AB/PO
AT/MEL
AT/ACR

Done on Thursday 03.05.2007

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Connection of the DC cables to the DFBMC leads of the two 120A circuits which passed the ElQA tests AT/MEL
DFBAO RQ6.L8B1
RQ6.L8B2
Balisage around the chimneys HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. Check of the short-circuit and the insulation. AB/PO
Sequencer tests on powering procedures for 6kA magnets with two power converters ("Q4Q5" procedure) AB/CO
Lock and switch off the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters AB/PO


Unlock and switch on the RPLB.UA83.RCBCVS5.L8B1 converter for sequencer Tests 80-120 A AB/PO

Sequencer tests 80-120A AB/CO AB/CO

DFBAO Installation of dry-air bags on the DFBMA chimneys AT/ACR
DFBMC Installation of two heaters in two 120 A chimneys and of the air blowers for the 6kA leads AT/ACR

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-05-03 - 12:59

RAT meeting 3/5/2007

PDF RAT_20070503.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x03x12x58/RAT_20070503.pdf

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-03 - 12:52

Planning of the day 03.05.2007

PDF Planning_of_the_day_03.05.2007.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x03x12x49/Planning_of_the_day_03.05.2007.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-03 - 02:01

Temperatures, relative humidity and condensation status in Sector 7-8, Wed 2.

Position T [°C] T_meteo [°C] RH_meteo [%] Condensation Comments
DFBMA          
120A-4 14     No Setpoint 40K - additional heaters
120A-8 13     No Setpoint 50K - additional heaters
6kA 8 (on top of G11)   Yes No condensation on chimneys - Setpoint 50K
DFBMC   20.5 44.2    
6kA 8 (on G11)   No Setpoint 70K
120A-8 11     No Setpoint 50K - additional heaters not working
120A-4 11     No Setpoint 40K - additional heaters not working
DFBAO_LCM          
Q6_L8 (120A) 8     Yes  
DFBAO_HCM          
Q9_L8 (60A) 12     No  
Q10_L8 (60A) 11     No  
Q11_L8 (60A) 10     No  
Q12_L8 (60A) 9 18.3 49.5 Yes  
Q13_L8 (60A) 11     No  
Q14_L8 (60A) 9     Yes  
Q15_L8 (60A) 10     Yes  
Q20_L8 (60A) 9     Yes  
Q25_L8 (60A) 11     Yes  
Q30_L8 (60A) 11     Yes  
Q30_R7 (60A) 10     -  
Q25_R7 (60A) 11     -  
Q20_R7 (60A) 10     No  
Q15_R7 (60A) 9     Yes  
Q14_R7 (60A) 9     Yes  
Q13_R7 (60A) 15?     No  
Q12_R7 (60A) 10 17.3 53.4 Yes  
Q11_R7 (60A) 9     Yes  
Q10_R7 (60A) 9     Yes  
Q9_R7 (60A) 10     No  
Q8_R7 (60A) 9     Yes  
Q7_R7 (60A) 9     Yes  
DFBAN   18 49.9 Yes  
600A 10     Yes Setpoint 150K
13kA 10     Yes Setpoint 110K
DFBMH       No  
600A 11     No Setpoint 120K

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-30 - 18:25

RAT meeting 30/4/2007

PDF RAT_20070430.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x30x18x20/RAT_20070430.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-27 - 19:39

RAT meeting 27/4/2007

PDF RAT_20070427.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x27x19x39/RAT_20070427.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-04-26 - 23:08

Powering Tests - Completed Activities on 26/04/07 and Circuits ready for next steps

X: previous day - X : last day - O : Proposal

  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

Other activities:

  • PNO3 60A on cells RCBH12_L8B2/RCBV12_L8B1 (manual) – executed by A.Cantone 12:15
  • ELQA 60 A cells 20 and 22 (10:00) & DFBAO
  • Transfer function Measurement at Nominal current on RCBH11_L8B1 (local test in the tunnel 10:00)
  • Calibration of the Cryo-valves (A.Bouillot)
  • Condensation appears on Q6/Q7/Q8 120 A Current leads
  • Connection of the 120 A PC cell 8. Tests not performed due to condensation.

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-26 - 12:27

RAT meeting 26/4/2007

PDF RAT_20070426.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x26x12x25/RAT_20070426.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-04-25 - 19:00

Powering Tests-Completed Activities in 25/04/07 and Program for 26/04/07

PDF 25th-26th_April_2007.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x25x18x38/25th-26th_April_2007.pdf

Comments:

Attached file contains the tasks completed on 25th April 2007 and Proposed Programme for 26th April 2007

Mirko Pojer and Blanca Perea for the HCC- 25 Apr 2007

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-25 - 18:14

RAT meeting 25/4/2007

PDF RAT_20070425.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x25x18x13/RAT_20070425.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-04-25 - 10:11

Sets of Circuits for 25/04/07

Morning

PCC Tests (AB/PO) - 10A

  RCBH13.L8B1
  RCBV13.L8B2
  RCBV14.L8B1
  RCBH14.L8B2
  RCBH15.L8B1
  RCBH16.L8B2
  RCBV15.L8B2
  RCBV16.L8B1

PNO.1

  RCBH13.L8B1
  RCBV13.L8B2
  RCBV14.L8B1
  RCBH14.L8B2

Afternoon
PCC Tests (AB/PO) - 10A

  RCBCH9.L8B1
  RCBCV9.L8B2
  RCBCH10.L8B2
  RCBCV10.L8B1
  
PNO.2

  RCBH11.L8B1
  RCBV11.L8B2 
  RCBV12.L8B1 
  RCBH12.L8B2

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-24 - 17:06

RAT meeting 24/4/2007

PDF RAT_20070424.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x24x17x05/RAT_20070424.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-23 - 17:38

RAT meeting 23/4/2007

PDF RAT_20070423.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x23x17x37/RAT_20070423.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-20 - 12:49

RAT meeting 20/4/2007

PDF RAT_20070420.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x20x12x49/RAT_20070420.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-18 - 19:10

RAT meeting 18/4/2007

PDF RAT_20070418.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x18x19x08/RAT_20070418.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-16 - 17:35

RAT meeting 16/4/2007

PDF RAT_20070416.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x16x17x33/RAT_20070416.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-13 - 16:15

RAT meeting 13/4/2007

PowerPoint ICC_070413.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x13x16x11/ICC_070413.ppt

PDF RAT_20070413.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x13x16x11/RAT_20070413.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-12 - 15:05

RAT meeting 12/4/2007

PDF RAT_20070412.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x12x15x04/RAT_20070412.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-11 - 17:34

RAT meeting 11/4/2007

PDF RAT_20070411.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x11x17x34/RAT_20070411.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-10 - 12:13

RAT meeting 10/4/2007

PDF RAT_20070410.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x10x12x11/RAT_20070410.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-05 - 12:42

RAT meeting 5/4/2007

PDF RAT_20070405.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x05x12x42/RAT_20070405.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-04 - 16:32

RAT meeting 4/4/2007

PDF RAT_20070404.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x04x16x31/RAT_20070404.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-03 - 16:33

RAT meeting 3/4/2007

PDF RAT_20070403.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x03x16x29/RAT_20070403.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-04-02 - 19:02

RAT meeting 2/4/2007

PDF RAT_20070402.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x04x02x19x02/RAT_20070402.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-03-30 - 16:24

RAT meeting, 30/03/2007

PDF RAT_20070330.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x30x16x23/RAT_20070330.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-03-28 - 16:40

RAT meeting, 28/3/2007

PDF RAT_20070328.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x28x16x39/RAT_20070328.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-27 - 16:16

RAT meeting 27/3/2007

PDF RAT_20070327.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x27x16x14/RAT_20070327.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-26 - 17:34

RAT meeting 26/3/2007

PDF RAT_20070326.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x26x17x20/RAT_20070326.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-23 - 15:46

RAT meeting 23/3/2007

PDF RAT_20070323.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x23x15x45/RAT_20070323.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-22 - 18:21

RAT meeting 22/3/2007

PDF RAT_20070322.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x22x18x07/RAT_20070322.pdf

PowerPoint ELQA_RAT_22-03-2007.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x22x18x07/ELQA_RAT_22-03-2007.ppt

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-21 - 18:43

RAT meeting 21/3/2007

PDF RAT_20070321.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x21x18x41/RAT_20070321.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-19 - 17:55

RAT meeting 19/3/2007

PDF RAT_20070319.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x19x17x48/RAT_20070319.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-16 - 17:14

RAT meeting 16/3/2007

PDF RAT_20070316.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x16x17x14/RAT_20070316.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-03-15 - 10:13

Electrical problem on dipole 1055 installed in sector 7-8 and recommendations on powering

EDMS document nr. 828962

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-14 - 19:32

RAT meeting 14/3/2007

PDF RAT_20070314.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x14x19x30/RAT_20070314.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-12 - 18:07

RAT meeting 12/3/2007

PDF RAT_20070312.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x12x18x06/RAT_20070312.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-09 - 18:54

RAT meeting 9/3/2007

PDF RAT_20070309.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x09x18x53/RAT_20070309.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-08 - 17:18

RAT minutes 8/3/2007

PDF RAT_20070308.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x08x17x17/RAT_20070308.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-07 - 15:47

RAT meeting 7/3/2007

PDF RAT_20070307.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x07x15x46/RAT_20070307.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-05 - 18:42

RAT meeting 5/3/2007

PDF RAT_20070305.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x05x18x41/RAT_20070305.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-02 - 17:45

RAT meeting point 8 - 2/3/2007

PDF RAT_20070302.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x02x17x44/RAT_20070302.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-03-01 - 10:01

RAT minutes 28/2/2007

PowerPoint sector78-vacinterlock26-2-07.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x01x09x59/sector78-vacinterlock26-2-07.ppt

PDF RAT_20070228.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x03x01x09x59/RAT_20070228.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-02-26 - 18:11

RAT meeting Point 8 26/2/2007

PDF RAT_20070226.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x02x26x18x09/RAT_20070226.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-02-21 - 18:37

RAT minutes 21 Feb 2007

Microsoft Word file RAT_20070221.doc:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x02x21x18x23/RAT_20070221.doc
RAT meeting minutes

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-02-19 - 17:05

Cool-Down Sector 7-8 - 19/02/2007

PDF RAT_20070219.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x02x19x17x04/RAT_20070219.pdf

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-02-15 - 12:39

RAT in point 8 - Minutes of 15 February 2007

PDF RAT_20070215.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x02x15x12x37/RAT_20070215.pdf

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-02-14 - 19:21

RAT point 8 - Minutes 14 February

PDF RAT_20070214.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x02x14x19x19/RAT_20070214.pdf

Comments:

  • Cooldown at 75K

  • D2 repair ongoing (safety measures ensured today)

  • Warm magnets tests finishing tomorrow

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-02-13 - 12:43

RAT point 8 minutes - 13 February 2007

PDF RAT_20070213.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x02x13x12x43/RAT_20070213.pdf

Warning:

  • Strong mechanical danger in LSS8L in the D2 magnet (DCUM 23180 - DCUM 23195)

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-02-12 - 14:44

RAT Point 8 - 12 February 2007

Main activities ongoing (check meeting minutes for further details):

  • LHCCOOL cards needed to access sector 78

  • Cooldown of Sec78 from 80K to 20K planned for this week

  • Cryo-leak being repaired in magnet 3002. The problem appeared also in other magnets (SSS349).

  • Powering Procedure Test from Wednesday 14th

  • Warm Magnets powering tests ongoing this week in LSS8L an LSS8R

PDF RAT_20070212_WithPPT_planning.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x02x12x14x24/RAT_20070212_WithPPT_planning.pdf

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-02-09 - 17:59

RAT Point 8 - 09/02/2007

PDF RAT_20070209.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x02x09x17x58/RAT_20070209.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-02-07 - 19:44

RAT in Point 8 of 7th February - Minutes

PDF RAT_20070207.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x02x07x19x43/RAT_20070207.pdf

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-02-06 - 15:52

RAT Point 8 - Preparation for Powering 06-02-2007

PDF RAT_20070206.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x02x06x15x51/RAT_20070206.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-02-05 - 18:34

Minutes - sector 45 - 5th February 2007

PDF RAT_20070205.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x02x05x18x32/RAT_20070205.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-02-05 - 15:03

RAT Point 8 Minutes 05/02/07

PDF RAT_20070205.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x02x05x15x01/RAT_20070205.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-02-01 - 17:52

RAT Minutes Point 8 01-Feb

PDF RAT_20070201.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x02x01x17x51/RAT_20070201.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-01-30 - 14:19

RAT in Point 8 - Tuesday 30/01

PDF RAT_20070130.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x30x14x18/RAT_20070130.pdf

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-29 - 17:06

RAT in Point 8 - Monday Jan 29

PDF RAT_20070129_withAttachments.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x29x17x02/RAT_20070129_withAttachments.pdf

  • The short to ground in Sector 78 disappeared last Saturday
  • The cooling capacity was restored to 80% today
  • The mobile camera run will be planned tomorrow from Point 7
  • Preparation for Powering Tests will take place from tomorrow in UA83

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-26 - 12:27

RAT in point 8 - 26th January 2007

PDF RAT_20070126.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x26x12x26/RAT_20070126.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-01-25 - 18:49

Preparation for the Powering Tests Sector 7-8

PDF RAT_78_070124.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x25x18x48/RAT_78_070124.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-01-25 - 11:47

Preparation for Cool-Down Sectors 7-8,4-5,8-1

PDF RAT_20070125_2.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x25x11x45/RAT_20070125_2.pdf

Comments:

Edit, attach file or add comment


Created by Main.JeanPierreMalodDognin on 2007-01-24 - 16:56

la mise en froid se deroule tres normalement...

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-23 - 14:31

RAT in point 8 - 23th January 2007

PDF RAT_20070123.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x23x14x25/RAT_20070123.pdf

Comments:

  • NEXT RAT will be on thursday 25th january 2007 at 8:30 -- Main.BorisBellesia - 23 Jan 2007 - 14:31

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-01-22 - 17:17

RAT in Point 8 - 22nd January 2007

PDF RAT_20070122.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x22x17x13/RAT_20070122.pdf

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-19 - 19:37

RAT in Point 8 - Minutes - 19th Jan 2007

PDF RAT_20070119.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x19x19x35/RAT_20070119.pdf

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-18 - 11:14

Cool-Down follow up - Jan 18th - Plots

PDF ELQA-HVQ_cooldown_2007-01-18.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x18x10x46/ELQA-HVQ_cooldown_2007-01-18.pdf

PDF Cooling_LHC_sector78_18Jan07.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x18x10x46/Cooling_LHC_sector78_18Jan07.pdf

Comments:

  • Inlet set temperature = 185K
  • Cooling Power = 300kW = 50% of nominal
  • Leakage Current Monitoring = ok

MircoCoccoli - 18 Jan 2007 - 10:51

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-17 - 17:24

RAT in point 8 - 17th January - Minutes

PDF RAT_20070117_with_NC_Mag_Planning.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x17x17x17/RAT_20070117_with_NC_Mag_Planning.pdf

> ATTENTION: next RAT in point 8 on FRIDAY 19th at 8.30am <

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-17 - 16:43

Minutes of the meeting on Point 8 Warm Magnets Commissioning - 15th Jan

PDF Minutes_WarmMag_LSS8_15Jan07.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x17x15x56/Minutes_WarmMag_LSS8_15Jan07.pdf

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-16 - 18:33

RAT in point 8 - 16th January - Minutes

PDF RAT_20070116.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x16x18x25/RAT_20070116.pdf

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 12.01.07 Cryo-instrumentation cables: 2 LSS8L connectors (A.Suraci-J-C.Guillaume)
  • 10.01.07 Continuity error with cable 1813004A and a wrong housing of 1702440 which should be replaced (D.Bozzini)
Cryogenics
  • 16.01.07 Vac-Cryo Interlock tests – window of 1 day needed (P.Cruikshank-F.Millet)
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 08.01.07 Re-installation of 3 BLMs interfering with pumping groups (B.Dehning)
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-01-16 - 10:17

First Cooldown QRL C to D

Images obtained from Francois Millet

PNG Cooldown.png:
Cooldown.png
First cooldown QRL 78
PNG First_cooldown_Q9R7.png:
First_cooldown_Q9R7.png
First cooldown Q9R7

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-15 - 18:42

RAT Meeting in Point 8 - Jan 15th

PDF RAT_20070115.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x15x18x38/RAT_20070115.pdf

MircoCoccoli - 15 Jan 2007 - 18:40

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 12.01.07 Cryo-instrumentation cables: 2 LSS8L connectors (A.Suraci-J-C.Guillaume)
  • 10.01.07 Continuity error with cable 1813004A and a wrong housing of 1702440 which should be replaced (D.Bozzini)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 08.01.07 Re-installation of 3 BLMs interfering with pumping groups (B.Dehning)
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-12 - 18:46

RAT Meeting in Point 8 - Jan 12th

PDF RAT_20070112.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x12x18x43/RAT_20070112.pdf

MircoCoccoli - 12 Jan 2007 - 18:45

Open Issues (Sect 78):

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 10.01.07 Connection of the NC Magnets in LSS8R (J-C.Guillaume)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 12.01.07 Cryo-instrumentation cables: 2 LSS8L connectors (A.Suraci-J-C.Guillaume)
  • 10.01.07 Continuity error with cable 1813004A and a wrong housing of 1702440 which should be replaced (D.Bozzini)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
DFB Commissioning
  • 29.11.06 DFBX Pumping of Current Leads insulation vacuum (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 08.01.07 Re-installation of 3 BLMs interfering with pumping groups (B.Dehning)
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2007-01-11 - 18:27

RAT Meeting P8 - 11th Jan Minutes

PDF RAT_20070111.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x11x18x26/RAT_20070111.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2007-01-10 - 15:13

RAT Meeting P8 - 10th Jan Minutes

PDF RAT_20070110.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x10x15x12/RAT_20070110.pdf

Comments:

Edit, attach file or add comment


Created by Main.IvanMoyaMartinez on 2007-01-09 - 11:05

Labels of the Water cooled cables following the ECR- EDMS 804188

Some pictures of the water-cooled cables with the labels showing they are following the ECR- EDMS 804188, when it's necessary

JPEG P1000209.JPG:
P1000209.JPG

JPEG P1000206.JPG:
P1000206.JPG

JPEG P1000207.JPG:
P1000207.JPG

JPEG P1000208.JPG:
P1000208.JPG

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-08 - 20:35

RAT Point 8 Minutes

PDF RAT_20070108.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x08x20x33/RAT_20070108.pdf

RAT regarding Cool Down of Sect 78 and preparation for Cool Down of sectors 45 and 81

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-20 - 12:01

RAT Preparation Cool Down sect 78 - Dec 20th

PDF RAT_20061220.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x11x10/RAT_20061220.pdf

Next RAT meeting Tuesday, January 9th 8:30 @ P8 2889-R-009

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 15.12.06 DFBXG non-conformity on the current lead heater cables (P.Chambouvet-J-C.Guillaume)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
DFB Commissioning
  • 29.11.06 DFBX Pumping of Current Leads insulation vacuum (S.Feher)
13 kA EE Systems
  • 16.12.06 Calibration of the Arc Detectors on the 13 kA EE System (K.Dahlerup-Petersen)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

MircoCoccoli - 20 Dec 2006 - 11:11

Edit, attach file or add comment


Created by Main.RobertoSaban on 2006-12-20 - 09:52

Incident Summary Report - Short on the MB circuit during TP4B

During the Electrical Quality Assurance (TP4B) of sector 78 a short to ground was identified. The short chronological report of the events is attached.

PDF short_21R7.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x09x48/short_21R7.pdf

PowerPoint FSavary.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x09x48/FSavary.ppt
Report to MARIC of the repair sequence
PowerPoint DBozzini.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x09x48/DBozzini.ppt
Report to MARIC of the ELQA tests

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-19 - 12:00

Preparation Cool-Down Sector 78-RAT 19th Dec

PDF RAT_20061219.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x19x11x13/RAT_20061219.pdf

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 15.12.06 DFBXG non-conformity on the current lead heater cables (P.Chambouvet-J-C.Guillaume)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
DFB Commissioning
  • 29.11.06 DFBX Pumping of Current Leads insulation vacuum (S.Feher)
13 kA EE Systems
  • 16.12.06 Calibration of the Arc Detectors on the 13 kA EE System (K.Dahlerup-Petersen)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

MircoCoccoli - 19 Dec 2006 - 11:56

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-12-18 - 12:28

Preparation Cool-Down Sector 78-RAT 18thDec

PDF RAT_20061218.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x18x12x27/RAT_20061218.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-12-15 - 12:15

Preparation Cool-Down Sector 78-RAT 15th Nov

PDF RAT_20061215.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x15x12x14/RAT_20061215.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-12-14 - 18:25

Preparation Cool-Down Sector 78-RAT 14th Nov

PDF RAT_20061214.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x14x18x20/RAT_20061214.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-14 - 15:15

60A converter tests - Thu. 14th Dec.

Advancement of today:

  • connection of the half arc L8 finished: ready for ELQA.doc
  • TS-EL, resolution of the problem on the short circuit insulation in R7, no problems for co activities.

Tomorrow activities:

  • TS-EL, still on resolution of the problem on the short circuit insulation in R7, no problems for co activities.
  • If it is possible: El_QA.doc 120A circuits in RR77 and UA83, no problems for co activities.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-13 - 20:28

60A converter tests - Wed. 13th Dec.

JPEG P1000122.JPG:
P1000122.JPG

Advancement of today:

  • Pre-connection of the cable of the current leads: on C28-C32_L8
  • All 60A cables disconnected PC side:
  • Problem on the short circuit insulation on the 60A cables in R7 (see pic)

Tomorrow activities:

  • TS-EL, resolution of the problem on the short circuit insulation in R7, no problems for coactivities.
  • if it is possible: El_QA.doc 120A circuits in RR77 and UA83, no problems for coactivities.

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-13 - 17:25

RAT Preparation Cool Down sec78 - Dec 13

PDF RAT_20061213.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x13x17x22/RAT_20061213.pdf

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 12.12.06 Short-to-ground in QBBI.A21R7 (Lyra side) (D.Bozzini, F.Seyvet)
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum
  • 12.12.06 One Pressure Gauge in DFBs cannot be read (F.Millet)
Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 12.12.06 Repositioning of 1 DC cable on DFBAO and of 2 DV cables on DFBAN (all 13kA) (J.C.Guillaume)
  • 08.12.06 DFBXG 100V test on current leads (S.Feher)
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

MircoCoccoli - 13 Dec 2006 - 17:22

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-12 - 20:28

60A converter tests - Tuesday 12th Dec.

Advancement Tue. 12 Dec.:

  • AB-PO activities ended in the whole arc
  • Polarity tests ended in the whole arc

Activities Wed. 13th Dec.:

  • TS-EL pre-conneciton in C32-34_L8 / C 12-34_R7; no problems for coactivities
  • if it is possible: El_QA.doc in C18-28 _L8; co activities limited - a trolley is used

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-12 - 12:40

Preparation for Cool Down RAT sec78 - Dec 12

PDF RAT_20061212.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x12x12x38/RAT_20061212.pdf

Comments:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum
  • 12.12.06 One Pressure Gauge in DFBs cannot be read (F.Millet)
Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 08.12.06 DFBXG 100V test on current leads (S.Feher)
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

MircoCoccoli - 12 Dec 2006 - 12:38

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-11 - 16:08

Preparation for Cool Down RAT - Dec 11

PDF RAT_20061211.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x11x16x04/RAT_20061211.pdf

Comments:

AC non-conformity

  • 07.12.06 Non conformity in DFBMH, DFBAN and DFBXG current lead heaters (P.Chambouvet-J-C.Guillaume)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

Main.MircoCoccoli - 11 Dec 2006 - 16:07

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-08 - 17:54

60A converter tests - Friday 7th Dec.

Advancement of today:

  • AB-PO activities on cells C24-30_R7
  • 4 hr run in C12_R7-C22_R7 and C32-C34_R7 and C28-C32_L8 simultaneously (shortened version of the 24 hr run)
  • AB-PO main activity ended today;
  • Polarity tests on cells C24-26_R7

Monday activities:

  • AB-PO - tests on faulty cells: - C34_L8 - 1 converter in C32_R7 - 1 converter in C12_R7 - no problems for coactivities.
  • Polarity and conneciton in C28-32_L8 / C30-34_R7; no problems for coactivities
  • ELQA in C18-26_L8; co activities limited - a trolley is used

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-08 - 10:27

Preparation for Cool Down RAT - Dec 8

PDF RAT_20061208.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x08x10x17/RAT_20061208.pdf

IMPORTANT: PLEASE COMMENT ON THE OPEN ISSUES IF YOU FIND THEM INCORRECT, INCOMPLETE OR NOT-PERFECTLY UPDATED.

AC non-conformity

  • 07.12.06 Non conformity in DFBMH, DFBAN and DFBXG current lead heaters (P.Chambouvet-J-C.Guillaume)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling

  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum -

Cryogenics

  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

MircoCoccoli - 08 Dec 2006 - 10:17

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-07 - 19:02

60A converter tests - Thursday 7th Dec.

Comments:

Advancement of today:

  • AB-PO activities on cells C12_R7 - C16-22_R7
  • Polarity tests on 120A converter alimenting cells C6-7-8-9-10_R7
  • Disconnection of the circuits of the 120A converter alimenting cells C6-7-8-9-10_R7
  • MME activities postponed to Monday

Tomorrow activities:

  • AB-PO activities on cells C24-30_R7
  • 4 hr run in C12_R7-C22_R7 and C32-C34_R7 and C28-C32_L8 simultaneously (shortened version of the 24 hr run)
  • ELQA tests - whole day / no co-activities

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-07 - 12:37

Preparation for Cool Down sec78 - Dec 7

PDF RAT_20061207.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x07x12x29/RAT_20061207.pdf

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 30.11.06 Non conformity: 4 temperature sensor cables need repair (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC (Flohe)
Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
DFB Commissioning
  • 04.12.06 Re-positioning collars used to fix the WRL to the CLs (all DFBs) (A.Perin-D.Bozzini-F.Millet)
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)

MircoCoccoli - 07 Dec 2006 - 12:36

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-06 - 15:17

Preparation for Cool Down sec78 - Dec 6

PDF RAT_20061206_wslides_wList.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x06x15x14/RAT_20061206_wslides_wList.pdf

Comments:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC (Flohe)
Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
DFB Commissioning
  • 04.12.06 Re-positioning collars used to fix the WRL to the CLs (all DFBs) (A.Perin-D.Bozzini-F.Millet)
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)

Main.MircoCoccoli - 06 Dec 2006 - 15:16

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-05 - 13:50

Preparation for Cool Down sec78 - 5th Dec 06

PDF RAT_20061205_wslides.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x05x13x47/RAT_20061205_wslides.pdf
Minutes with Presentation by D.Bozzini attached

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 3 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 04.12.06 Access & Safety Cable Installation (full sector) (J-C.Guillaume; end by Dec 7th)
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 04.12.06 Re-positioning collars used to fix the WRL to the CLs (all DFBs) (A.Perin-D.Bozzini-F.Millet)
  • 29.11.06 Pumping of insulation vacuum for DFBX
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

-- Main.MircoCoccoli - 05 Dec 2006 - 13:47

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-04 - 19:03

Preparation for Cool Down sec78 - RAT 4th Dec

PDF RAT_20061204.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x04x17x39/RAT_20061204.pdf

PDF Slide_w49.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x04x17x39/Slide_w49.pdf
Week 49 Usage Table

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 3 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 04.12.06 Access & Safety Cable Installation (full sector) (J-C.Guillaume; end by Dec 7th)
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 04.12.06 Re-positioning collars used to fix the WRL to the CLs (all DFBs) (A.Perin-D.Bozzini-F.Millet)
  • 29.11.06 Pumping of insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

-- Main.MircoCoccoli - 04 Dec 2006 - 17:40

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-03 - 23:20

Sector 78 - Week 49 - Zones Usage

This tentative planning will be discussed at Monday 4th RAT meeting in Point 8.

Thanks, Mirco

PDF Plan78w49.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x03x23x13/Plan78w49.pdf

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-01 - 13:16

RAT - Preparation for Cool Down sec 78 - 1 Dec 06

PDF RAT_20061201.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x01x13x13/RAT_20061201.pdf

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 3 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (A. Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected (A.Suraci: 80% done)
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 29.11.06 Definition of owner of activity and pumping needs for current leads insulation vacuum (DFBX)
  • 29.11.06 Pumping of insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

-- Main.MircoCoccoli - 01 Dec 2006 - 13:13

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-11-30 - 20:07

preparation for cool down sec 78 - Nov. 30

Microsoft Excel Spreadsheet Sector_78_Planning_60A.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x30x20x01/Sector_78_Planning_60A.xls

PDF RAT_20061130.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x30x20x01/RAT_20061130.pdf

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 3 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (A. Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected (A.Suraci: 80% done)
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 29.11.06 Definition of owner of activity and pumping needs for current leads insulation vacuum (DFBX)
  • 29.11.06 Pumping of insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX) -- boris.bellesia - 30 Nov 2006 - 20:05

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-11-29 - 18:50

Preparation for Cool Down sec.78 - Nov 29

PDF RAT_20061129.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x29x18x25/RAT_20061129.pdf

Mpp Sector_78_Preparation_of_Cool_Down_w48.mpp:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x29x18x25/Sector_78_Preparation_of_Cool_Down_w48.mpp
UPDATED PLANNING (comments are welcome!)

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 3 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected (A.Suraci: 80% done)
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 29.11.06 Definition of owner of activity and pumping needs for current leads insulation vacuum (DFBX)
  • 29.11.06 Pumping of insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 22.11.06 Pretest of 60A converters in sector 78, location C16L8. (B.Bellesia: stopped due to LQA hardware problems; expected end by Tuesday 28 Nov)
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-11-28 - 16:40

Preparation for Cool Down sec. 78 - Nov 28

PDF RAT_20061128.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x28x16x37/RAT_20061128.pdf

-- Main.MircoCoccoli - 28 Nov 2006 - 16:38

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 4 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG, DFBMH
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected (A.Suraci: 80% done)
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: (done: DFBAO; next(w48): DFBAN, DFBMA, DFBMC, DFBMH, DFBX))
  • 17.11.06 Pumping of insulation vacuum for all DFBs between 7 and 8 (P.Chambouvet: DFBAN done; next: DFBMA, DFBMC, DFBX)
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 22.11.06 Pretest of 60A converters in sector 78, location C16L8. (B.Bellesia: stopped due to LQA hardware problems; expected end by Tuesday 28 Nov)
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-11-27 - 19:08

RAT - Preparation for Cool Down sec.78 - Nov 27

PDF RAT_20061127.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x27x19x02/RAT_20061127.pdf

Main.MircoCoccoli - 27 Nov 2006 - 19:02

Open Issues:

AC non-conformity

  • 21.11.06 Transport and installation of the heater system of the current leads for RR77 and UA83 (P.Chambouvet: about 2 weeks of work starting Mon 27 Nov)
  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 4 cables are in repair)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 25.11.06 Functional test on the leads heaters: DFBMA, DFBMC, DFBX
  • 25.11.06 120A and 600A positioning on DFBAN and DFBMH
  • 23.11.06 Cable temporary positioning on the transport side of the tunnel (A.Vergara: next year)
  • 22.11.06 Pre-connection of 120A cables in LSSR7
  • 20.11.06 Galvanic insulation installation on all DFBs
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected (D.Bozzini: 80% done)
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: ongoing according to schedule (next: DFBMA, DFBMC, DFBX); in week 48 will begin ElQA-TP4-B, i.e. HV tests)
  • 17.11.06 Pumping of insulation vacuum for all DFBs between 7 and 8 (P.Chambouvet: DFBAN done; next: DFBMA, DFBMC, DFBX)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J.C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 23.11.06 Planning of consignations/deconsignations
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 22.11.06 Pretest of 60A converters in sector 78, location C16L8. (B.Bellesia: stopped due to LQA hardware problems; expected end by Tuesday 28 Nov)
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-11-24 - 20:37

Preparation Cool-Down Sector 78-RAT 24th Nov

PDF RAT_20061124.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x24x20x25/RAT_20061124.pdf

Comments: Open Issues:

  • AC non-conformity
    • 21.11.06 Transport and installation of the heater system of the current leads for RR77 and UA83 (P.Chambouvet: about 2 weeks of work starting Mon 27 Nov)
    • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 4 cables are in repair)
    • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
    • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
  • DC cabling
    • 23.11.06 Cable temporary positioning on the transport side of the tunnel (A.Vergara: next year)
    • 22.11.06 Pre-connection of 120A cables in LSSR7
    • 20.11.06 Galvanic insulation installation on all DFBs
    • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
    • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
  • Assembly -
  • Vacuum -
  • Cryogenics
    • 22.11.06 Cryo-valves remote control in CCC by AB/CO
    • 22.11.06 Control of quench valve
    • 20.11.06 DFB instrumentation cables to be connected (D.Bozzini: 80% done)
    • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
  • DFB Commissioning
    • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: ongoing according to schedule; next week will begin ElQA-TP4-B, i.e. HV tests)
    • 17.11.06 Pumping of insulation vacuum for all DFBs between 7 and 8 (P.Chambouvet: DFBAN done)
  • Safety
    • General remark: written communication in advance to announce operations!!!
    • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
    • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J.C.Guillaume and approval by J.Etheridge.
    • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
    • 23.11.06 Planning of consignations/deconsignations
    • 16.11.06 Cool down safety procedure & access conditions
  • Preparation of powering test
    • 22.11.06 Pretest of 60A converters in sector 78, location C16L8. (B.Bellesia: stopped due to LQA hardware problems; expected end by Tuesday 28 Nov)
    • 22.11.06 Installation of flexible cables by CV (UA83)
    • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
  • AOB
    • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
    • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

-- Main.BlancaPereaSolano - 24 Nov 2006 - 20:29

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-11-23 - 15:23

Preparation for Cool-Down Sec78 - Thu Nov, 23rd

PDF RAT_20061123.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x23x14x53/RAT_20061123.pdf

Comments:

    • AC non-conformity
  • 21.11.06 Transport and installation of the heater system of the current leads for RR77 and UA83
  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.

    • DC cabling
  • 23.11.06 Cable temporary positioning on the transport side of the tunnel
  • 23.11.06 Adjustment of cable length on DFBAO
  • 22.11.06 Missing labels on DFBAO instrumentation cables
  • 22.11.06 Pre-connection of 120A cables in LSSR7
  • 20.11.06 Galvanic insulation installation of 600 A cables to the DFBAO
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
    • Assembly: -

    • Vacuum: -

    • Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected
  • 17.11.06 Instrumentation cable HV tests in LSS8L

    • DFB Commissioning
  • 13.11.06 ElQA-TP4-A in the DFBAO, DFBAN, DFBAX
  • 17.11.06 Pumping of insulation vacuum DFBAN, DFBAO, DFBMC, DFBMA, DFBXG

    • Safety: -

    • General remark: written communication in advance to announce operations!!!
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype plastic protection or the monorail (protection derived from a plastic tube) by J.C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 23.11.06 Planning of consignations/deconsignations
  • 16.11.06 Cool down safety procedure & access conditions

    • Preparation of powering test
  • 22.11.06 Pretest of 60A converters in sector 78, location C16L8.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.

    • AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

-- Main.MircoCoccoli - 23 Nov 2006 - 14:55

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-11-22 - 16:19

Preparation for Cool-Down of Sector 78, 22 Nov

PDF RAT_20061122.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x22x16x14/RAT_20061122.pdf

Comments:

AC non-conformity

  • 21.11.06 Transport and installation of the heater system of the current leads for RR77 and UA83
  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
  • 20.11.06 Galvanic insulation installation of 600 A cables to the DFBAO
  • 22.11.06 Pre-connection of 120A cables in LSSR7
  • 22.11.06 Missing labels on DFBAO instrumentation cables
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected
  • 17.11.06 Instrumentation cable HV tests in LSS8L
DFB Commissioning
  • 17.11.06 Prototype of galvanic insulation to be tested at Point 6
  • 13.11.06 ElQA-TP4-A in the DFBAO, DFBAN, DFBAX
  • 17.11.06 Pumping of insulation vacuum DFBAN, DFBAO, DFBMC, DFBMA, DFBXG
Safety General remark: written communication in advance to announce operations!!!
  • 16.11.06 Cool down safety procedure & access conditions
  • 16.11.06 Definition of Safety procedure for Powering Procedure Test (aka Dry Runs)
Preparation of powering test
  • 22.11.06 Pretest of 60A converters in sector 78, location C16L8.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-11-23 - 12:20

Preparation for Cool-Down RAT - Tuesday, November 21st

PDF RAT_20061121.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x21x12x15/RAT_20061121.pdf

Open Issues

  • AC non-conformity
    • 21.11.06 Transport and installation of the heater system of the current leads for RR77 and UA83
    • 21.11.06 Non-conformity on instrumentation cables (LT, PT, Level Gauges, …)
    • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current)
    • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
    • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • DC cabling
    • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
    • 17.11.06 600 A cable positioning on DFBA
    • 20.11.06 Connections (made by TS/MME) of 120 A cables to be verified (TS/EL)
    • 20.11.06 Connections of 600 A cables to the DFBAO to be verified
  • Assembly

  • Vacuum
    • 20.11.06 21 leak detectors to be placed all along the arc 78
  • Cryogenics
    • 20.11.06 DFB instrumentation cables to be connected
    • 20.11.06 Leaks on DFBAO gas recovery line in repair/installation of helicoflex
    • 17.11.06 Instrumentation cable HV tests in LSS8L
  • DFB Commissioning
    • 17.11.06 Prototype of galvanic insulation to be tested at Point 6
    • 13.11.06 ElQA -TP4-A in the DFBAO, DFBAN, DFBAX
    • 17.11.06 Pumping of insulation vacuum DFBAN, DFBAO, DFBMC, DFBMA, DFBXG
  • Safety
  • General remark: written communication in advance to announce operations!!!
    • 16.11.06 Cool down safety procedure & access conditions
    • 16.11.06 Definition of Safety procedure for Powering Procedure Test (aka Dry Runs)
  • Preparation of powering test
    • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be cleaned!
  • AOB
    • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
    • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-11-20 - 19:55

Preparation cool-down of Sector 78, 20 Nov

PDF Sector_78_Preparation_of_Cool_Down_w47.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x20x19x42/Sector_78_Preparation_of_Cool_Down_w47.pdf

PDF RAT_20061120.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x20x19x42/RAT_20061120.pdf

Mpp Sector_78_Preparation_of_Cool_Down_w47.mpp:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x20x19x42/Sector_78_Preparation_of_Cool_Down_w47.mpp

Comments:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converetrs & VAC.
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
DC cabling
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
  • 13.11.06 120 A cable connections in DFBAO, DFBMC, DFBMA and DFBX LSS8L
  • 17.11.06 600 A cable positioning on DFBA
  • 20.11.06 Connections (made by TS/MME) of 120 A cables to be verified (TS/EL)
  • 20.11.06 Connections of 600 A cables to the DFBAO to be verified
  • 20.11.06 Flohe on DFBAO or DFBAM? To be decided

Assembly - Vacuum

  • 20.11.06 21 leak detectors to be placed all along the arc 78
  • 17.11.06 Bellows damaged in Q6-Q7 LSS8L
Cryogenics
  • 20.11.06 DFB instrumentation cables to be connected
  • 20.11.06 leaks on DFBAO in reparation
  • 17.11.06 Instrumentation cable HV tests in LSS8L
  • 17.11.06 Pressurized air supply to valve on WRL
  • 13.11.06 Support on DFBAO WRL
DFB Commissioning
  • 17.11.06 Prototype of galvanic insulation to be tested at Point 6
  • 13.11.06 ElQA-TP4-A in the DFBAO, DFBAN, DFBAX
  • 17.11.06 Pumping of insulation vacuum will start next week in 7R
Safety General remark: written communication in advance to announce operations!!!
  • 16.11.06 Cool down safety procedure & access conditions
  • 16.11.06 Definition of Safety procedure for Powering Procedure Test (aka Dry Runs)
  • 17.11.06 Monorail in Point 8 locked off on Monday (20.11) starting from 10.30!!!
Preparation of powering test
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be cleaned!
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX) -- Main.bellesia - 20 Nov 2006 - 19:53

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-11-17 - 16:54

Preparation of cool-down in sector 78, 17 Nov

PDF RAT_20061117.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x17x16x48/RAT_20061117.pdf

Comments:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converetrs & VAC.
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
DC cabling
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
  • 13.11.06 120 A cable connections in DFBAO, DFBMC, DFBMA and DFBX LSS8L
  • 17.11.06 600 A cable positioning on DFBA
Assembly -

Vacuum

  • 17.11.06 Leak in arc (Q11-8L)
  • 17.11.06 Bellows damaged in Q6-Q7 LSS8L
Cryogenics
  • 17.11.06 Instrumentation cable HV tests in LSS8L
  • 17.11.06 Cryogenic valve manipulation on DFBX
  • 17.11.06 Pressurized air supply to valve on WRL
  • 13.11.06 Support on DFBAO WRL
DFB Commissioning
  • 17.11.06 Prototype of galvanic insulation to be tested at Point 6
  • 13.11.06 ElQA-TP4-A in the DFBAO, DFBAN, DFBAX
  • 17.11.06 Pumping of insulation vacuum will start next week in 7R
Safety General remark: written communication in advance to announce operations!!!
  • 16.11.06 Cool down safety procedure & access conditions
  • 16.11.06 Definition of Safety procedure for Powering Procedure Test (aka Dry Runs)
  • 17.11.06 Monorail in Point 8 locked off on Monday starting from 10.30!!!
Preparation of powering test
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be cleaned!
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-11-16 - 17:10

Preparation of Cool-Down in Sector 78, 16 Nov

PDF RAT_20061116.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x16x17x00/RAT_20061116.pdf

Comments:

  • Open Issues
    • 13.11.06 Cryo. Valves manipulation on the DFBMH, DFBAN and DFBAO
    • 13.11.06 Non-conformity of the mini-crates of cryo instrumentation
    • 13.11.06 Cabling non-conformity affecting QPS and 60A Orbit Correctors IST
    • 13.11.06 WRL DFBAO missing support LSS8L
    • 13.11.06 3 x 600 A power cable positioning LSS8L
    • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
    • 13.11.06 ElQA-TP4-A in the DFBAO, DFBAN, DFBAX
    • 13.11.06 LSS8L Q6-Q7 bake-out
    • 13.11.06 120 A Cables connection in DFBMC, DFBMA and DFBX LSS8L
    • 13.11.06 6 kA and 13 kA cables positioning at DFBMA and DFBMC
    • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)
    • 13.11.06 DS7R and DS8L leak test
    • 13.11.06 Leak in R7
    • 13.11.06 Galvanic Insulation 600 A, 6 kA, 13 kA for LSS8L and LSS7R available
    • 15.11.06 Leak in arc
    • 15.11.06 Leak test to be performed after the replacement of bellows

-- Main.MirkoPojer - 16 Nov 2006 - 17:52

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-11-15 - 16:05

Preparation cool-down of Sector 78, 15 Nov

PDF RAT_20061115.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x15x14x56/RAT_20061115.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-11-14 - 17:45

Preparation of the Cool-Down in Sector 78, 14 Nov.

PDF RAT_20061114.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x14x17x42/RAT_20061114.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-11-13 - 17:55

Preparation Cool-Down Sector 78-RAT 13th Nov

PDF RAT_20061113.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x13x17x54/RAT_20061113.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-10-30 - 17:06

RAT Minutes 30-10-2006

PDF RAT_061030.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x10x30x17x05/RAT_061030.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-10-27 - 15:06

RAT Minutes 27-10-2006

PDF RAT_061027.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x10x27x14x28/RAT_061027.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-10-23 - 18:45

Status of LSS8

The final Access system is being installed in P8. We could see the personnel and material SAS in the UA (Angels and Demons' eye recognition !!). The bake-out of the kickers in LSS8R is taking place, and should be at nominal temperature by wednesday ! The TDI is still not in place, though. We show here also the new vacuum valves.

JPEG Kickers-Bake-Out2-P8.JPG:
Kickers-Bake-Out2-P8.JPG
Kickers being baked-out
JPEG SAS-Eye.JPG:
SAS-Eye.JPG
SAS in UA87
JPEG VacuumValves-P8.JPG:
VacuumValves-P8.JPG
Vacuum Valve

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-10-16 - 10:58

Long Straight Section in P8

Friday 13th October I visited the Long Straight Section in P8 up to the TI8 injection line. Here I post some photos of the status of the installation and particularly about the bake-out preparation of the MKIs. You can see how TI8 enters at the MSI level (beautiful !!). Details about collimators, and BLM are also shown. The TDI was not down yet !

JPEG IMG_2010.JPG:
IMG_2010.JPG
Bake-out preparation for the Kickers in the LHC ring
JPEG IMG_2008.JPG:
IMG_2008.JPG
TI8 entering the ring at the level of MSIs
JPEG IMG_2009.JPG:
IMG_2009.JPG
BLM
JPEG IMG_2006.JPG:
IMG_2006.JPG
TI8 warm magnets running behind the QRL
JPEG IMG_2007.JPG:
IMG_2007.JPG
TI8 MCIs and MQIs

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-05-29 - 15:13

First results for the 24 hrs run in UA87

First comments from the field for the 24 hrs run in UA87. Minutes of the last RAT meeting...

Regs

Rosario

PDF RAT_20060529.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x05x29x15x10/RAT_20060529.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-05-24 - 12:00

24hrs test ongoing...

Extension of the 24 hrs test decided this morning...

Last news in the RAT minutes.

Rosario

PDF RAT_20060524.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x05x24x11x58/RAT_20060524.pdf

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-05-24 - 08:21

fond decran

il faut aimer le vert...

JPEG P5230077.JPG:
P5230077.JPG

JPEG P5230079.JPG:
P5230079.JPG

JPEG P5230076.JPG:
P5230076.JPG

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-05-23 - 11:13

Les medecins au chevet de RB..in UA87...

Les transformateurs siamois du RB accusaient , hier, une difference de leurs temperatures de 20 degres (detectee hier lors du test de debit de courant), handicap au tests de 24 h programes pour 14h ce 23 mai. Hugues et son equipe de medecins se relaient a son chevet pour en determiner la cause... Un des transformateurs du RB aurait -il contracte un rechauffement/refroidissement, au point d en faire une maladie benigne?

JPEG blindagemobileUJ86230506.JPG:
blindagemobileUJ86230506.JPG

JPEG RBUA87230506.JPG:
RBUA87230506.JPG

JPEG laissez-vousguider.JPG:
laissez-vousguider.JPG

JPEG suivezlesflechesenUA86.JPG:
suivezlesflechesenUA86.JPG

JPEG UJ86230506.JPG:
UJ86230506.JPG

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-05-22 - 12:04

Ready for the 24 hrs test tomorrow..

Here you are, RAT fans... The minutes of the last meeting.

And remember tomorrow is the G day, 24 hrs run in UA87.

PDF RAT_20060522.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x05x22x12x01/RAT_20060522.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-05-19 - 15:02

Run CCC

Last RAT minutes from CCC...

PDF RAT_20060519.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x05x19x15x01/RAT_20060519.pdf

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-05-19 - 13:53

Tests 8h 19 mai UA87 et instantanes en images..

JPEG mesuresIRdescablesDCUA87.JPG:
mesuresIRdescablesDCUA87.JPG

JPEG RQ6UA87.jpg:
RQ6UA87.jpg

JPEG RQDsoushtesurveillance.JPG:
RQDsoushtesurveillance.JPG

JPEG P5180071.JPG:
P5180071.JPG

Comments:

-- Main.JeanPierreMalodDognin - 19 May 2006

Julien Kiss en pleine mesure des temperatures des cables DC..par infra rouge

-- Main.JeanPierreMalodDognin - 19 May 2006

-- Main.JeanPierreMalodDognin - 19 May 2006

RQ6 en plein effort debitant son/ses/miliers d ampere(s) vers 17h...

-- Main.JeanPierreMalodDognin - 19 May 2006

RQD sous controle ...comme aux urgences...

-- Main.JeanPierreMalodDognin - 19 May 2006

-- Main.JeanPierreMalodDognin - 19 May 2006

Less rafraichissements sont servis....

-- Main.JeanPierreMalodDognin - 19 May 2006

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-05-18 - 15:24

Finally ready for 8hrs run, 120, 600 and >2kA

News from the field...

The Rats

PDF RAT_20060518.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x05x18x15x22/RAT_20060518.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-05-16 - 11:15

400kV affects the schedule of the tests in UA87

Minutes of the RAT meeting....

PDF RAT_20060516.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x05x16x11x13/RAT_20060516.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-05-15 - 15:33

Test CCC cancelled today...

Please find enclose the minutes of the RAT meeting.

Unfortunately the schedule has been a bit reorganised in the light of the 400kV trip.

Regards

Rosario

PDF RAT_20060515.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x05x15x15x30/RAT_20060515.pdf

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-05-11 - 15:22

Mesures de courant DC en UA87

les parties qui ont chauffeees ont change de couleur..mais lesquelles ont chauffe exagerement..?

JPEG mesuresdeIDCUA87.JPG:
mesuresdeIDCUA87.JPG

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-05-08 - 15:03

Minutes of RAT

Voila le minutes du meeting quotidien

PDF RAT_20060508.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x05x08x15x01/RAT_20060508.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-05-04 - 16:15

8hrs run today for 120 et 600A

8hrs run today for 120 et 600A, news in the minutes of meeting.

Rosario

PDF RAT_20060504.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x05x04x16x12/RAT_20060504.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-05-02 - 11:53

Tunnel RATs news....

Here you are the daily news from the front line.

RAT minutes of meeting Mai 2, 2006.

Rosario

PDF RAT_20060502.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x05x02x11x51/RAT_20060502.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-04-27 - 14:24

Tunnel RATs as usual

Here you are your daily news from the UA87.

PDF RAT_20060427.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x04x27x14x22/RAT_20060427.pdf

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-04-27 - 14:07

on respire en RA87 ou....

..au-dessus de la QRL ,les boites a cables etaient venues .. se reposer tout contre les vannes cryogeniques qu 'on apercoit sur la photo. Fort heureusement aucun degat n 'est a deplorer et tout est rentre dans l 'ordre apres que Jean-Claude Guillaune ait fait corriger le probleme en faisant raccourcir les corgages nylon qui soutiennent l 'ensemble cables en court-circuit et boites isolantes , cordages maintenant cet ensemble a bonne distance des fragiles vannes en question qui peuvent desormais vaquer a leurs occupations de vannes... Tout va etre mis en oeuvre pour que ce probleme ne se renouvelle pas.

JPEG RA87boitesacablesCC.jpg:
RA87boitesacablesCC.jpg

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-04-26 - 17:01

501 ieme aimant connecte

Comme on peut le constater sur cette photo le 501ieme aimant est enfin connecte a la ligne THT !! Souhaitons que son fonctionnement soit irreprochable ...

JPEG 501iemeaimantconnecte.jpg:
501iemeaimantconnecte.jpg

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-04-26 - 16:59

Fresh news from the tunnel RATs

Minutes of the daily meeting in point 8.

PDF RAT_20060426.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x04x26x16x55/RAT_20060426.pdf

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-04-26 - 15:37

DFB posee en UJ86...

Elle est deja imposante...une fois raccordee elle le sera encore plus. JPEG DFBXUJ86.JPG:
DFBXUJ86.JPG

JPEG DFBUA87.JPG:
DFBUA87.JPG

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-04-25 - 10:57

Tunnel RATs news April 25

April 25, fete de la liberation pour les Italiens. Here you are last news from UA87.

Regards

PDF RAT_20060425.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x04x25x10x55/RAT_20060425.pdf

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-04-24 - 14:33

nouvelles du front en images...

Par manque de temps... quelques images jetees un peu en vrac ...la suite de la moisson d images un peu plus tard peut etre dans la soiree ,ou l apres midi. JPEG lemeilleurdesmondesUA87.JPG:
lemeilleurdesmondesUA87.JPG

JPEG vuedessusRQDRQFUA.JPG:
vuedessusRQDRQFUA.JPG

JPEG RBdedosUA87.JPG:
RBdedosUA87.JPG

JPEG IMGP4730.JPG:
IMGP4730.JPG

JPEG connectionsdeRB.JPG:
connectionsdeRB.JPG

JPEG RQDRQFbisUA87.JPG:
RQDRQFbisUA87.JPG

JPEG IMGP4747.JPG:
IMGP4747.JPG

JPEG imposantesconnectionsRQDRQFUA87.JPG:
imposantesconnectionsRQDRQFUA87.JPG

JPEG deshistoiresdRQD.JPG:
deshistoiresdRQD.JPG

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-04-24 - 11:48

Last from tunnel RATs April 24

Minutes from point 8...

PDF RAT_20060424.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x04x24x11x46/RAT_20060424.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-04-21 - 11:07

News from the tunnel RATs, April 21.

Usual minutes of meeting in point 8.

Regards

Rosario PDF RAT_20060421.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x04x21x11x06/RAT_20060421.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-04-20 - 18:24

RAT news April 20, 2006

Hi tunnel RATs, here you are last news from the field.

Rosario PDF RAT_20060420.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x04x20x18x22/RAT_20060420.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-04-19 - 17:30

RAT meeting Wed April 19

As usual the RAT meeting took place in point 8 this morning. Main issues: safety and coactivities (kickers, DFBX)

Regards

Rosario

Comments:

PDF RAT_20060419.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x04x19x17x27/RAT_20060419.pdf

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-04-18 - 18:06

News from the point 8. RAT meeting minutes.

Here you are the last RAT meeting minutes. Have a look if something concerns you..

Regards

Rosario

Comments:

PDF RAT_20060418.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x04x18x18x03/RAT_20060418.pdf

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-04-13 - 11:55

Photos fraiches du 13 avril en UA87 et RA87

Sans aucun commentaire autre que ...celui des images .

Comments:

JPEG RBUA87.JPG:
RBUA87.JPG

JPEG RQDA81UA87.JPG:
RQDA81UA87.JPG

JPEG IMGP4381.JPG:
IMGP4381.JPG

JPEG pourquisontcesserpentsenRA87.JPG:
pourquisontcesserpentsenRA87.JPG

JPEG RA87encouleurs.JPG:
RA87encouleurs.JPG

JPEG lamachineprendformeRA87.JPG:
lamachineprendformeRA87.JPG

JPEG posetiroir600A.JPG:
posetiroir600A.JPG

JPEG DQSQUA87.JPG:
DQSQUA87.JPG

JPEG lescouleursenUA87.JPG:
lescouleursenUA87.JPG

JPEG DQRCenUA87.JPG:
DQRCenUA87.JPG

JPEG RQFRQDUA87.JPG:
RQFRQDUA87.JPG

JPEG lapieuvreenRA87.JPG:
lapieuvreenRA87.JPG

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-04-12 - 17:29

Start up of the RAT activity in UA87

Dear all, Heare we are, next tuesday, April 18, 8h30 as usual, officially start the short-circuit tests for the commissioning of the UA87. A preparatory meeting took place in point 8 this morning. Minutes in attachment.

Rosario

Comments:

PDF RAT_20060412.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x04x12x17x22/RAT_20060412.pdf

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-04-10 - 23:12

tests HT des cables refoidis en UA43: OK

Les problèmes de fuites ayant été resolus par CV qui a travaillé "ferme" ce week-end,les tests HT des cables de puissance(avec eau de refroidissement) ont pu se dérouler comme prévu ce soir 10 avril( avec bien sur la bonne qualité de l 'eau en ce qui concerne sa résistivité). CV a signalé également que l'équilibrage sera revu quand les branchements "eau" sur les convertisseurs de type RYM seront réalisés. Les mesures d 'isolement s 'étalent de 20 (environ) à 100 Mohm dans le meilleur des cas pour les cables refoidis. Un élément important à corriger (photo): un court-circuit avec la masse sur les piquages desservant RA / RB a été mis en évidence pendant ces tests: raccord eau touchant le caisson de protection métallique des tubes cuivre de puissance. Une solution Mac Gyver TRES(!)provisoire a été trouvée (ligature fibre de verre d'emballage...) pour vérifier que ce court-circuit était bien le seul sur cette ligne ....

Une photo montre le foisonnement des câbles sortant des RYM et également que les dessus des racks n 'ont pour l' instant pas reçu leurs doses de copeaux métalliques comme ce fut le cas en UA 47, on croise les doigts....

...Et demain est un autre jour....

Comments:

JPEG UA87RBccaveclecaissonbis.jpg:
UA87RBccaveclecaissonbis.jpg

JPEG UA87RYMsorties100406.JPG:
UA87RYMsorties100406.JPG

JPEG RARBUA87.jpg:
RARBUA87.jpg

JPEG sortiesRYMUA87.JPG:
sortiesRYMUA87.JPG

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-04-07 - 17:32

Acqua Alta en UA87

Des problèmes de serrage, de joints plats et donc d étanchéite sur les connections de sortie des convertisseurs RQ7....RQ10 en UA87 ont tenté de plonger momentanément l 'UA87 dans une période de hautes eaux.... (photo) ce vendredi 7 avril, alors qu' on mettait la dernière main à la réalisation des connections de puissance sur DQRB.......

Cependant que Noel (Fournier) et des techniciens, à l'abri dans l'enceinte insonorisée de DQSB poursuivaient leur travaux, et continuaient de poser les sondes de températures et voltmètriques pour les connections de puissance (photo)...

Mais les tests HT sur les cables - initialement prevus lundi soir- vont peut-ètre devoir ètre reportés au mardi soir,.et le personnel CV devra s'affairer ce week-end afin que toutes les raccords "eau" soient parfaits...

Comments:

JPEG Noel070406.JPG:
Noel070406.JPG

JPEG UA87070406.JPG:
UA87070406.JPG

JPEG AquaaltaUA87070406.JPG:
AquaaltaUA87070406.JPG

JPEG UA87DQSB.JPG:
UA87DQSB.JPG

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-04-06 - 16:44

Next RAT meeting

Dear all,
Please remeber next RAT meeting in point 8 will be held on Wednesday, April 12, 2006, at 8h30.
Regards.
Rosario

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-04-06 - 15:01

DFB...DFB....

elle attend sagement d etre enfin connectee , toute brodee d or qu elle semble etre..(un peu de poesie ne peut pas faire de mal?)

Comments:

JPEG DFBRJ84.JPG:
DFBRJ84.JPG

JPEG sortiesRYMUA87.JPG:
sortiesRYMUA87.JPG

Edit, attach file or add comment


Created by Main.RobertoSaban on 2006-03-13 - 13:37

The validation of the cryoplant in Point 8

Les tests de validation du cryoplant du Point 8 (stockages helium et azote, refrigerateurs 4.5 K, unites de refrigeration 1.8 K et boite d'interconnection) démarreront ce mercredi 15 mars 2006. La fin des tests est programmee le 30 mai 2006.

En prevision des tests de Hardware Commissioning des secteurs 7-8 et 8-1, ces runs cryoplant P8 permettront de valider les dernieres consolidations mecaniques apportees au systeme cryogenique et ils permettront aussi de valider les differents modes d'operation normaux et degrades de l'ensemble de nos equipements cryogeniques presents au P8.

Pour informer le personnel circulant au P8, le message d'information ci-joint sera affiche sur les differents points d'acces des batiments cryo de surface et des zones cryo en caverne. Tout commentaire de fond et/ou de forme sur la note d'information jointe est le bienvenu.

A votre disposition pour tout complement, meilleures salutations,

Francois Millet Resp. operation cryo P8

* Message_P8cryorun_spring2006.ppt: Sign to be posted in Point 8

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-07-10 - 15:17

Warm Magnets Commissioning of Point 7 - Minutes from meeting_10/07/07

PDF RAT_20070710_WM7.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x07x10x15x13/RAT_20070710_WM7.pdf

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-06-22 - 15:48

Warm Magnets Commissioning at point 7. Minutes 22-06-2007

PDF WARM_MAGNETS_7.v03.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x06x22x15x40/WARM_MAGNETS_7.v03.pdf

PDF RAT_20070622_WM7.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x06x22x15x40/RAT_20070622_WM7.pdf

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-06-18 - 12:34

Warm Magnets Commissioning at point 7. Minutes 18-06-2007

PDF RAT_20070618_WM7.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x06x18x12x33/RAT_20070618_WM7.pdf

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-06-11 - 19:06

Warm Magnets Commissioning at point 7. Minutes 11-06-07

PDF RAT_20070611_WM7.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x06x11x19x01/RAT_20070611_WM7.pdf

PDF PLANNING_WM7.v02.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x06x11x19x01/PLANNING_WM7.v02.pdf

Comments:

Please, find in the minutes actions affecting the planning that have come out after the meeting.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-03-01 - 15:49

SCT in RR73 and UJ76 - Mar. 1st

PDF RAT_20070301.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x03x01x15x39/RAT_20070301.pdf

********************************************

* All converters in RR73 and UJ76 VALIDATED during 24hr heat run

********************************************

07.02 DONE Water circuit filters cleaning in RR73
12.02 DONE ERD1 powering, FGCs configuration and first tests in RR73
13.02 DONE 120A (RR73) powering tests
14.02 DONE 600A (RR73) powering tests
14.02 DONE Full power ventilation test
15.02 DONE 600A-EE interlock and powering tests
15.02 DONE 600A-EE interlock tests
20.02 DONE 600A (UJ76) powering tests
21.02 DONE Water circuit filters cleaning in UJ76
22.02 DONE Polarity test
23.02 DONE 8h heat run in RR73 and UJ76
26.02 DONE Water circuit filters cleaning in RR73/UJ76
27.02 DONE CCC test for 24h run
28.02 DONE 24h heat run test
01.03 DONE Electrical services-Inom-Imax
01.03 DONE EE-QPS test at Imax
02.03 - 13kA EE tests
02.03 - AUG&UPS

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-02-28 - 15:13

SCT in RR73 and UJ76 Feb. 28th

PDF RAT_20070228.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x02x28x15x10/RAT_20070228.pdf

07.02 DONE Water circuit filters cleaning in RR73
12.02 DONE ERD1 powering, FGCs configuration and first tests in RR73
13.02 DONE 120A (RR73) powering tests
14.02 DONE 600A (RR73) powering tests
14.02 DONE Full power ventilation test
15.02 DONE 600A-EE interlock and powering tests
15.02 DONE 600A-EE interlock tests
20.02 DONE 600A (UJ76) powering tests
21.02 DONE Water circuit filters cleaning in UJ76
22.02 DONE Polarity test
23.02 DONE 8h heat run in RR73 and UJ76
26.02 DONE Water circuit filters cleaning in RR73/UJ76
27.02 DONE CCC test for 24h run
28.02 ON GOING 24h heat run test
01.03 - Electrical services-Inom-Imax
01.03 - EE-QPS test at Imax
01.03 - AUG&UPS

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-02-26 - 12:23

SCT in RR73 and UJ76 Feb. 26th

PDF RAT_20070226.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x02x26x12x02/RAT_20070226.pdf

  • TOMORROW 27.02: 24hr heat run preparation from CCC

  • Heat Runs and Test Schedule:

07.02 DONE Water circuit filters cleaning in RR73
12.02 DONE ERD1 powering, FGCs configuration and first tests in RR73
13.02 DONE 120A (RR73) powering tests
14.02 DONE 600A (RR73) powering tests
14.02 DONE Full power ventilation test
15.02 DONE 600A-EE interlock and powering tests
15.02 DONE 600A-EE interlock tests
20.02 DONE 600A (UJ76) powering tests
21.02 DONE Water circuit filters cleaning in UJ76
22.02 DONE Polarity test
23.02 DONE 8h heat run in RR73 and UJ76
26.02 ON GOING Water circuit filters cleaning in RR73/UJ76
27.02 - CCC test for 24h run
28.02 - 24h heat run test
01.03 - Electrical services-Inom-Imax
01.03 - EE-QPS test at Imax
01.03 - AUG&UPS

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-02-26 - 12:01

SCT in

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-02-23 - 12:04

SCT in RR73 and UJ76 - Feb. 23

PDF RAT_20070223.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x02x23x11x54/RAT_20070223.pdf

!!8hr HEAT RUN ON-GOING!!

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-02-22 - 22:28

SCT in RR73 and UJ76 Feb. 22nd

PDF RAT_20070222.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x02x22x22x24/RAT_20070222.pdf

Comments:

* 23.02 - 8hr RUN.

Heat Runs and Test Schedule

07.02 DONE Water circuit filters cleaning in RR73
12.02 DONE ERD1 powering, FGCs configuration and first tests in RR73
13.02 DONE 120A (RR73) powering tests
14.02 DONE 600A (RR73) powering tests
14.02 DONE Full power ventilation test
15.02 DONE 600A-EE interlock and powering tests
15.02 DONE 600A-EE interlock tests
20.02 DONE 600A (UJ76) powering tests
21.02 DONE Water circuit filters cleaning in UJ76
22.02 DONE Polarity test
23.02 - 8h heat run in RR73 and UJ76
26.02 - PIC interface test
26.02 - Water circuit filters cleaning in RR73/UJ76
27.02 - CCC test for 24h run
28.02 - 24h heat run test
01.03 - Electrical services-Inom-Imax
01.03 - EE-QPS test at Imax
01.03 - AUG&UPS

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-02-20 - 19:12

SCT in RR73 and UJ76, Feb. 20th

PDF RAT_20070220.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x02x20x18x57/RAT_20070220.pdf

Heat Runs and Test Schedule

07.02 DONE Water circuit filters cleaning in RR73
12.02 DONE ERD1 powering, FGCs configuration and first tests in RR73
13.02 DONE 120A (RR73) powering tests
14.02 DONE 600A (RR73) powering tests
14.02 DONE Full power ventilation test
15.02 ON GOING 600A-EE interlock and powering tests
15.02 ON GOING 600A-EE interlock tests
20.02 ON GOING 600A (UJ76) powering tests
21.02 DONE Water circuit filters cleaning in UJ76
22.02 - Polarity test
23.02 - 8h heat run in RR73 and UJ76
26.02 - PIC interface test
26.02 - Water circuit filters cleaning in RR73/UJ76
27.02 - CCC test for 24h run
28.02 - 24h heat run test
01.03 - Electrical services-Inom-Imax
01.03 - EE-QPS test at Imax
01.03 - AUG&UPS

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-02-16 - 16:18

SCT RR73 and UJ76, Feb. 16

PDF RAT_20070216.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x02x16x16x17/RAT_20070216.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-02-14 - 18:21

SCT - RR73 UJ76 - Feb. 14th

PDF RAT_20070214.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x02x14x18x20/RAT_20070214.pdf

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-06-20 - 11:24

une vue de la protection des cables 13kA

une bache armee de fibres de verre a ete enroulee en 4 epaisseurs autour des cables refroidis et offre une bonne protection mecanique et electrique pendant les tests de courant ..quelques volts seulement....

JPEG RR77cables13kA.JPG:
RR77cables13kA.JPG

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-06-15 - 13:51

Balisage RR77 du 15 juin au matin

Les tests vont pouvoir commencer puisque tout semble pret...

JPEG RR77balisee2.JPG:
RR77balisee2.JPG

JPEG RR77balisee.JPG:
RR77balisee.JPG

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-06-14 - 17:18

...RR77 le 14 juin. a 16h ..

JPEG IMGP5971.JPG:
IMGP5971.JPG

JPEG IMGP5980.JPG:
IMGP5980.JPG

JPEG IMGP5974.JPG:
IMGP5974.JPG

JPEG IMGP5977.JPG:
IMGP5977.JPG

JPEG IMGP5986.JPG:
IMGP5986.JPG

Comments:

-- Main.JeanPierreMalodDognin - 14 Jun 2006

Mise sous tension du tableau ERD -source d energie des convertisseurs- prevue pour demain 15 juin en premiere heure....

-- Main.JeanPierreMalodDognin - 14 Jun 2006

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-06-07 - 16:15

En RR77, l echeance approche...

JPEG RR77070606.JPG:
RR77070606.JPG

JPEG TRSUJ76070606.JPG:
TRSUJ76070606.JPG

JPEG RR77bis.JPG:
RR77bis.JPG

Comments:

-- Main.JeanPierreMalodDognin - 07 Jun 2006

noter la position occupee par la maquette grandeur nature de la DFB et les cables DC y aboutissant(manque la piece de court circuit)

-- Main.JeanPierreMalodDognin - 07 Jun 2006

Et bien sur pour alimenter RR77 en energie, il faut un transfo de puissance AC que l equipe EL s active a finir de tester.(UJ76)

-- Main.JeanPierreMalodDognin - 07 Jun 2006

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-05-30 - 14:58

demandez les dernieres nouvelles d RR77..demandez..

JPEG IMGP5718.JPG:
IMGP5718.JPG
recherche contorsionnistes pour connections tres difficiles voir presque impossibles en RR77, sur convertisseurs
JPEG IMGP5723.JPG:
IMGP5723.JPG
un mur sur la frontiere s est eleve.. la resistance aux envahisseurs(les radiations..) s organise, vu en arrivant en RR77 depuis LHC7
JPEG IMGP5733.JPG:
IMGP5733.JPG
vue du ciel(Arthus-Bertrand..) du 1 ier etage d RR77..Noel aux prises avec un chassis lie a DQRB.???
JPEG IMGP5721.JPG:
IMGP5721.JPG
convertisseurs tous raccordes!!!
JPEG IMGP5730.JPG:
IMGP5730.JPG
vous avez un message(a gauche au premier plan..)
JPEG LEswitchenRR77.JPG:
LEswitchenRR77.JPG
recherche appart tres insonorise pour scenes de menage repetees(faire offre sous chiffre RR77 DQS
JPEG IMGP5742.JPG:
IMGP5742.JPG
A proximite de RR77, comme deux crustaces se defient de leurs pinces..deux aimants se font face..le combat va commencer.

Comments:

ce matin vers 10h en RR77....

-- Main.JeanPierreMalodDognin - 30 May 2006

vue du 1ier etage de RR77..Noel Fournier aux prises avec un chassis lie a DQF???

-- Main.JeanPierreMalodDognin - 30 May 2006

cherche appart tres insonorise pour scenes de menage reguliere(faire offre sous chiffre RR77. DQF....)

-- Main.JeanPierreMalodDognin - 30 May 2006

Ce matin la pose des panneaux d isolation phonique se poursuivait activement a l etage de RR77...ce n est qu ensuite que seront connectes au switch les cables de laisons de puissance DC

-- Main.JeanPierreMalodDognin - 30 May 2006

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-05-12 - 11:22

Petite idee de la RR77 le10 mai 2006

JPEG P5100028.JPG:
P5100028.JPG

JPEG P5100026.JPG:
P5100026.JPG

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-04-12 - 17:34

LSS7R short-circuit tests soon

Dear all, Once UA87 will be completed, we are immediately supposed to start the short circuit tests in RR77. In less than two months. A status of the installation works in LSS7R seems adequate at this stage. Proposal: next HCC meeting, April 20 15h30 in bld. 30.

Rosario

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-03-27 - 12:14

letter from Jean-Pierre Malod-Dognin

Comme je suis allé faire un’’ tour de chauffe’’ dans la RR77 ce PM, je me permet de vous faire profiter de ces quelques photos (une fois n est pas

coutume) que j ai cru bon de faire, comme une carte postale du bout du monde (faut pas pousser quand même… )

Les choses sont moins ‘’étendues’’ que je ne le pensais a priori.. .

J ai matérialise les emplacements des convertisseurs par des flèches rouges (c est leurs seule position possible, je présume..)

Les ‘’ contacteurs ‘’ sont au premier étage et accessibles seulement par une …échelle (flèche verte)

JP Malod-Dognin

Comments:

JPEG RR771.JPG:
RR771.JPG

JPEG RR773bis.jpg:
RR773bis.jpg

JPEG RR775.JPG:
RR775.JPG

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-10-31 - 20:49

Installation and Commissioning of the Beam Dump System _ Minutes of 30/10/07 meeting

PDF RAT__20071030.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x10x31x20x48/RAT__20071030.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-10-30 - 19:39

From Pressure Test to Cool-down _ Sector 6-7 _ 30/10/2007

JPEG 6-7_october30.jpg:
6-7_october30.jpg

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-10-30 - 19:38

From Pressure Test to Cool-down _ Sector 5-6 _ 30/10/2007

JPEG 5-6_october30.jpg:
5-6_october30.jpg

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-10-17 - 23:15

From pressure test to cool-down - Sector 5-6 - 16/10/07

JPEG Sector_5-6_status_161007.jpg:
Sector_5-6_status_161007.jpg

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-10-09 - 21:28

From Pressure test to Cool-down - Sector 5-6 _ 09/10/2007

JPEG 5-6.jpg:
5-6.jpg

Feedback of the meeting:

  • Problems with the soldering of the flexible of DFBAJ - To be resolded (after meeting: soldered, waiting for radiography)
  • ElQA on DFBXF.3R5 not yet done (after meeting: possibly done on Thursday)
  • Cool-down could be delayed 1 week (start on w.46)

After meeting:

  • Partial coherence test successfully performed
  • Leak in 15L6 will be repaired on Wednesday; 3 bars will be put during night to check
  • Rail for re-routing of 120 A cables on DFBLE.RR57 is in place; connection on Wednesday

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-10-03 - 15:32

Warm Magnets Commissioning in point 6. Action list for the 24h run and polarity test

-Thursday 4th morning: Setup ongoing (Bernard Dubois) and pre-test from the CCC before 10:30h

-Thursday 4th,10:30h: Consignation of the converters in order to remove the covers from the magnets (Jacky Mazet)

-Thursday 4th,≈14:00h: Once the covers are removed, the converters will be deconsignée (Jacky Mazet)

-Thursday 4th,after deconsignation of the converters: Start of the 24h heat run from the CCC

-Thursday 4th,≈ 17:00h: General verification on the cables (Baptiste Lebeau)

-Friday 5th, before finishing the 24h heat run: Polarity test (Jacky Mazet)

-Friday 5th, ≈ 13:30h: Magnets and cables temperatures measurements (Jacky Mazet, Baptiste Lebeau)

-Friday 5th, ≈ 14:00h: Stop of the heat run from the CCC

-Friday 5th, just after stopping the heat run, temperatures measurements on the converters (Bernard Dubois)

-Friday 5th , consignation of the converters (Jacky Mazet)

-Friday 5th, after consignations of the converters: the covers will be put in place (Jacky Mazet)

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-10-03 - 15:15

Warm Magnets Commissioning Poin 6. Action list for the 24h heat run and polarity test

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-09-28 - 18:30

Warm Magnets Commissioning in point 6. Minutes 28.09.2007

To be found together with the Minutes 28.09.2007 in point 2

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-09-14 - 19:07

Beam Dump System _ Installation and Commissioning - Minutes of 13/09/2007.

PDF RAT__20070913.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x09x14x18x57/RAT__20070913.pdf

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-09-07 - 16:26

Warm Magnets Commissioning at point 6. Minutes 03-09-2007

PDF RAT_20070903_WM6.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x09x07x15x38/RAT_20070903_WM6.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-08-28 - 19:03

Beam Dump System - Installation and Commissioning - Minutes from 28/08/2007

PDF RAT_20070828.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x08x28x19x02/RAT_20070828.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-08-06 - 01:03

Beam Dump System - Installation and Commissioning - 31/07/2007

PDF RAT_20070731.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x08x06x00x59/RAT_20070731.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-06-26 - 19:00

Beam Dump System - Installation and Commissioning - 19/06/2007

PDF RAT_20070619.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x06x26x18x59/RAT_20070619.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-06-18 - 16:24

Beam Dump System - Installation & Commissioning - 05/06/2007

PDF RAT_20070605.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x06x18x16x21/RAT_20070605.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-31 - 17:54

Beam Dump System - Installation & Commissioning - 22nd May

PDF RAT_20070522.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x05x31x17x53/RAT_20070522.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-15 - 09:54

Beam Dump System Installation and Commissioning, RAT 08 May 2007

PDF RAT_20070508.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x05x15x09x53/RAT_20070508.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-04-19 - 19:12

Beam Dump System Commissioning - RAT of April 17, 2007.

Mpp LSS6_v2.mpp:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x04x19x18x50/LSS6_v2.mpp

PDF RAT_20070417.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x04x19x18x50/RAT_20070417.pdf

PLANNED INTERVENTIONS IN THE AREA

  • The bake-out of the whole MKDs-to-MKDs sector will be completed by w28.
  • A cabling campaign (including fire detection system) is foreseen for w16 to w19. No interferences have been pointed out during the inspection visit of AB/BT, TS/EL and TS/HDO.
  • A water maintenance intervention has been scheduled for the w18 to w20 (demineralised water, cryogenic services and iced water will be affected).
  • Test of the MSD foreseen for w.29-30.
  • Cool-down of sector 5-6 for mid August.
  • Cool-down of sector 6-7 for mid October.
Open Issues
  • 10.04 Installation of BPMSD AB/BI
  • 10.04 Installation of BCTFD AB/BI
  • 10.04 Modifications on PDC of MKB-67 requested by Norisko AB/BT
  • 10.04 Mounting and alignment of beam dump line in RA63 AT/VAC
  • 10.04 Installation of MKD magnets in RA63 AB/BT
  • 10.04 Installation of junction modules between MKDs in RA67 AT/VAC
  • 03.04 Inspection by Norisko on PDC of UA63
  • 03.04 Assembly of PDC for MKDs in UA63 AB/BT
  • 03.04 Assembly of PDC for MKDs in UA67 AB/BT
  • 20.03 Mismatch between TCDS and MSD AT/VAC&AB/BT
  • 13.03 Grounding of instrumentation/after deconsignation AB/BT
  • 27.02 TCS: installation on w.18
  • 27.02 Vacuum tube between carbon window and TDE (fabrication on going-ready for the end of June) AB/ATB
  • 27.02 VMTAB bellows are needed for w18 AB/BT
  • 20.02 VAC: bake-out of MKD-6R AT/VAC
  • 20.02 TDE in UD68 ready for end of June
  • 20.02 BPM installation AB/BI
  • 20.02 Vacuum pipe on Q4 AT/VAC-AB/BT
  • 20.02 Safety inspection by Norisko on PDC for MKD63 and MKB63
  • 20.02 Optical fibers for pulsing K. Loukili-TS/EL
Closed Issues
  • 17.04 VAC: bake-out of MKD-6L 20.02
  • 10.04 Beam dump line in TD62 10.04
  • 10.04 BLMs integration 13.03
  • 10.04 Smoothing of the MSD 10.04
  • 03.04 Assembly of PDC for MKBs in UA63 03.04
  • 27.03 Safety inspection by Norisko on power distr.rack (MKB67) 20.02
  • 27.03 Grounding of instrumentation/before deconsignation 13.03
  • 10.04 Mounting supports for second beam (on drawings) (6L&R) 13.03
  • 27.03 Place for storing MKDs in UX65 20.03
  • 20.03 Grounding of magnets 13.03
  • 13.03 Intermediate bellows availability 27.02
  • 13.03 Interference between vacuum support and MKD 27.02
  • 27.02 Space in UJ62 and UJ67 for installation of spare TDE (and transport equipment) - late installation of that part of dump line 20.02

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-04-13 - 15:29

Beam Dump System Commissioning - RAT of April 10, 2007.

Comments: PLANNED INTERVENTIONS IN THE AREA

  • The bake-out of the whole MKDs-to-MKDs sector will be completed by w28.
  • A cabling campaign (including fire detection system) is foreseen for w16 to w19.
  • A water maintenance intervention has been scheduled for the w18 to w20 (demineralised water, cryogenic services and iced water will be affected).
  • Test of the MSD foreseen for w.21-22.
  • Cool-down of sector 5-6 for mid August.
  • Cool-down of sector 6-7 for mid October.
Open Issues
  • 10.04 Installation of BPMSD AB/BI
  • 10.04 Installation of BCTFD AB/BI
  • 10.04 Modifications on PDC of MKB-67 requested by Norisko AB/BT
  • 10.04 Mounting and alignment of beam dump line in RA63 AT/VAC
  • 10.04 Installation of MKD magnets in RA63 AB/BT
  • 10.04 Installation of junction modules between MKDs in RA67 AT/VAC
  • 03.04 Inspection by Norisko on PDC of UA63
  • 03.04 Assembly of PDC for MKDs in UA63 AB/BT
  • 03.04 Assembly of PDC for MKDs in UA67 AB/BT
  • 20.03 Mismatch between TCDS and MSD AT/VAC&AB/BT
  • 13.03 Grounding of instrumentation/after deconsignation AB/BT
  • 27.02 TCS: installation after dump line
  • 27.02 Vacuum tube between carbon window and TDE AB/ATB
  • 27.02 VMTAB bellows are needed for w18 AB/BT
  • 20.02 VAC: bake-out of 6L AT/VAC
  • 20.02 VAC: bake-out of 6R AT/VAC
  • 20.02 TDE in UD68 ready for end of June
  • 20.02 BPM installation AB/BI
  • 20.02 Vacuum pipe on Q4 AT/VAC-AB/BT
  • 20.02 Safety inspection by Norisko on PDC for MKD63 and MKB63
  • 20.02 Optical fibers for pulsing K. Loukili-TS/EL
Closed Issues
  • 10.04 Beam dump line in TD62 10.04
  • 10.04 BLMs integration 13.03
  • 10.04 Smoothing of the MSD 10.04
  • 03.04 Assembly of PDC for MKBs in UA63 03.04
  • 27.03 Safety inspection by Norisko on power distr.rack (MKB67) 20.02
  • 27.03 Grounding of instrumentation/before deconsignation 13.03
  • 10.04 Mounting supports for second beam (on drawings) (6L&R) 13.03
  • 27.03 Place for storing MKDs in UX65 20.03
  • 20.03 Grounding of magnets 13.03
  • 13.03 Intermediate bellows availability 27.02
  • 13.03 Interference between vacuum support and MKD 27.02
  • 27.02 Space in UJ62 and UJ67 for installation of spare TDE (and transport equipment) – late installation of that part of dump line 20.02

-- Main.MirkoPojer - 13 Apr 2007

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-04-04 - 14:44

Beam Dump System Commissioning - RAT of April 3, 2007.

PDF RAT_20070403.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x04x04x14x34/RAT_20070403.pdf

PLANNED INTERVENTIONS IN THE AREA

  • The bake-out of the whole MKDs-to-MKDs sector will be completed by w28.
  • Smoothing of the MSD is foreseen for this week (w.12).
  • A cabling campaign (including fire detection system) is foreseen for w16 to w19.
  • A water maintenance intervention has been scheduled for the w18 to w20 (demineralised water, cryogenic services and iced water will be affected).
  • Test of the MSD foreseen for w.21-22.
  • Cool-down of sector 5-6 for mid August.
  • Cool-down of sector 6-7 for mid October.
Open Issues
  • 20.03 Mismatch between TCDS and MSD AT/VAC&AB/BT
  • 13.03 Mounting supports for second beam (on drawings) (6R) AB/BT
  • 13.03 BLMs integration AB/BT&ICL
  • 13.03 Grounding of instrumentation/after deconsignation AB/BT
  • 27.02 TCS: installation after dump line
  • 27.02 Vacuum tube between carbon window and TDE AB/ATB
  • 27.02 VMTAB bellows are needed for w18 AB/BT
  • 20.02 VAC: bake-out of 6L AT/VAC
  • 20.02 VAC: bake-out of 6R AT/VAC
  • 20.02 TDE in UD68 ready for end of June
  • 20.02 BPM installation AB/BI
  • 20.02 Vacuum pipe on Q4 AT/VAC-AB/BT
  • 20.02 Safety inspection by Norisko on PDC for MKD63 and MKB63
  • 20.02 Optical fibers for pulsing K. Loukili-TS/EL
Closed Issues
  • 27.03 Safety inspection by Norisko on power distr.rack (MKB67) 20.02
  • 27.03 Grounding of instrumentation/before deconsignation 13.03
  • 27.03 Mounting supports for second beam (on drawings) (6L) 13.03
  • 27.03 Place for storing MKDs in UX65 20.03
  • 20.03 Grounding of magnets 13.03
  • 13.03 Intermediate bellows availability 27.02
  • 13.03 Interference between vacuum support and MKD L. Ducimetiere 27.02
  • 27.02 Space in UJ62 and UJ67 for installation of spare TDE (and transport equipment) – late installation of that part of dump line 20.02

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-03-28 - 16:25

Beam Dump System Commissioning - RAT of March 27, 2007.

PDF RAT_20070327.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x03x28x14x18/RAT_20070327.pdf

PLANNED INTERVENTIONS IN THE AREA

  • The bake-out of the whole MKDs-to-MKDs sector will be completed by w28.
  • Smoothing of the MSD is foreseen for this week (w.12).
  • A cabling campaign (including fire detection system) is foreseen for w16 to w19.
  • A water maintenance intervention has been scheduled for the w18 to w20 (demineralised water, cryogenic services and iced water will be affected).
  • Test of the MSD foreseen for w.21-22.
  • Cool-down of sector 5-6 for mid August.
  • Cool-down of sector 6-7 for mid October.
Open Issues
  • 20.03 Mismatch between TCDS and MSD AT/VAC&AB/BT
  • 13.03 Mounting supports for second beam (on drawings) (6R) AB/BT
  • 13.03 BLMs integration AB/BT&ICL
  • 13.03 Grounding of instrumentation/after deconsignation AB/BT
  • 27.02 TCS: installation after dump line
  • 27.02 Vacuum tube between carbon window and TDE AB/ATB
  • 27.02 VMTAB bellows are needed for w18 AB/BT
  • 20.02 VAC: bake-out of 6L by w-13 AT/VAC
  • 20.02 VAC: bake-out of 6R by w-15 AT/VAC
  • 20.02 TDE in UD68 ready for end of June
  • 20.02 BPM installation AB/BI
  • 20.02 Vacuum pipe on Q4 AT/VAC-AB/BT
  • 20.02 Safety inspection by Norisko on power distr. Racks for MKDs
  • 20.02 Optical fibers for pulsing K. Loukili-TS/EL
Closed Issues
  • 27.02 Safety inspection by Norisko on power distr.Racks (MKBs) 20.02
  • 27.03 Grounding of instrumentation/before deconsignation 13.03
  • 27.03 Mounting supports for second beam (on drawings) (6L) 13.03
  • 27.03 Place for storing MKDs in UX65 20.03
  • 20.03 Grounding of magnets 13.03
  • 13.03 Intermediate bellows availability 27.02
  • 13.03 Interference between vacuum support and MKD L. Ducimetiere 27.02
  • 27.02 Space in UJ62 and UJ67 for installation of spare TDE (and transport equipment) – late installation of that part of dump line 20.02

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-03-23 - 09:55

Beam Dump System Commissioning - RAT of March 20, 2007.

PDF RAT_20070320.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x03x23x09x40/RAT_20070320.pdf

Planned interventions in the area

  • Smoothing of the MSD is foreseen for this week (w.12).
  • A campaign for fire detection system cabling is foreseen for w14 to w17.
  • A water maintenance intervention has been scheduled for the w18 to w20 (demineralised water, cryogenic services and iced water will be affected).
  • Test of the MSD foreseen for w.21-22.
  • Cool-down of sector 5-6 for mid August.
  • Cool-down of sector 6-7 for mid October.
Open Issues
  • 20.03 Mismatch between TCDS and MSD AT/VAC&AB/BT
  • 20.03 Place for storing MKDs in UX65 AB/BT
  • 13.03 Mounting supports for second beam (on drawings) AB/BT
  • 13.03 BLMs integration AB/BT&ICL
  • 13.03 Grounding of instrumentation AB/BT
  • 27.02 TCS: installation after dump line
  • 27.02 Vacuum tube between carbon window and TDE AB/ATB
  • 27.02 VMTAB bellows are needed for w18 AB/BT
  • 20.02 VAC: bake-out of 6L by w-13 AT/VAC
  • 20.02 VAC: bake-out of 6R by w-15 AT/VAC
  • 20.02 TDE in UD68 ready for end of June
  • 20.02 BPM installation AB/BI
  • 20.02 Vacuum pipe on Q4 AT/VAC-AB/BT
  • 20.02 Safety inspection by Norisko on power distr. racks
  • 20.02 Optical fibers for pulsing (compl. W11) K. Loukili-TS/EL
Closed Issues
  • 20.03 Grounding of magnets 13.03
  • 13.03 Intermediate bellows availability 27.02
  • 13.03 Interference between vacuum support and MKD L. Ducimetiere 27.02
  • 27.02 Space in UJ62 and UJ67 for installation of spare TDE (and transport equipment) – late installation of that part of dump line 20.02

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-03-15 - 18:34

Beam Dump System Commissioning - RAT of Tuesday, 13.

PDF RAT_20070313.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x03x15x18x18/RAT_20070313.pdf

Text MTF_STEPS.txt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x03x15x18x18/MTF_STEPS.txt

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-03-08 - 18:07

Beam Dump System Commissioning - RAT of Tuesday, 6.

PDF RAT_20070306.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x03x08x18x06/RAT_20070306.pdf

PDF LSS6_v2.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x03x08x18x06/LSS6_v2.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-02-21 - 18:18

Beam Dumping System Commissioning RAT, Feb. 20

PDF RAT_20070220.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x02x21x18x16/RAT_20070220.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-02-14 - 18:23

SCT - UA67 UJ67 - OPEN ISSUES

Comments: Open Issues SCT in UA67 and UJ67

  • 13.02 Report on the AUG
  • 12.02 PM generated but stacked in local
  • 12.02 Few labels missing for cables identification
  • 05.02 Little ducts between UA and RA not closed during the 24hr
  • 02.02 RSF2.A67B1 and RSF2.A67B2 PIC cables to extend
  • 02.02 NC opening on the bad connection on the 6kA racks
  • 08.01 EE fan rotation problem – all ring

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-02-12 - 17:32

SCT in UA67/UJ67 and Preparation for SCT in RR73/UJ76, Feb 12

PDF RAT_20070212.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x02x12x17x28/RAT_20070212.pdf

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-02-08 - 20:19

SCT in UA67/UJ67 - Feb, 8

PDF SCT-planning-RR73-UJ76_v02.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x02x08x20x17/SCT-planning-RR73-UJ76_v02.pdf

PDF RAT_20070208.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x02x08x20x17/RAT_20070208.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-02-06 - 19:38

SCT UA67-UJ67 Feb. 6th

PDF RAT_20070206.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x02x06x19x37/RAT_20070206.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-02-05 - 15:30

SCT UA67-UJ67 Feb. 5th

PDF RAT_20070205.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x02x05x15x20/RAT_20070205.pdf

  • 24h PC heat run ON GOING.

Phone numbers for operators:

Hardware commissioning coordination:

  • Boris Bellesia 162483

Power converters :

  • Responsable AB/PO: David Nisbet 164223
  • (en cas de absence) Hugues Thiesen 164205
  • Numéro général AB/PO-LHC-SCT-24H: 162333

For AB/PO – in the tunnel:

  • Journée 8:00-18:00 heures: Emmanuel Garde 164035
  • Journée 17:00 -20:00 heures: Vincent Barbet 165148
  • Nuit 20:00–02:00 heures: Halil Oruc 165193
  • Nuit 02:00-8:00 heures: Okbi Zenasni 163678

Energy extraction system:

  • Reiner Denz 165289

DC cables:

  • Journée jusqu’a 12:00 heures: D.Camus 164903
  • Journée 12:00-20:00 heures: B.Lebeau 160913
  • Nuit 20:00-02:00 heures: S.Boutte 163390
  • Nuit 02.00-08.00 heures: D.Camus 164903

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-02-02 - 21:01

SCT UA67 - UJ67, Feb. 2nd

PDF RAT_20070202.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x02x02x20x50/RAT_20070202.pdf

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-31 - 20:38

SCT UA 67 - UJ 67, Jan. 31st.

PDF RAT_20070131.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x31x20x37/RAT_20070131.pdf

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-30 - 12:35

SCT_UA 67- UJ 67, Jan. 30th 2006

PDF RAT_20070130.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x30x12x14/RAT_20070130.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-29 - 12:45

SCT_UA 67- UJ 67, Jan. 29th 2006

PDF RAT_20070129.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x29x12x44/RAT_20070129.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-01-26 - 18:03

Beam Dumping System_RAT - January, 26

PDF RAT_20070126.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x26x18x01/RAT_20070126.pdf

PDF LSS6_v1.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x26x18x01/LSS6_v1.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-26 - 16:43

SCT_UA 67- UJ 67, Jan. 26th 2006

PDF RAT_20070126.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x26x16x41/RAT_20070126.pdf

Comments:

  • At 15:00 8hr run of the 13 kA circuits have been STOPPED due to water cooling circuit.
  • TS-CV intervention scheduled for Monday 29.01
  • 8hr run for the 13 kA circuits foreseen for the next Tuesday (30.01) after TS-CV intervention.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-25 - 17:05

AUG scheduling meeting

PowerPoint Tests_annuels_des_AUG_LHC.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x25x16x36/Tests_annuels_des_AUG_LHC.ppt

Present :

P. Anderssen, B. Bellesia, C. Castillo Trello, G. Cumer, C. Dehavay, P. Fessia, J.P. Malod Dognin, M. Pojer, R. Saban, E. Barbero Soto, J.Toullieux, S. Weisz


* Presentation of G. Cumer (see attached file) - The procedures adopted during an AUG have been mentioned. - Resume of the affected Surface and Underground Areas during the AUG in Point 6.

* Friday 9 Feb. can be a possible date for a AUG in point 6: Waiting for the “ok” of Installation (S. Weisz/ TS-IC) and Interconnections (P. Fessia/ AT-MCS)

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-25 - 16:31

SCT_UA 67- UJ 67, Jan. 25th 2006

PDF RAT_20070125.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x25x16x30/RAT_20070125.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-24 - 19:21

SCT_UA 67- UJ 67, Jan. 24th 2006

PDF RAT_20070124.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x24x19x21/RAT_20070124.pdf

Comments:

Edit, attach file or add comment


Created by Main.JeanPierreMalodDognin on 2007-01-19 - 16:45

Choses vues avant les tests...

JPEG IMGP9979.JPG:
IMGP9979.JPG

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-01-19 - 12:24

RAT for SCT in UA67/UJ67 - Jan, 19

PDF RAT_20070119.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x19x11x25/RAT_20070119.pdf

PDF SCT-Planning-UA67UJ67-v6.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x19x11x25/SCT-Planning-UA67UJ67-v6.pdf

Open Issues

  • 17.01 Broken thermostat on 13 kA EE
  • 10.01 QPS racks connection
  • 08.01 EE fan rotation problem – all ring
  • 21.12 Aluminum panel for the cables ducts/ shipping and install.

Heat Runs and Test Schedule

08.01 DONE 600A and 13 kA EE system IST.
08.01 DONE RB IST. ON GOING
11.01 DONE 120A powering tests.
15.01 DONE 2 kA powering tests.
15.01 DONE 600A powering tests.
17.01 DONE 600A-EE powering tests.
22.01 - PIC-PC-QPS Interface tests.
23.01 - RQD&F powering tests.
24.01 - RB powering tests.
26.01 - 120A-600A-2kA-RQD/F-RB 8h heat run.
05.02 - 24h heat run test.
07.02 - Polarity test.
14.02 - AUG&UPS test.
- Main.MirkoPojer - 19 Jan 2007 - 11:28

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-17 - 11:24

SCT_UA 67- UJ 67, Jan. 17th 2007

PDF RAT_20070117.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x17x11x19/RAT_20070117.pdf

Open Issues

  • 17.01 Broken thermostat on 13 kA EE
  • 12.01 Power cut in ERD-R13
  • 12.01 Cleaning of filters by CV (planned for 17.01)
  • 10.01 QPS racks connection
  • 08.01 EE fan rotation problem – all ring
  • 21.12 Aluminum panel for the cables ducts/ shipping and install.
  • 21.12 HV Insulation test of the DQS poles

Heat Runs and Test Schedule

08.01 ON GOING 600A and 13 kA EE system IST
08.01 ON GOING RB IST
11.01 DONE 120A powering tests.
15.01 DONE 2 kA powering tests.
17.01 ON GOING PIC-PC-QPS Interface tests.
19.01 DONE 600A powering tests.
23.01 - 600A-EE powering tests.
26.01 - 120A-600A-2kA 8h heat run.
30.01 - RQD&F powering tests.
02.02 - RB powering tests.
07.02 - RQ&RB 8h heat run.
12.02 - 24h heat run test.
14.02 - AUG&UPS test.
14.02 - Polarity test.

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-01-15 - 11:21

RAT for SCT in UA67/UJ67 - Jan, 15

PDF RAT_20070115.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x15x11x07/RAT_20070115.pdf

Open Issues:

  • 15.01 Loosen bolts on RCBYH4.R6B1 and RCBYV4.R6B2 PCs.
  • 12.01 Power cut in ERD-R13
  • 12.01 Cleaning of filters by CV (planned for 17.01)
  • 10.01 QPS racks connection
  • 21.12 Aluminum panel for the cables ducts/ shipping and install.
  • 21.12 HV Insulation test of the DQS poles
Heat Runs and Test Schedule:
  • 08.01 600A and 13 kA EE system IST. ON GOING
  • 08.01 RB IST. ON GOING
  • 11.01 120A powering tests. COMPLETED
  • 15.01 2 kA powering tests. ON GOING
  • 17.01 PIC-PC-QPS Interface tests.
  • 19.01 600A powering tests. COMPLETED
  • 23.01 600A-EE powering tests.
  • 26.01 120A-600A-2kA 8h heat run.
  • 30.01 RQD&F powering tests.
  • 02.02 RB powering tests.
  • 07.02 RQ&RB 8h heat run.
  • 12.02 24h heat run test.
  • 14.02 AUG&UPS test.
  • 14.02 Polarity test.

- Main.MirkoPojer - 15 Jan 2007 - 11:15

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-01-12 - 15:48

RAT for SCT in UA67/UJ67 - Jan, 12

PDF RAT_20070112.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x12x15x43/RAT_20070112.pdf

Open Issues:

  • 12.01 Power cut in ERD-R13
  • 12.01 Cleaning of filters by CV (planned for 17.01)
  • 10.01 QPS racks connection
  • 08.01 EE fan rotation problem to be fixed for the whole machine
  • 21.12 Aluminum panel for the cables ducts/ shipping and install.
  • 21.12 HV Insulation test of the DQS poles
Heat Runs and Test Schedule:
  • 08.01 600A and 13 kA EE system IST. ON GOING
  • 08.01 RB IST.
  • 11.01 120A powering tests.
  • 15.01 2 kA powering tests.
  • 17.01 PIC-PC-QPS Interface tests.
  • 19.01 600A powering tests.
  • 23.01 600A-EE powering tests.
  • 26.01 120A-600A-2kA 8h heat run.
  • 30.01 RQD&F powering tests.
  • 02.02 RB powering tests.
  • 07.02 RQ&RB 8h heat run.
  • 12.02 24h heat run test.
  • 14.02 AUG&UPS test.
  • 14.02 Polarity test.

-- Main.MirkoPojer - 12 Jan 2007 - 15:46

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-10 - 17:47

SCT_UA 67- UJ 67, Jan. 10th 2007

PDF RAT_20070110.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x10x17x43/RAT_20070110.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-08 - 18:46

SCT_UA 67- UJ 67, Jan. 8th 2007

PDF RAT_20070108.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x08x18x45/RAT_20070108.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-21 - 19:11

SCT_UA 67- UJ 67, Dec. 21st 2006

PDF RAT_20061221.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x12x21x19x03/RAT_20061221.pdf

Open Issues:

  • 21.12 EE Fan rotation to be verified
  • 21.12 Aluminum panel for the cables ducts
  • 21.12 HV Insulation test of the DQS poles
  • 18.12 ERD rack powering
  • 18.12 EOD rack powering
  • 18.12 PC powering
  • 18.12 Elettas calibration
  • 18.12 Ventilation station
  • 18.12 Cooling station test by Roese

-- Main.BorisBellesia - 21 Dec 2006 - 19:10

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-11-01 - 09:22

RAT meeting this morning in point 6

Please find enclose the minutes of our last RAT meeting, held this morning in SX6.

Main issue: the 24-h heat run scheduled from Wed 01.11 to Thu 02.11.

Regards

PDF RAT_20061031.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x11x1x09x21/RAT_20061031.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-17 - 12:09

Test on the DQSQD switch, today Friday Nov 17.

Please find enclose a small report of the tests held in UA63 concerning the 13kA switch DQSQD.

Regards

PDF RAT_20061117.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x11x17x12x07/RAT_20061117.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-16 - 14:27

Event analysis, DQSQD switch behavior during the AUG simulation on Friday Nov 3

Please find enclose a short description of the events related to the behavior of the DQSQD 13kA switch during the AUG simulation on Friday Nov 3.

Regards Rosario

PDF MinutesFailureEE_20061117.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x11x16x14x16/MinutesFailureEE_20061117.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-03 - 16:18

Last RAT this morning, Nov 3, in point 6

Please find enclose the RAT meeting minutes of our last meeting in point 6.

Main issues: a) results of the 24-h heat run b) UPS test c) AUG simulation

Regards

PDF RAT_20061103.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x11x03x16x16/RAT_20061103.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-11-02 - 18:02

RAT meeting 02-11-2006

PDF RAT_061102.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x11x02x18x01/RAT_061102.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-01 - 10:11

RAT meeting Nov 1st in CCC

Please find enclose the minutes of the last RAT meeting held today in the CCC.

Remember 24-h heat run today in UA6.

Regards.

PDF RAT_20061101.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x11x01x10x09/RAT_20061101.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-31 - 16:58

RAT meeting this morning in point 6

Please find enclose the minutes of our last RAT meeting, held this morning in SX6.

Main issue: the 24-h heat run scheduled from Wed 01.11 to Thu 02.11.

Regards

PDF RAT_20061031.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x31x16x55/RAT_20061031.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-10-25 - 18:35

RAT minutes 25-10-06

PDF RAT_061025.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x25x18x32/RAT_061025.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-10-23 - 18:18

Visit to P6: LSS and UA63

During the visit today to P6, we took a look at the Beam Dumping System (some collimator, kickers and the TDE line and dump) and the situation in UA63, where we found Giges sorting out the problem with the filters and the Elettas !

JPEG TDE-DumpLine-P6.JPG:
TDE-DumpLine-P6.JPG
TDE line tunnel
JPEG Collimators2-P6.JPG:
Collimators2-P6.JPG
Beam Dumping System Collimators
JPEG TDE-P6.JPG:
TDE-P6.JPG
TDE
JPEG Giges-ElettasRegulation-P6.JPG:
Giges-ElettasRegulation-P6.JPG
Giges fine-tuning of the Elettas in UA63
JPEG Kickers-P6.JPG:
Kickers-P6.JPG
Beam Dumping System Kickers

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-20 - 10:07

RAT minutes Oct 19

Enclose you can find the minutes of our last RAT meeting held in point 6 on Thursday OCt 19.

Main issue: new schedule for the 8-h and 24-h heat runs.

Regards

PDF RAT_20061019.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x20x10x04/RAT_20061019.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-18 - 16:34

RAT in point 6 on Tue Oct 17

Please find enclose the minutes of the last RAT meeting held in point 6 on Tuesday Oct 17.

PDF RAT_20061017.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x18x16x32/RAT_20061017.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-13 - 11:32

RAT for the UA63

Please find enclose the minutes of the RAT meeting held today in point 6.

NB: TSIC proposes to start the next endurance test (RR53) on Dec 04. Please comment.

PDF RAT_20061013.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x13x11x29/RAT_20061013.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-11 - 15:35

RAT of the day: Wed Oct 11.

Please find enclose your RAT minutes of the day...

Main issue: mise en tension of the PCs in UA63.

Regards

PDF RAT_20061011.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x11x15x34/RAT_20061011.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-09 - 10:52

Schedule of the SCT in UA63

PDF SCT-planning-UA63-UJ63-v0.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x09x10x51/SCT-planning-UA63-UJ63-v0.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-09 - 10:51

SCT in UA63. First RAT meeting in Point 6 this morning.

First RAT meeting in Point 6 this morning for the preparation of the SCT in UA63. Enclose the minutes.

Regards

PDF RAT_20061009.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x09x10x49/RAT_20061009.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-10-06 - 09:32

First visit to point 6: supervision before S/C

First visit to point 6. The reason of the visit was to know the situation of the power converters, the cables in the tunnel side and in the power converter side. Also, the objetive was to make a first inspection of what we'll find in the short circuit test.

JPEG P1000018.JPG:
P1000018.JPG
The air cooled cables was already prepared for the short circuit test
JPEG P1000003.JPG:
P1000003.JPG
The water cooled cables still not conected to power converter
JPEG P1000004.JPG:
P1000004.JPG
The water cooled cables still not conected to power converter
JPEG P1000012.JPG:
P1000012.JPG
The labels of the connection between the power converter and the energy extraction system was made with one single label, the cablotheque id number
JPEG P1000020.JPG:
P1000020.JPG
The air cooled cables was already prepared for the short circuit test
JPEG P1000016.JPG:
P1000016.JPG
In the tunnel side, the blocks of the water cooled cables were put in all the cables, and placed in the DFB situation
JPEG P1000002.JPG:
P1000002.JPG
The water cooled cables still not conected to power converter
JPEG P1000017.JPG:
P1000017.JPG
Example of the identification of the water cooled cables.
JPEG P1000010.JPG:
P1000010.JPG
The air cooled cables was already connected to the power converters
JPEG P1000021.JPG:
P1000021.JPG
The water cooled cables of the DFBMM.4L6 was being hung up in the moment of the visit.

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-06-06 - 16:40

Choses vues en UA63 et 67 le 06 06 06...

JPEG espaceursdecablesDCfourreaux.JPG:
espaceursdecablesDCfourreaux.JPG

JPEG IMGP5835.JPG:
IMGP5835.JPG

JPEG IMGP5830.JPG:
IMGP5830.JPG

JPEG IMGP5821.JPG:
IMGP5821.JPG

JPEG UA63060606.jpg:
UA63060606.jpg

JPEG IMGP5833.JPG:
IMGP5833.JPG

JPEG IMGP5837.JPG:
IMGP5837.JPG

JPEG UA67.jpg:
UA67.jpg

JPEG IMGP5827.JPG:
IMGP5827.JPG

Comments:

instantanes ramenes des profondeurs.

-- Main.JeanPierreMalodDognin - 06 Jun 2006

les transfos de RB a decouvert...

-- Main.JeanPierreMalodDognin - 06 Jun 2006

-- Main.JeanPierreMalodDognin - 06 Jun 2006

la derniere photo est faite en UA43,ou on installe "l habillage" de RB...

-- Main.JeanPierreMalodDognin - 06 Jun 2006

les futures connection d eRQD RQF s' impatientent

-- Main.JeanPierreMalodDognin - 06 Jun 2006

espaceurs de cable en fourreaux recherchent travail a temps plein, faire offre sous chiffre UA67

-- Main.JeanPierreMalodDognin - 06 Jun 2006

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-04-03 - 17:12

inauguration ...

Content d'etre le premier a fouler l'espace vierge de ce blog ! Monde non interdit des 2 UA du point 6... ces photos du 3 avril montrent l avancement (?) des travaux...en UA 63..UA67 et UJ63(noter la presence du bel escalier en colimacon en UA 63...)

  • UJ63030406.JPG:
    UJ63030406.JPG

  • RQDUA67.jpg:
    RQDUA67.jpg

  • RQDavenirUA67.JPG:
    RQDavenirUA67.JPG

  • UA67.JPG:
    UA67.JPG

  • UA630304.JPG:
    UA630304.JPG

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-09-28 - 18:29

Warm Magnets Commissioning in point 5. Minutes 28.09.2007

To be found together with the Minutes 28.09.2007 in point 2

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-16 - 17:02

Warm Magnets Commissioning at point 5 - Minutes 11-05-2007

PDF Planning.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2007x05x16x16x40/Planning.pdf

PDF RAT_planning_20070511_WM5.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2007x05x16x16x40/RAT_planning_20070511_WM5.pdf

Comments:

-- Main.MariaPazCasasLino - 16 May 2007

Edit, attach file or add comment


Created by Main.IvanMoyaMartinez on 2006-12-20 - 10:57

The Flags are Up

Polarity Check done in RR57-UA56

JPEG P1000172.JPG:
P1000172.JPG

JPEG P1000173.JPG:
P1000173.JPG

JPEG P1000174.JPG:
P1000174.JPG

JPEG P1000157.JPG:
P1000157.JPG

JPEG P1000154.JPG:
P1000154.JPG

JPEG P1000155.JPG:
P1000155.JPG

Comments:

  • Updated. Cables labeled following the ECR(EDMS 804188) -- Main.IvanMoyaMartinez - 09 Jan 2007 - 11:00

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-18 - 14:27

RAT for Point 5, Dec 18

PDF RAT_20061218.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x18x14x23/RAT_20061218.pdf

Comments: NEXT MEETING WILL BE HELD AT POINT 6 ON THURSDAY, 21

Open Issues for RR57/UJ56

  • 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)
  • 06.12 Permanent labels for RQ converters in RR57 (according to ECR)
  • 07.12 Access key to electrical distribution room in UJ56

Open Issues for UA67/UJ67

  • 18.12 ERD rack powering
  • 18.12 EOD rack powering
  • 18.12 Connection of cable extremities on DQS
  • 18.12 HV test with circulating water
  • 18.12 Water-cooled cables supports
  • 18.12 PC powering
  • 18.12 Water circulation
  • 18.12 Elettas calibration
  • 18.12 Ventilation station
  • 18.12 CYNET01 and CYNET02 powering
  • 18.12 WorldFip cabling for PO
  • 18.12 Cooling station test by Roese
  • 18.12 Electrical protection IP2 on LEMs
  • 18.12 Cleaning of the areas

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-12 - 15:21

RAT for Point 5 in CCC, Dec 12

PDF RAT_20061212.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x12x15x08/RAT_20061212.pdf

PDF UJ56-RR57_IR-Report_2006-12-12_24H-Burnin.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x12x15x08/UJ56-RR57_IR-Report_2006-12-12_24H-Burnin.pdf

Comments: NEXT MEETING WILL BE HELD ON DEC,18 AT POINT 5

Please, find also included the IR analysis rapport, kindly provided by Jeff Thomsen.

Open Issues

  • 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)
  • 06.12 Permanent labels for RQ converters in RR57 (according to ECR)
  • 07.12 Access key to electrical distribution room in UJ56

Heat Runs and Test Schedule

  • 01.12 Squeeze test DONE (all the power converters involved have been tested)
  • 06.12 8-h heat run DONE
  • 07.12 Fire detection test DONE
  • 08.12 Full power ventilation tests in RR57 – from 8:30 till 12:00 DONE
  • 08.12 Afternoon: Polarity Test_1st part TO BE RE-SCHEDULED
  • 11.12 10am: 24-h heat run DONE
  • 13.12 9.00: UPS test and Electric cut (AUG simulation).
  • 13.12 Polarity Test_2nd part TO BE RE-SCHEDULED
  • 14.12 13kA energy extraction test
  • 15.12 14.00: New sequencer test

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-11 - 13:18

RAT for Point5 in CCC, Dec 11

PDF RAT_20061211.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x11x13x11/RAT_20061211.pdf

Comments: Persons on call for assistance during the 24h heat run:

  • Hardware commissioning coordination: Mirko Pojer 162482
  • Power converters :
    • Responsable ABPO: David Nisbet 164223
    • (en cas de absence) Hugues Thiesen 164205
    • Numéro général ABPO-LHC-SCT-24H: 162333
    • Permanences ABPO: Journée 8h a 18 heures 11-12-2006 : Emmanuel Garde (16 4035) / Philippe Marchon Journée 18h a 20 heures 11-12-2006 : David Nisbet (164223) / Jeff Thomsen (165008) Nuit 20h 11-12-2006 a 2h00 12-12-2006 : Okbi Zenasni tel (16 3678) Nuit 2h a 8h 12-12-2006 : Cédric Caffier tel (16 4133) Journée 8h a 17 h 12-12-2006 : Emmanuel Garde (16 4035) / Philippe Marchon
  • Energy extraction system: Knud Dahlerup-Petersen 163279
  • DC cables: J. Kis (10h-18h), B. Lebeau (18h-00h), S. Cotte (00h-7h), J. Kis (7h-end of test) 164879

Open Issues

  • 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)
  • 06.12 Permanent labels for RQ converters in RR57 (according to ECR)
  • 07.12 Access key to electrical distribution room in UJ56

Heat Runs and Test Schedule

  • 01.12 Squeeze test DONE (all the power converters involved have been tested)
  • 06.12 8-h heat run DONE
  • 07.12 Fire detection test DONE
  • 08.12 Full power ventilation tests in RR57 – from 8:30 till 12:00 DONE
  • 08.12 Afternoon: Polarity Test_1st part TO BE RE-SCHEDULED
  • 11.12 10am: 24-h heat run
  • 13.12 UPS test and Electric cut (AUG simulation).
  • 13.12 Polarity Test_2nd part TO BE RE-SCHEDULED
  • 14.12 13kA energy extraction test.
  • 15.12 New sequencer test

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-07 - 12:36

RAT for Point 5 in CCC, Dec 7

PDF RAT_20061207.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x07x12x32/RAT_20061207.pdf

Comments: Next meeting in CCC on Monday Dec,11.

Open Issues

  • 22.11 Fire detection – awaiting certification
  • 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)
  • 06.12 Permanent labels for RQ converters in RR57 (according to ECR)
  • 07.12 Cables re-distribution for 600 A
  • 07.12 Access key to electrical distribution room in UJ56

Heat Runs and Test Schedule

  • 07.12 Fire detection test
  • 08.12 Full power ventilation tests in RR57 – from 8:30 till 12:00
  • 08.12 Afternoon: Polarity Test_1st part
  • 11.12 10am: 24-h heat run
  • 12.12 UPS test.
  • 13.12 Electric cut (AUG simulation).
  • 13.12 Polarity Test_2nd part
  • 14.12 13kA energy extraction test.

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-06 - 12:47

RAT for Point 5 in CCC, Dec 6

PDF RAT_20061206.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x06x12x42/RAT_20061206.pdf

Comments: 8-H HEAT RUN ON GOING

Next meeting: Thursday, 7 at 8.30 in CCC.

Open Issues

  • 22.11 Fire detection – awaiting certification
  • 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)

Heat Runs and Test Schedule

  • 07.12 Fire detection test
  • 08.12 Full power ventilation tests in RR57 – from 8:30 till 12:00
  • 11.12 10am: 24-h heat run
  • 12.12 UPS test.
  • 13.12 Electric cut (AUG simulation).
  • 13.12 Polarity Test
  • 14.12 13kA energy extraction test.

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-05 - 12:58

Special RAT on SCT in RR57/UJ56 and ECR application, Dec 5

PDF RAT_20061205.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x05x12x51/RAT_20061205.pdf

PowerPoint RQ4-RQ10-ECR.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x05x12x51/RQ4-RQ10-ECR.ppt

Microsoft Excel Spreadsheet RR57-UJ56-8H-24H-PC_LIST.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x05x12x51/RR57-UJ56-8H-24H-PC_LIST.xls

Comments:

Please find enclosed the list of PC which will be powered during the 8h heat run and the 24h heat run. The list of steps to be implemented following the ERC (EDMS No.804188) is also included

Open Issues

  • 22.11 Fire detection – awaiting certification
  • 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-01 - 11:49

RAT Point 5, Dec 1

PDF RAT_20061201.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x01x11x42/RAT_20061201.pdf

Comments: Next meeting oin Wednesday, Dec 6, at 8.30 in CCC

Open Issues

  • 23.11 Ripples in RQT12R5B1 power converter.
  • 22.11 Full power ventilation test in RR57 – Friday 8th Dec. 8:30-12:00
  • 22.11 Fire detection – awaiting certification

Heat Runs and Test Schedule

  • 01.12 Squeeze test
  • 06.12 8-h heat run
  • 07.12 Polarity Test
  • 08.12 Full power ventilation tests
  • 11.12 10am: 24-h heat run
  • 12.12 UPS test.
  • 13.12 Electric cut (AUG simulation).
  • 14.12 13kA energy extraction test.

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-11-29 - 12:57

RAT at Point 5, Nov 29

PDF RAT_20061129.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x29x12x54/RAT_20061129.pdf

Comments:

Open Issues

  • 29.11 WIFI problems
  • 29.11 ELETTAS threshold setup
  • 23.11 Ripples in RQT12R5B1 power converter.
  • 22.11 Full power ventilation test in RR57 – Friday 8th Dec. 8:30-12:00
  • 22.11 Fire detection – awaiting certification

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-11-27 - 11:34

RAT at Point 5, Nov 27

PDF RAT_20061127.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x27x11x29/RAT_20061127.pdf

Comments:

Open issues

  • 23.11 Ripples in RQT12R5B1 power converter.
  • 23.11 Cables inversion on rack RYMCA04.
  • 20.11 Problem with water conductivity in DQS: reason known, insulation test still to be performed.
  • 13.11 DC cables high voltage test (RR57 done. Waiting UJ56)
  • 13.11 Test de circulation ED (waiting UJ56)
  • 22.11 Full power ventilation test in RR57 – Friday 8th Dec. 8:30-12:00
  • 22.11 Fire detection – awaiting certification

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-11-23 - 11:53

RAT at Point5, Nov 23

PDF RAT_20061123.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x23x11x08/RAT_20061123.pdf

Comments:

Next RAT on Monday 27

Open Issues

  • 20.11 Problem with water conductivity in DQS
  • 13.11 DC cables high voltage test (RR57 done. Waiting UJ56)
  • 13.11 Ventilation in UJ56.
  • 13.11 Test de circulation ED (waiting UJ56)
  • 20.11 Water conductivity in DQS – checked tomorrow (23.11)
  • 22.11 DQS acousting shielding and cabling – No news
  • 22.11 Full power ventilation test in RR57 – Friday 8th Dec. 8:30-12:00
  • 22.11 Fire detection – awaiting certification
  • 22.11 Wi-Fi damaged cable has to be repaired – no news

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-11-22 - 15:45

Last RAT meeting at Point5 - Wednesday 22

PDF RAT_20061122.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x22x15x44/RAT_20061122.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-20 - 16:21

Last RAT meeting in pint 5, Monday Nov 20.

Please find enclose the minutes of our last RAT meeting in point 5.

Regards

PDF RAT_20061120.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x20x16x18/RAT_20061120.pdf

Comments:

Edit, attach file or add comment


Created by Main.JeanPierreMalodDognin on 2006-11-20 - 10:33

tests en eau des cables refroidis RR57 du 20 novembre: TOUT est OK

JPEG tests_en_eau_cables_refoid_RR57_.jpg:
tests_en_eau_cables_refoid_RR57_.jpg

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-11-17 - 14:41

Essais en court-circuits UJ56 et RR57

Bonjour,

Les visites des installations électriques des ouvrages cités en objet ont été réalisées à ce jour. Aucune non conformité ne s'oppose à ce que les essais en court-circuit soient effectués à partir du 20 novembre 2006. Les rapports de visite parviendront ultérieurement aux responsables de chaque installation.

Cordialement,

Caroline BENOIT-GONIN Norisko Equipements pour SC/GS Office 57-2-024 caroline.benoit-gonin@cernNOSPAMPLEASE.ch

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-15 - 12:00

RAT today in point 5, Nov 15

Please find enclose the last RAT minutes. Meeting held today, Wed Nov 15, in point 5.

Main issue: schedule in UJ56 and RR57.

Regards

PDF RAT_20061115.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x15x11x52/RAT_20061115.pdf

Comments:

Edit, attach file or add comment


Created by Main.JeanPierreMalodDognin on 2006-11-14 - 16:37

RR77, ca avance...

lundi 13 , tests des cables refroidis sans eau?...tout est OK....

JPEG cable_refroidi_RR57JPG.JPG:
cable_refroidi_RR57JPG.JPG

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-13 - 16:35

RAT meeting, Mon Nov 13

Please find enclose the minutes of the last RAT meeeting held this morning, Nov 13, in point 5.

Main issue: the readiness of the RR57 and UJ56, DC cables, elettas, etc..

Regards

PDF RAT_20061113.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x13x16x33/RAT_20061113.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-08 - 16:10

RAT in point 5, Nov 8

Please find enclose the minutes of the last RAT meeting, held this morning in point 5.

Main issues:

- Preparation of the SCT in RR57 and UJ56

- Last tests in UA63

Reg

PDF RAT_20061108.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x08x16x08/RAT_20061108.pdf

Comments:

Edit, attach file or add comment


Created by Main.JeanPierreMalodDognin on 2006-11-03 - 15:39

RR57 en finition

JPEG Tests_cables_refroidis_RR57.jpg:
Tests_cables_refroidis_RR57.jpg

JPEG IMGP8910.JPG:
IMGP8910.JPG

JPEG Redimensionnement_de_IMGP8903.JPG:
Redimensionnement_de_IMGP8903.JPG

JPEG IMGP8938.JPG:
IMGP8938.JPG

JPEG IMGP8950.JPG:
IMGP8950.JPG

JPEG IMGP8944.JPG:
IMGP8944.JPG

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-06 - 11:48

Last RAT meeting in point 5

Enclose the minutes of our last RAT meeting in point 5.

Main subject: 13kA EE switch test.

Also includea a small status of the works in UA63.

Regards

PDF RAT_20061006.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x10x06x11x46/RAT_20061006.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-10-06 - 10:12

Cable and Power Converter Temperatures during the 24-run in RR53 & USC55

Please find Infra Red Analysis of RR53 & USC55 during 24H burnin tests at ultimate current of 2006-10-04. https://edms.cern.ch/file/781901/1/RR53-USC55_IR-Report_2006-10-04_24H-Burnin.pdf

All raw pictures (IR or not) are located in EDMS doc: https://edms.cern.ch/document/781901/1

IR Tunnel Tests EDMS doc should be located in a dedicated place under EDMS Tunnel installation structure to compare results if needed between locations measured. Note that UA83 is at following address: https://edms.cern.ch/document/672600/1.

Jeff Thomsen & Yves Thurel

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-05 - 10:11

24 hours heat run in point 5

Please find enclose the minutes of the last RAT meeting in point 5.

Main subject: first results of the 24-h heat run oin RR53 and USC55.

Regards

PDF RAT_20061004.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x10x05x10x09/RAT_20061004.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-03 - 17:48

Last news from point 5 - 24-h heat run

Please find enclose the last minutes of the RAT meeting concerning the 24-h heat run in RR53 and USC55.

PDF RAT_20061003.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x10x03x17x46/RAT_20061003.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-10-02 - 11:23

DC connection sur 5kA Inner Triplet à USC55

Allo Roberto!

Le systeme de detection de defaut de terre interne au convertisseur a encore passe avec succes le TEST DE CONTROLE...

Bonne soiree Hugues

PS: JC Guillaume est en train de traiter le pb.

Bonjour,

Suite à un défaut "Earth Fault" affiché sur le convertisseur 5kA Inner Triplet HCRPHGC__-KP000003, nous avons investigué le problème. Il s'est avéré être un court-circuit franc entre le DC Output du convertisseur et le GND (Il n'y avait pas la tension Vearth en mode actif et un test à l'ohmmètre a indiqué 0ohm entre le DC- et le GND). Après un rapide coup d'oeil, nous nous sommes aperçu que les vis permettant de raccorder le câble à eau sur DC-, touchaient le capot de protection des diodes de roue-libre (voir photos jointes). Le nécessaire a été fait pour contacter les personnes responsables de ces connections et les modifications seront faites dès demain.

Bonne soirée

Amitiés

Vincent

JPEG IMGP2675.JPG:
IMGP2675.JPG

JPEG IMGP2673.JPG:
IMGP2673.JPG

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-10-02 - 11:20

eau gl. 4-5

And here it is the demin. water plot.

The power is very low (mean power around 200 kW), which is creating some pumping of the regulation (the reg. valve is opening and closing) as the nominal power of the circuit is 1400 kW. We are going to slow the PI to see if we get a smoother plot.

JPEG glacee_sect45_29_09_2006.jpg:
glacee_sect45_29_09_2006.jpg

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-10-02 - 11:19

hwc

Please find below the plot for RR53, today at 13:21 hours. The system is working according to spec. (high speed since about 10:00 h). Keep in mind that the nominal values requested by specification are: Temp (mean) ground floor: 28 (+- 5) C Temp. (mean) top floor: 30 (+- 5) C

The measured values are (at 14:04 h) Temp. (mean) GF: 24.2 C Temp (mean) top : 30 C (vault) Temp mean regulation: 27.1 C

Water monitoring in UW45 is also underway and according to spec. Plot in the coming minutes.

We are at this moment discussing with J.C. Guillaume (on site) some adjustmemts to try and improve the temperature of his cables.

JPEG Outlook.jpg:
Outlook.jpg

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-02 - 10:38

RAT meeting, october 2, 2006

Please find enclose the minutes of the last RAT meeting, held in CCC this morning.

Main subject: last 8-h heat run and cables DC overheating.

Regards

PDF RAT_20061002.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x10x02x10x37/RAT_20061002.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-09-29 - 11:03

RAT Sept 29 8-h heat run

Please find enclose the minutes of the last RAT meeting held this morning in CCC.

Main subject: 8-h heat run launched this morning 8h30. It concerns 120, 600A and ineer triplets.

Regards

PDF RAT_20060929.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x09x29x11x01/RAT_20060929.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-09-26 - 17:10

RAT 26 sept 2006

Minutes from the last RAT meeting in point 5.

Regards

PDF RAT_20060926.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x09x26x17x09/RAT_20060926.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-09-21 - 12:22

RAT du 21septembre 06

Please find enclose the minutes of the last RAT meeting held uin point 5 this morning 8h30.

NB: new schedule for the tests. 8h heat run moved to Friday sept 29.

Regards

PDF RAT_20060921.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x09x21x12x18/RAT_20060921.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-09-19 - 10:57

RAT 120 and 600A powering tests

Last minutes of the RAT meeting held today in point 5.

Main subjects: AC cabling, cooling and preparation of the next 8h run.

Regards

PDF RAT_20060919.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x09x19x10x55/RAT_20060919.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-09-18 - 14:26

RAT today...

Last news from the field.

Minutes of the meeting held today in point 5.

Regards

PDF RAT_20060918.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x09x18x14x24/RAT_20060918.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-09-14 - 18:17

RAT minutes..

Enclosed the last RAT minutes.

PDF RAT_20060914.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x09x14x18x16/RAT_20060914.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-09-13 - 16:49

RAT in point 5 today

Please find enclose the minutes of the RAT meeting, hed today in point 5.

New version of the SCT schedule in this area also included .

Regards

PDF RAT_20060913.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x09x13x16x47/RAT_20060913.pdf

PDF SCT-planning-RR53-USC55_v20060913rp.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x09x13x16x47/SCT-planning-RR53-USC55_v20060913rp.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-09-11 - 15:12

First RAT in point 5

Next RAT meeting will be held in point 5...

The first one in this area.

The SCT will concern RR53 and USC55.

8h30, bld. 3595/R-013.

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-11-20 - 19:26

Shift HC teams and Piquet numbers - 20th Nov. 2007

PDF shifts.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x11x20x19x25/shifts.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-11-19 - 20:48

RAT CCC: Powering tests of Sector45 - 19 Nov. 2007

PDF RAT_20071119.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x11x19x20x47/RAT_20071119.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-11-14 - 18:30

RAT CCC: preparation for Powering test: Sector45 - 14th Nov.

PDF RAT_20071114.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x11x14x18x29/RAT_20071114.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-11-13 - 19:09

RAT Point4 - Preparation for Powering Sector45 - 9th Nov. 2007

PDF RAT_20071112.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x11x13x19x07/RAT_20071112.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-10-31 - 10:40

From Pressure Test to Cooldown - Sectors 45 and 34 - 30th October

Pptx Sector45_34_status_20071030.pptx:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x10x31x10x10/Sector45_34_status_20071030.pptx

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-10-23 - 12:24

From pressure test to Cooldown - Sector45 and 34 - 23rd October

Pptx Sector45_34_status_20071023.pptx:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x10x23x12x22/Sector45_34_status_20071023.pptx

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-10-23 - 12:21

From pressure test to cooldown - 23rd October 2007

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-10-23 - 12:15

Preparation for cooldown - RF-Powering tests-Installation activities

PDF RAT_20071022.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x10x23x12x14/RAT_20071022.pdf

PowerPoint Sector45_34_RF-PT_planning.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x10x23x12x14/Sector45_34_RF-PT_planning.ppt

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-10-16 - 16:06

From Pressure test to Cooldown - 16th October - Point4

Pptx Sector45_34_status_20071015.pptx:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x10x16x15x57/Sector45_34_status_20071015.pptx

Sectors 45 and 34

  • Minutes: see attached presentation

  • Sector34 Open Issue: Needed strict coordination among AT-MCS, AT-VAC and AT-ACR to define priorities of DSL interconnection and 7L3 leaks localization/repair. A. Perin is the expert to who address technical questions.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-10-16 - 14:24

From Pressure test to CooLdown - 16th Oct.

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-10-10 - 20:21

Warm Magnets Commissioning Point 4. Minutes 09-10-2007

PDF Warm_Magnets_Point4_Minutes.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x10x10x20x20/Warm_Magnets_Point4_Minutes.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-10-09 - 20:16

From Pressure Test to Cool-Down - Tuesday 09 - Point 4

SECTOR45

General planning:

  • W41-W42: L5 to 80K
  • W42-W43: Cool-down of the whole sector from 80K up to 4.5K
  • W44: Cool-down of the arc45 from 4.5K up to 1.9K / ElQA of the Circuits of the matching sections / Powering tests of the matching sections
  • W45: ElQA of the circuits of the arc and Powering tests of the matching sections
  • W46: Powering tests of the circuits of the arc

Activitites - w41:

  • Today -09/10 - cooldown L5 to 80K
    • Sector closed since 08/10
    • Access allowed with key and tokens except RE52-UL55
    • See document LHC-S-ES-0017 for particular access conditions
  • ElQA break-down monitoring all along cooldown; AT-MEL
  • QPS activities; AT-MEL
  • CMS vacuum installation/tests; AT-VAC
  • Magnet alignment; TS-SU
  • Cable campaign point5; TS-EL
  • Inner triplet interconnection; AT-MCS

Open Issues + Long term activitites:

  • Swapped cables of RB circuit on DFBA RR53
  • Cleaning – W43/44
  • Collimator installation W44
  • CMS-ZDC optical fiber installation; W44

SECTOR34

General planning+Activities:

  • W40: Pressure test of the arc / DSL still disconnected
  • W41-W42:
    • Consolidation
      • All lines under nominal pressure up to Wed. 10/10 - AT-VAC/ACR
      • Cryo-valves reconfiguration Thu.-Fri. 11-12/10 - AT-ACR
      • Opening and repairs of leaks (two leaks found: 7R4 ~2*10^-6 [mbar*l/s] and 7L4 ~10^-5 [mbar*l/s]) - AT-MCS/VAC
    • Checks the jacks - stable pressure conditions needed - 2 days needed - date to be decided, possibly on Friday 12/10 - TS-IC
    • ElQA.DOC from Wed. 10/10 - AT-MEL
    • Manometer installation on the ceiling at the level of RE38 – trolley passage difficult – TS-IC
  • W43: DFBLC-DSLC-DFBAF interconnections start (possibility to start during W42 if compatibility with consolidation)
  • W50: Flushing

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-09-25 - 16:44

Sector45 - Access conditions and cool-down co-activities

PDF Access_restart_cd_20070925.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x09x25x16x43/Access_restart_cd_20070925.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-06-29 - 16:33

Activities foreseen in sector 45 from Monday 2nd July up to the cooldown

Activities in sector 45 from Monday 2nd July

  • CRYO
    • Purging on Friday 29th evening and all the week end long;
    • Refilling on Monday 2nd July;
  • ElQA-TP4B and Monitoring installation
    • On Tuesday 3th July, once 6Bar conditions are reached, ElQA will be repeated (only on RB) maybe from 8am to 1pm. RESTRICTED ACCESS.
  • CRYO-instrumentation
    • Coherence test of all the valves in ARC and LSS (Thursday 28th and Friday 29th);
    • Commissioning;
    • Initialization of the CV910 valves on Tuesday 26th and Wednesday 27th;
  • TS/EL intervention
    • Changing of the connectors of a BPM (LSS4R);
    • Connection check of the instrumentation cables of BCT;
  • TIM camera
    • General check on Monday 2nd July;
    • Run from P4 to P5, during the Cryo Patrouille the morning of the start of the cool-down;
  • Safety Inspection
    • Monday 2nd July at 2pm;
  • AT/ACR Patrouille
    • A Patrouille will be carried out on Wednesday 4th July just before the cooldown.
  • AT/VAC
    • Short access to LSS4R and LSS5L on Monday 2nd July.
  • AB/BI
    • LSS4R short intervention on the BI (capeau).

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-06-27 - 18:52

going on activities in sector 45

Going on activities in sector 45

  • DFBAH and MB (diode)
    • checked and closed by Wednesday 27th;
    • Leak test on Thursday 28th;
    • Purging on Friday 29th evening and all the week end long;
    • Refilling on Monday 2nd July;
    • On Tuesday 3th July, once 6Bar conditions are reached, ElQA will be repeated (only on RB);
  • CRYO-instrumentation
    • Initialization of RF valves on Wednesday 27th or Thursday 28th;
    • Coherence test of all the valves in ARC and LSS (Thursday 28th and Friday 29th);
    • Commissioning;
    • Initialization of the CV910 valves on Tuesday 26th and Wednesday 27th;
  • TS/EL intervention
    • Changing of the connectors of a BPM (LSS4R);
    • Placing of some Radmon cables on Wednesday 27th, Thursday 28th and Friday 29th (LSS5L);
    • Connection check of the instrumentation cables of BCT;
  • TIM camera
    • General check on Monday 2nd July;
    • Run from P4 to P5, during the Cryo Patrouille the morning of the start of the cool-down;
  • Safety Inspection
    • Monday 2nd July at 2pm;
  • RAMSES
    • Check of the equipment on Tuesday 26th and Wednesday 27th.
  • CV Intervention
    • Thursday 28th June from 2pm to 4pm
  • QUI
    • Caps removal, sleeve installation and then closure of the QUI performed on Thursday 28th.
    • The new mechanism for the header B valve is being prepared and should be ready for Friday 29th or early next week.

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2007-06-26 - 18:47

Diagnosis of MB.A45 circuit shorted to ground - Courtesy of D.Bozzini & J.P.Tock

This morning MCS has open the diode box of the dipole 3186. AT/MEL has connected an ohmmeter during the operation to follow the status of the short. At the time of opening the big flange of the diode box the resistance to ground started to change (increase). After the opening it could be possible to first see a quite big quantity of dust on the diode metallic box and second some big pieces of welding residuals.

After having disconnected the diode at the level of the half moons the short disappeared (infinite resistance) and in at the same time some debris felled down from the top part of the half moons. As you can see in the attached pictures the size of one of this residual fits perfectly between the live part and the metallic body of the bus bar duct. In addition in one picture you will see a carbon trace on the G10 insulating piece. This could be provoked by the current of 1A we injected to better localize the fault position flowing through the debris supported onto the G10 piece.

The system will be set back to the requested pressure condition (6 bars in the magnets, 2.2 bars in the DFBA's - which should be achieved by Tuesday 3rd July), when the ELQA-TP4B qualification of this faulty circuit will be repeated.

JPEG IMG_0871.jpg:
IMG_0871.jpg

JPEG IMG_0933.jpg:
IMG_0933.jpg

JPEG IMG_0909.jpg:
IMG_0909.jpg

JPEG IMG_0882.jpg:
IMG_0882.jpg

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2007-06-13 - 18:14

Beam Instrumentation - BSRT Installed and Aligned

The two BSRT telescopes were Installed last Monday in LSS4.

LSS4L: The BSRTD has been connected to the BSRTM and the system successfully aligned with the laser.

LSS4R: The vacuum chamber in IP4 right side is still intact after the BGI bake-out incident and therefore connection and alignment of the BSRT right side will be done tomorrow !

J.J. Gras for the BI Team

JPEG BSRTD_C5L4_Front1.jpg:
BSRTD_C5L4_Front1.jpg

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-06-08 - 18:42

RAT in Point 4 - Minutes of 8th June 2007

PDF RAT_20070608.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x06x08x18x41/RAT_20070608.pdf

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2007-06-04 - 00:37

RAT in P4 - Minutes 1st June 2007 (3 weeks for Cool-Down !)

PDF RAT_20070601.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x06x04x00x35/RAT_20070601.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-22 - 18:34

RAT in Point 4 - Minutes of 22nd May 2007

PDF RAT_20070522.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x05x22x18x33/RAT_20070522.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-22 - 18:33

RAT in Point 4 - Minutes of 11th May 2007

PDF RAT_20070511.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x05x22x18x32/RAT_20070511.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-07 - 18:38

RAT in Point 4 - Minutes of 4th May 2007

PDF RAT_20070504.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x05x07x18x37/RAT_20070504.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-04-27 - 17:47

RAT in Point 4 - Minutes of 27th April 2007

PDF RAT_20070427.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x04x27x17x39/RAT_20070427.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-04-20 - 18:36

RAT in Point 4 - Minutes of 20th April 2007

PDF RAT_20070420.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x04x20x18x26/RAT_20070420.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-04-17 - 18:52

RAT in Point 4 - Minutes of 13th April 2007

PDF RAT_20070413.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x04x17x18x51/RAT_20070413.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-03-30 - 18:47

RAT in Point 4 - Minutes of 30th March 2007

PDF RAT_20070330.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x03x30x18x45/RAT_20070330.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-03-28 - 09:50

RAT in Point 4 - Minutes of 23th March 2007

PDF RAT_20070323.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x03x28x09x49/RAT_20070323.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-03-21 - 10:56

RAT in point 4 - Minutes of 16th March

PDF RAT_20070316.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x03x21x10x54/RAT_20070316.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-03-12 - 18:57

RAT in Point 4 - Minutes of 9th March 2007

PDF RAT_20070309.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x03x12x18x56/RAT_20070309.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-02-23 - 17:03

RAT in point 4 - minutes of the 23th February 2007

PDF RAT_20070223.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x02x23x17x01/RAT_20070223.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-02-16 - 18:46

RAT in point 4 - minutes of 16th February 2007

PDF RAT_20070216.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x02x16x18x45/RAT_20070216.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-09-25 - 17:30

Dry-Run RAT 06-09-21

PDF Z:\Antonio\Dry-Run\RAT\DR_RAT_060921.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x09x25x17x28/Z:\Antonio\Dry-Run\RAT\DR_RAT_060921.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-30 - 10:33

24-h run UA43: RATs first round table

Minutes of the meeting held this morning as usual 8h30 in CCC.

First round table related to the 24 hours heat run in UA43 and UJ43.

Have a nice day

PDF RAT_20060830.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x30x10x31/RAT_20060830.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-29 - 16:04

RATs: 24-h heat run completed today 29.08 at 15h20

Last news from the field (UA43 and UJ43)

24-h run completed today at 15h20

Please find enclose few more details concerning these last two days of HC activity in point 4.

Regards

PDF RAT_20060829.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x29x16x01/RAT_20060829.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-28 - 11:17

RAT 24-hours test today

Minutes of the last RAT meeting

Today 24-h heat run in UA43 from 14h00

PDF RAT_20060828.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x28x11x14/RAT_20060828.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-25 - 11:07

RAT: first analysis of the 8h-run and RR53

Minutes of the meeting, including:

1) First analysis related to the 8h-run (24.08 from 9h27 to 17h05)

2) Round table concerning the work advancement in RR53

PDF RAT_20060825.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x25x11x04/RAT_20060825.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-24 - 11:03

RAT: 8 hours run today..

Minutes from the last meeting in CCC, mainly related to the 8 hours run (in progress)..

PDF RAT_20060824.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x24x11x01/RAT_20060824.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-23 - 11:29

Last from RATs...

Please find enclose the minutes of the last RATs meeting, held in point 4 this morning 8h30, for the preparation of the LHC short-circuit test in UA43.

PDF RAT_20060823.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x23x11x27/RAT_20060823.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-21 - 10:34

News from tunnel RATs

Here you are fresh news from tunnel RATs.

Enclose the minutes of last RAT meeting, held today in SX4.

Infos related to the interlock tests and the next 8h-run.

Regards

PDF RAT_20060821.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x21x10x28/RAT_20060821.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-17 - 11:18

RAT 4you

Minutes of last RAT meeting this morning in point 4. Main issues: next 8h run and 24h run end of month.

PDF RAT_20060817.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x17x11x16/RAT_20060817.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-15 - 10:59

Last RAT news: interlock test RQD and RQF

Please find enclose the minutes of the last RAT meeting, held in point 4 this morning. The doc includes infos related to the last interlock test and the future 8hrs-run.

PDF RAT_20060815.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x15x10x57/RAT_20060815.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-14 - 11:00

RAT meeting today..

Please find enclose the last RAT meeting minutes.. 1st 8h run in UA43, interlock test for RQD and RQF, and next steps.

PDF RAT_20060814.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x14x10x57/RAT_20060814.pdf

Comments:

Edit, attach file or add comment


Created by Main.JeanPierreMalodDognin on 2006-07-17 - 15:25

des tests HTdes cables refroidis de UA43

Les tests d isolement des cables refroidis de l UA43 ne s averent pas tous positifs... Apres bien des interrogations , il semble que l eau contenue dans les cables soit de tres mauvaise qualite et perturbe considerablement les mesures ... Apres discutions avec Loic Bougeois, il ne semble pas d actualite de vidanger ces cables refoidis mais plutot de faire circuler l eau demineralisee en vue de la debarrasser des impuretes qu elle contient... Travail effectue apres la pose des eletas...et le raccordement aux switchs.

JPEG sertissage__trop__pres___RA43.jpg:
sertissage__trop__pres___RA43.jpg

JPEG sertissage_trop_pres__RA43.jpg:
sertissage_trop_pres__RA43.jpg

Comments:

  • et puis nous avons finalement vidange les cables et tout est rentre dans l ordre le jour suivant....les resistances d isolement sont conformes a ce qu on attendait... -- Main.JeanPierreMalodDognin - 21 Jul 2006 - 16:43

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-04-06 - 15:17

proche des masques Africains?

un detail de DQS , pour rever un peu...

Comments:

JPEG DQSdetail.JPG:
DQSdetail.JPG

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-03-30 - 08:53

RAT meeting minutes 3rd of March 2006

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-03-30 - 08:50

RAT meeting minutes 2nd of March 2006

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-03-30 - 08:48

RAT meeting minutes 1st of March 2006

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-03-30 - 08:49

RAT meeting minutes 27th of February 2006

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-03-30 - 08:46

RAT meeting minutes 24th of February 2006

Comments:

Edit, attach file or add comment


Created by Main.RobertoSaban on 2006-03-30 - 08:36

Field Control Room in UA43

At the request of Sebastien Pelletier from TS-MEL/Cabling Section, I asked Pierre Charrue and his team to remove all the computing equipment in UA43. The PCs, monitors and the big screen are now stored by AB-CO. A request to re-install this must be made to them a month in advance of the start of the Short Circuit Tests in UA43 en of July 2006.

Comments:

-- Main.JeanPierreMalodDognin - 05 Apr 2006

Et, hier 4 avril...pour en rajouter une couche(!) dans le demontage(provisoire?) de le SdeC,j ai -apres demande de Rosario- demonte les baches qui en cachait l interieur ...la video surveillance est toujours operationnelle et l acces par code conserve.

-- Main.JeanPierreMalodDognin - 05 Apr 2006

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-04 - 18:55

SCT in UJ33

PowerPoint GlobalAnalysis24hHR_UJ33.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint3x2007x05x04x18x27/GlobalAnalysis24hHR_UJ33.ppt

  • During the 24hr heat run major problems were found (for a more detailed description see the attached file):
1 - Over-heating of the systems in UJ33 (central part) – TS-EL will re-routing the cables inside the UJ33.

2 - Over- heating of the cables in the tunnel side RZ33 – a vertical cable tray will be positioned parallel to the TS-CV cooling water supports. The tray will host one of the three cable layers that are now placed on the same tray.

  • The re-routing cable works will last approximately up to the third week of June. The integration/cost studies are on going and the possible co-activities interfering with the preparation for the cooldown of the A34 are being taken into account.
A meeting to up-date the "status of the art" is foreseen for Monday 14th. Details will be given.

  • During the inspection of the UJ33 P. Proudlock commented:

1 - ODH system must be foreseen in UJ33 - the DFBLC inside the UJ makes it indispensable

2 - Rectangular ducts (for the geometers) must be closed during the heat-run

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-04-18 - 13:19

SCT in UJ33 - 18th April

PDF RAT_20071704.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint3x2007x04x18x13x15/RAT_20071704.pdf

********************************** **********************************

In the Minutes:

* Resume of the two 8hr heat runs

* List of the contact people of the concerned system during the 24hr heat run

********************************** **********************************

Schedule:

22.03 DONE Polarity Check
30.03 DONE CCC tests: preparation for the 8hr run
16.04 DONE First 8hr HEAT RUN ; re-scheduled
17.04 DONE Second 8hr HEAT RUN
18.04 - CCC tests: preparation for the 24hr run
19.04 - 24hr HEAT RUN
20.04 - AUG - UPS test
21.05 - EE 4hr run

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-04-13 - 15:28

SCT in UJ33 - 12th April

PDF RAT_20071204.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint3x2007x04x13x15x22/RAT_20071204.pdf

PDF SOC_settings_8hrun_UJ33.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint3x2007x04x13x15x22/SOC_settings_8hrun_UJ33.pdf

********************************** **********************************

1 - TS-CV confirmed that the ED circuit of the sector 2-3 will be normally operated the 16th-17th and the 19th.

********************************** **********************************

2 - In attachment: the current of each PC in UJ33 in the 4 steps of the 8hr run.

* HEAT RUNS Schedule:

22.03 DONE Polarity Check
30.03 DONE CCC tests: preparation for the 8hr run
16.04 - First 8hr HEAT RUN ; re-scheduled
17.04 - Second 8hr HEAT RUN
18.04 - CCC tests: preparation for the 24hr run
19.04 - 24hr HEAT RUN
20.04 - AUG - UPS test
21.05 - EE 4hr run

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-04-04 - 12:31

SCT in UJ33 - 4th April

PDF RAT_20070404.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint3x2007x04x04x12x30/RAT_20070404.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-04-02 - 19:02

SCT in UJ33 - 2nd April

PDF RAT_20070402.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint3x2007x04x02x18x56/RAT_20070402.pdf

Comments: HEAT RUN CANCELLED

  • Maintenance of the water-cooling circuits coming form Point 2 providing water also to the cooling circuits of the power converters in UJ33 was planned to be performed between 02.04 (today) and the 20.04. This manipulation was not taken into account during the planning of the SCT in UJ33.

  • The maintenance caused the stop of the tests and was irreversible: it will be not possible recuperate the cooling power until the end of the intervention.

  • The tests has been rescheduled in agreement with TS-CV. The priority will be given to the work in SX2 (the primary circuit) in order that the water in UJ33 will be available the 16.04.

HEAT RUNS Schedule:

22.03 DONE Polarity Check
30.03 DONE CCC tests: preparation for the 8hr run
2.04 CANCELLED First 8hr HEAT RUN
16.04 - First 8hr HEAT RUN ; re-scheduled
17.04 - Second 8hr HEAT RUN
18.04 - CCC tests: preparation for the 24hr run
19.04 - 24hr HEAT RUN